### v2

# Finite-Time Behavior of Erlang-C Model: Mixing Time, Mean Queue Length and Tail Bounds

Abstract.
Service systems like data centers and ride-hailing are popularly modeled as queueing systems in the literature. Such systems are primarily studied in the steady state due to their analytical tractability. However, almost all applications in real life do not operate in a steady state, so there is a clear discrepancy in translating theoretical queueing results to practical applications. To this end, we provide a finite-time convergence for Erlang-C systems (also known as $M/M/n$ queues), providing a stepping stone towards understanding the transient behavior of more general queueing systems. We obtain a bound on the Chi-square distance between the finite time queue length distribution and the stationary distribution for a finite number of servers. We then use these bounds to study the behavior in the many-server heavy-traffic asymptotic regimes. The Erlang-C model exhibits a phase transition at the so-called Halfin-Whitt regime. We show that our mixing rate matches the limiting behavior in the Super-Halfin-Whitt regime and matches up to a constant factor in the Sub-Halfin-Whitt regime.
To prove such a result, we employ the Lyapunov-Poincaré approach, where we first carefully design a Lyapunov function to obtain a negative drift outside a finite set. Within the finite set, we develop different strategies depending on the properties of the finite set to get a handle on the mixing behavior via a local Poincaré inequality. A key aspect of our methodological contribution is in obtaining tight guarantees in these two regions, which when combined, give us tight mixing time bounds. We believe that this approach is of independent interest for studying mixing in reversible countable-state Markov chains more generally.
Queuing Theory, Drift Method, Markov Chain mixing
copyright: acmlicensedjournalyear: 2025copyright: rightsretainedconference: Abstracts of the 2025 ACM SIGMETRICS International Conference on Measurement and Modeling of Computer Systems; June 9–13, 2025; Stony Brook, NY, USA.booktitle: Abstracts of the 2025 ACM SIGMETRICS International Conference on Measurement and Modeling of Computer Systems (SIGMETRICS Abstracts ’25), June 9–13, 2025, Stony Brook, NY, USAisbn: 979-8-4007-1593-8/25/06doi: 10.1145/3726854.3727287journal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Mathematics of computing Probability and statisticsccs: Applied computing Operations research
1. Introduction
Queueing theory has proved to be a successful tool for studying several service systems like analyzing data centers (Maguluri and Srikant, 2014; Maguluri et al., 2014, 2012), wireless networks (Nguyen and Modiano, 2023, 2024a, 2024b), and ride-hailing and EV systems (Varma et al., 2021, 2023b, 2023a). A popular approach is to study such systems in the steady-state owing to the tractability of analysis. However, these systems rarely operate in a steady state and so, the theoretical analysis does not directly translate into real-life performance guarantees. Furthermore, a typical goal in the performance analysis of service systems is to understand queue length/waiting time behavior. However, it is not possible to obtain the exact queue length distribution except in special cases such as the $M/M/1$ queue (Robert, 2003; Leguesdron et al., 1993; Morse, 1955). So typically, one uses the steady-state behavior as a proxy for the finite-time behavior. However, this approximation is useful only when the system mixes quickly, i.e. approaches the steady-state behavior quickly. These discrepancies motivate us to analyze the rate of convergence to the steady state. Such a result allows one to better understand the applicability of the steady-state results. As the transient analysis of queueing systems is known to be challenging, we restrict our attention to one of the simplest systems, the Erlang-C model (also known as the $M/M/n$ queue).
Moreover, sometimes it is not even possible to obtain the closed form steady-state distribution of the system in many queuing systems, and therefore one resorts to asymptotic analysis such as heavy traffic analysis (Hurtado-Lange and Maguluri, 2020). Recent works have characterized the rate of convergence in these regimes, thus obtaining a handle on the pre-asymptotic queue length behavior (Jhunjhunwala et al., 2023; Braverman et al., 2017). Since these approximation errors are larger when one is further away from the asymptotic regime, and therefore one has a better handle on the stationary distribution when one is closer to the asymptotic regime.
However, prior works suggest that the queuing systems mix slower as one gets closer to the asymptotic regimes, for example, in the Heavy Traffic regime (Pang et al., 2007; Robert, 2003; Gamarnik and Goldberg, 2013). And so, we have a dilemma where the system has a larger steady-state approximation error if it is further away from the asymptotic regime, but mixes slower when it is closer to the asymptotic regime. And so, it is important to characterize the goodness of the steady-state analysis and asymptotic analysis to the finite-time queue length behavior. Therefore, the goal of this paper is to obtain queue length bounds in finite-time and pre-asymptotic regimes (i.e. with a finite number of servers) for the $M/M/n$ queue.
1.1. Contributions
In this work, we consider the Erlang-C system with $n$ servers whose service time is exponentially distributed with mean $1$ and the arrival rate $\lambda_{n}=n-n^{1-\alpha}$ for some many-server heavy traffic parameter $\alpha>0$. It is known that the Erlang-C system exhibits a phase transition at $\alpha=1/2$, and the regimes $\alpha\in(0,1/2)$ and $\alpha\in(1/2,\infty)$ are called the Sub-Halfin-Whitt and the Super-Halfin-Whitt regime respectively, each with a distinct behavior. Therefore, our results and our analysis are different for each regime. That being said, our contributions can be summarized as follows.
Tight mixing of $M/M/n$: We obtain tight convergence to stationarity results at time $t$ in terms of Chi-square distance for finite time and a finite number of servers. In particular, let $\pi_{n,t}$ and $\nu_{n}$ be the queue length distribution of the $M/M/n$ system at time $t$ and at the steady-state (i.e. $t\rightarrow\infty$) respectively, we obtain the convergence results in the form of
$\chi\left(\pi_{n,t},\nu_{n}\right)\leq e^{-\square t}\chi\left(\pi_{n,0},\nu_{n}\right)$
where the $\square$ term is called the mixing rate of the system and $\chi$ is the square root of the Chi-square distance, which would be defined in Subsection 2.2. Our results are true for all time $t$ and for all $n$ (number of servers). Furthermore, our results are tight in the sense that as $n$ goes to infinity, our rate matches the mixing rate of $M/M/1$ and $M/M/\infty$, depending on the regime. We obtain finite-time convergence results for the following regimes:
  • Super-Halfin-Whitt regime ($\alpha\in(1/2,\infty)$): In this regime, we show the convergence rate to stationarity to be $e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}$ where $C_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}C_{n}=1$ for $\alpha\in(1/2,1)$ and $C_{n}=1$ for $\alpha\in[1,\infty)$. In particular, our result for $\alpha\in[1,\infty)$ recovers the bound in (Zeifman, 1991). Moreover, to the best of our knowledge, our result in the regime $\alpha\in(1/2,1)$ is new in the literature and is a contribution of this paper, covering the gap $\alpha\in(1/2,1)$ between (Gamarnik and Goldberg, 2013; Zeifman, 1991). Our results are generalizations of the finite-time behavior of $M/M/1$ since they match the limiting behavior of $M/M/n$ in this regime, which is expected to resemble the behavior of $M/M/1$ with a service rate $n$ (whose mixing rate is known to be $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ (Morse, 1955; Van Doorn, 1985; Robert, 2003; Gamarnik and Goldberg, 2013)).
  • Sub-Halfin-Whitt regime ($\alpha\in(0,1/2)$): We show that the mixing rate of $M/M/n$ queue in the Sub-Halfin-Whitt regime is $D_{n}$, where $D_{n}$ is a positive parameter satisfying $\lim_{n\rightarrow\infty}D_{n}=1/25$, which matches the asymptotic behavior up to a constant factor. To the best of our knowledge, we are the first to achieve a constant mixing time guarantee for $M/M/n$ in the Sub-Halfin-Whitt regime. Moreover, if we have $\lambda_{n}\in\mathbb{Z}^{+}$ then we show that one can match the asymptotic behavior at the limit, i.e. the mixing rate approaches $1$ as in the $M/M/\infty$ system (Chafaï, 2006). We conjecture that the condition $\lambda_{n}\in\mathbb{Z}$ is not required, and getting a mixing rate approaching $1$ without this condition will be a future work.
  • Mean Field regime $\left(\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c\in[0,1)\right)$: Finally, in this regime, we show the mixing rate of the system approaches $1-\sqrt{c}$ as $n$ goes to infinity. When $c=0$, we have the mixing approaches $1$, which matches the mixing rate of $M/M/\infty$ at the asymptotic. Notably, this result does not rely on the arrival rate being an integer to match the mixing rate of $M/M/\infty$ at the asymptotic.
These results are formally stated in Theorem 3. To the best of our knowledge, our work is the first to establish Chi-square convergence results for queueing systems in contrast to most previous works that used TV distance and Wasserstein distance which would not be sufficient to establish finite-time statistics. On the other hand, the obtained Chi-square distance will allow us to obtain finite-time behavior characterizations of $M/M/n$ systems (i.e. the Erlang-C system), such as:
  • Mean queue length: we obtain an upper bound of the distance between the finite-time mean queue length and the steady-state mean queue length. The rate of which the finite-time mean queue length converges to the steady-state mean queue length is dictated by the mixing rate of the system. We formally state this result in Corollary 4.
  • Finite-time concentration bound of the number of customers: we obtain a finite-time concentration bound for the queue length. Our tail bound results match the steady-state and asymptotic behavior in (Jhunjhunwala et al., 2023; Braverman et al., 2017). The queue length concentration bound results are formally stated in Corollary 5.
  • Probability of having an idle server: We obtain a bound on the probability of having an idle server at time $t$ (i.e. the queue length is less than $n$). Moreover, we also recover that as $n$ goes to infinity, this probability goes to 00 in the Super-Halfin-Whitt regime and goes to a value close to $1$ in the Sub-Halfin-Whitt regime. The full result is stated in Corollary 6.
In summary, our finite-time behavior characterizations match the steady-state and asymptotic results previously established in (Jhunjhunwala et al., 2023; Braverman et al., 2017) and provide a holistic insight for Erlang-C systems.
Regime Mean Field Sub-HW Halfin-Whitt Super-HW
$\alpha\approx 0$ $\alpha\in(0,1/2)$ $\alpha=1/2$ $\alpha\in(1/2,1)$ $\alpha\in[1,\infty]$
$\lambda_{n}/n\rightarrow c$ $\lambda_{n}=n-B\sqrt{n}$
$\mathbb{E}\left[\#ofwaitingjobs\right]$ $O(1)$ $O(\sqrt{n})$ $O(n^{\alpha})$
Asymp. behavior Gaussian (Jhunjhunwala et al., 2023; Braverman et al., 2017) Exponential (Jhunjhunwala et al., 2023; Braverman et al., 2017)
Mixing rate $L_{n}$ $D_{n}$ $1-o(1)$ $\frac{B^{2}}{4}$ $C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$
$e^{-\square t}$ $L_{n}\rightarrow 1-\sqrt{c}$ $D_{n}\rightarrow\frac{1}{25}$ $C_{n}\rightarrow 1$
This work (van Leeuwaarden and Knessl, 2011; Gamarnik and Goldberg, 2013) This work This work, (Zeifman, 1991)
Table 1. Overview of mixing time analysis for $M/M/n$ for $\lambda_{n}=n-n^{1-\alpha},\mu=1$ where $\alpha$ is a parameter. The number of waiting jobs is denoted as $\left[q-n\right]^{+}=\max\{q-n,0\}$, and the asymptotic behavior is the queue length distribution behavior as $n\rightarrow\infty$. When $\alpha\approx\infty$, this is called the Classical Heavy Traffic regime. For the Mean Field regime, we write $\alpha\approx 0$ as an intuitive explanation, whereas the actual condition is $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c\in[0,1)$.
1.2. Our approach: The Lyapunov-Poincaré method
It is well-known that one can obtain exponential ergodicity for a stochastic system if the system admits some Lyapunov drift outside of some finite set (Meyn and Tweedie, 1994; Down et al., 1995; Rosenthal, 1995; Lund et al., 1996; Baxendale, 2005; Douc et al., 2018; Meyn and Tweedie, 2012; Raginsky et al., 2017; Taghvaei and Mehta, 2022; Andrieu et al., 2022). In contrast, our goal is to get a handle on the mixing rate. While one can easily get a handle on the mixing outside of the said set using the negative drift, it is much harder to handle the behavior inside the finite set. The special case when the finite set is a singleton (which is the case in an $M/M/1$ queue) is well understood (Lund et al., 1996). When the finite set is not a singleton, one typically uses a ”stitching theorem” to combine knowledge of the drift outside the finite set and local mixing behavior inside the finite set to obtain a mixing time bound for the entire system. Prior work has explored the use of minorization (Down et al., 1995; Rosenthal, 1995; Baxendale, 2005; Taghvaei and Mehta, 2022), contraction (Durmus and Moulines, 2014; Hairer et al., 2011) within the finite set or obtain the local Poincaré constant by using the radius of the finite set (Dominique et al., 2008; Raginsky et al., 2017; Divol et al., 2024), and then a corresponding stitching theorem to obtain mixing rates. However, these approaches either do not lead to a tight mixing rate inside the finite set (especially for high-dimensional settings (Qin and Hobert, 2020)) or are hard to obtain in practice. To this end, we propose the Lyapunov-Poincaré method, where we combine the Lyapunov drift outside of a finite set with a local Poincaré inequality inside the finite set to establish mixing results. Our work extends the previously established Lyapunov drift framework (Meyn and Tweedie, 1994; Down et al., 1995; Rosenthal, 1995; Lund et al., 1996; Baxendale, 2005; Douc et al., 2018; Meyn and Tweedie, 2012; Raginsky et al., 2017; Taghvaei and Mehta, 2022; Andrieu et al., 2022) to allow a more fine-grained mixing analysis, which eventually enables us to obtain tight mixing bounds. In particular, we develop new stitching theorems so that depending on the properties of the finite set, we can better combine the drift properties outside the finite set with the local mixing behavior inside the finite set (which is characterized in terms of local Poincaré inequalities) so as to obtain more fine-grained mixing analysis.
Handling of the finite set: Our work takes novel approaches to obtain tight mixing bounds inside the finite set. To obtain tight mixing bounds inside the finite set, we propose two novel finite set techniques: 1) the canonical path method and 2) the truncation method. These two methods correspond to two different regimes, which allow us to obtain mixing time bounds that match the asymptotic behaviors. These methods can be summarized as follows:
  • Local canonical path method: The canonical path method, previously developed in (Jerrum and Sinclair, 1989), is a powerful technique used to analyze the mixing time of finite discrete Markov chains by bounding how ”congested” the state space is with respect to moving from one state to another. In particular, the canonical path method estimates the mixing time of a finite discrete Markov chain with the congestion ratio, which scales with the stationary distribution and the sum of the path length of paths passing through a particular edge. As such, the canonical path method can be viewed as a conductance-typed approach to obtain mixing. Notably, the canonical path method obtains a tight mixing bound when the Markov chain is a symmetric random walk on a path graph (Berestycki, 2020; Levin and Peres, 2017), which is also the behavior of the Erlang-C system in the finite set when an appropriate Lyapunov function is chosen. By applying the local canonical path method, we obtain a tight weighted local Poincare inequality of the Erlang-C system inside the finite set, which allows us to use our stitching theorems to obtain the mixing rate bound for the entire system. We formally state our local canonical path lemma in Lemma 1.
  • Truncation: On the other hand, the aforementioned local canonical path method is only viable when the distribution inside the finite set is roughly uniform. For other finite sets whose distributions are not necessarily uniform, we propose the truncation method to obtain local mixing results inside such sets. In essence, if we can find another system with known Poincaré constant, which behaves identically inside the finite set, then we can use it to get a handle on the local Poincaré constant of the original system. With this method, we obtain tight mixing time bounds up to a constant factor for $M/M/n$ queues in the Mean Field regime, i.e. when $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c\in(0,1)$. For $c=0$ (the so-called Light Traffic regime), we have the mixing rate approaches to $1$, which matches the limiting behavior of the system.
1.3. Related works
Unlike many mixing time works which focus on the discrete-time finite state space Markov chains (Jerrum and Sinclair, 1989; Lovasz and Simonovits, 1990; Berestycki, 2020; Levin and Peres, 2017; Wilson, 2004), analyzing the mixing behavior of queuing systems is a non-trivial endeavor as it involves handling an infinite, countable state space. Indeed, while there have been prior works on the mixing time of $M/M/n$ queues (Robert, 2003; Gamarnik and Goldberg, 2013; Zeifman, 1991; van Leeuwaarden and Knessl, 2011) and other queuing systems (Lorek and Szekli, 2015; Mao and Xia, 2015; Gurvich, 2014; Braverman, 2020, 2023; Luczak and McDiarmid, 2006; Rutten and Mukherjee, 2023), each of them can either only obtain tight finite-time bounds that match the asymptotic behavior only for a specific regime or setting, or can only show exponential ergodicity without an explicit rate (Gurvich, 2014; Braverman, 2020, 2023), geometric convergence to a ball (Luczak and McDiarmid, 2006) or subgeometric convergence (Rutten and Mukherjee, 2023). In particular, we shall provide a recap of mixing results for each of these regimes and settings.
$M/M/1$ queue: The paper (Morse, 1955) is the first to characterize the transient behavior of an $M/M/1$ queue. Following this work, (Van Doorn, 1985) uses the spectral representation of birth-and-death processes to obtain the rate of convergence. On the other hand, (Robert, 2003) uses a coupling argument to establish the distance to stationarity for the $M/M/1$ queue where the author upper bounds the distance to stationarity with the probability that the two processes will coalesce within time $T$, which can be further upper bounded by the probability of hitting 00 within time $T$. Unfortunately, both of these techniques are difficult to generalize to the $M/M/n$ queue given its more complicated dynamics.
$M/M/\infty$ queue: The explicit characterization of the transient distribution of the queue length is known for the $M/M/\infty$ queue. Such a result is obtained by considering a two-dimensional Poisson point process representing the arrival time and the service time. A different approach is to bound the Entropy of the system (Chafaï, 2006). In this work, the author uses a diffusion approximation of the $M/M/\infty$ queuing system and uses the binomial-Poisson entropic inequalities to establish the rate of convergence.
Now, we will look at the mixing results for specific $M/M/n$ regimes and queueing settings.
$M/M/n$ queue with $\alpha\in[1,\infty)$: In (Zeifman, 1991), the author uses the log-norm to characterize the convergence rate of the birth-and-death process, which applies to the $M/M/n$ setting. This method obtains the tight mixing rate of $(\sqrt{n\mu}-\sqrt{\lambda})^{2}$ for $\alpha\geq 1$ and a sub-optimal mixing rate of $\mu-\frac{\lambda}{n-1}\ll(\sqrt{n\mu}-\sqrt{\lambda})^{2}$ for $\alpha\in(1/2,1)$. The log-norm approach is also applied to analyze queueing models with queue-length-dependent admission control (Natvig, 1975), however, with a sub-optimal mixing rate (Van Doorn, 1981). This suggests that while simple, the log-norm approach is not strong enough to obtain sharp convergence rates for a wide variety of settings.
$M/M/n$ queue with $\alpha=1/2$: The papers (Gamarnik and Goldberg, 2013; van Leeuwaarden and Knessl, 2011) uses the spectral representation of birth-and-death processes, previously established by (Karlin and McGregor, 1957), to obtain the rate of convergence of the $M/M/n$ queue in the Halfin-Whitt regime, i.e. $\alpha=1/2$. Most notably, the authors were able to characterize the phase transition of the mixing rate. In particular, as $\frac{n\mu-\lambda}{\sqrt{n}}$ increases, the mixing rate transitions from $(\sqrt{n\mu}-\sqrt{\lambda})^{2}$ to $\approx 1$, establishing a phase shift from $M/M/1$-like behavior to $M/M/\infty$-like behavior. However, these results do not establish convergence in the Chi-square sense and are applicable only for $\alpha=1/2$.
$M/M/n/n+R$ queue: For a variant of $M/M/n$ with $n$ servers and $R$ waiting positions, (van Doorn and Zeifman, 2009; van Doorn, 2011) establish the rate of convergence to stationarity using the spectral representation method.
To the best of our knowledge, tight mixing results for $\alpha\in(0,1/2)$ and $\alpha\in(1/2,1)$ are still open prior to our work. And so, the overall mixing picture for $M/M/n$ is not complete. Our work fills the gap in the literature as we characterize tight mixing rates in the Super Halfin-Whitt regime and a tight mixing rate up to a constant for the Sub Halfin-Whitt regime. Moreover, (Gamarnik and Goldberg, 2013) is complementary to our work since this work focuses on the $\alpha=1/2$ regime while we obtain mixing results for every other regime, thus completing the mixing literature for $M/M/n$.
Jackson networks: For Jackson networks, (Lorek and Szekli, 2015) shows exponential ergodicity but their bound using Cheeger’s inequality is not tight in terms of traffic, and the dependence of the mixing time on the spectral gap of the communication graph is not discussed. The spectral gap of the open Jackson network is established in (Mao and Xia, 2015) using the Markov chain decomposition method.
Load-balancing systems: In addition to $M/M/n$ and Jackson networks, mixing results have been previously established for some load-balancing algorithms. The paper (Rutten and Mukherjee, 2023) obtained sub-geometric convergence for the spatial load-balancing system and (Ma and Maguluri, 2025) obtained $O(1/t)$ convergence rate for the join-the-shortest-queue system with $2$ servers. On the other hand, (Luczak and McDiarmid, 2006; Brightwell et al., 2018) obtained exponential convergence to a ball in the sub-Halfin-Whitt regime for the join-the-shortest-queue and power-of-$d$ policies.
1.3.1. Lyapunov drift method
It is well-known that an irreducible Markov chain is positive recurrent if and only if it has a negative drift outside of a bounded set (Foster, 1953), which is famously known as the Foster-Lyapunov Theorem. Since then, several works have used the drift assumption to establish the convergence rate of the Markov chain. A long line of literature (Douc et al., 2018; Meyn and Tweedie, 2012, 1994; Down et al., 1995; Rosenthal, 1995; Lund et al., 1996; Baxendale, 2005; Taghvaei and Mehta, 2022; Hairer et al., 2011; Durmus and Moulines, 2014) establish geometric ergodicity for general Markov chains whenever we have the Lyapunov drift condition outside of some set and some control inside of that set (i.e. the set is a singleton or the set admits a minorization or a contraction condition). Recently, (Anderson et al., 2020) proposes a drift and hitting time approach instead of using the minorization assumption, and in a similar spirit, (Cattiaux et al., 2010) uses the exponential tail assumption of the hitting time to obtain Poincaré inequalities. However, these approaches do not allow us to obtain tight convergence rates for our case.
The Lyapunov drift method is known to be very effective in establishing steady-state queueing results (Eryilmaz and Srikant, 2012; Hurtado-Lange and Maguluri, 2020; Jhunjhunwala et al., 2023) and beyond that, convergence and steady-state results in Stochastic Approximation and Optimization (Chen et al., 2021, 2019, 2020; Nguyen and Maguluri, 2023, 2024). As Lyapunov drift has been well-studied in the queuing literature, we naturally want to leverage this drift to prove mixing results. Previously, (Gurvich, 2014) obtained exponential convergence results for Markovian queues but the focus is not on obtaining tight mixing rates. For JSQ, (Braverman, 2020, 2023) uses the Lyapunov drift mixing framework by (Down et al., 1995). All of these works, however, only simply show exponential ergodicity and do not get a characterization on the mixing rate. This suggests that to get a tight mixing rate bound, a more fine-grained application of the Lyapunov drift method is required.
In addition to the negative drift approach and the convergence in $\ell_{2}$ distance or TV distance, some other works have opted for convergence in Wasserstein space instead (Qu et al., 2023; Qin and Hobert, 2021; Durmus et al., 2015, 2023) since many systems such as the constant step size SGD may not converge in TV distance but does converge in Wasserstein distance (Qu et al., 2023). In this metric, (Qu et al., 2023) proposes the notion of contractive drift rather than the Foster-Lyapunov drift to obtain the finite-time convergence of various systems such as the single server queue and the constant step size SGD. Notably, this notion allows us to bypass the bounded set problem, where the drift and minorization approach is known to have limitations (Qin and Hobert, 2020). On the other hand, the approach of using a contraction condition to handle the bounded set (also known as drift and contraction) often leads to a tradeoff in terms of the drift rate and the bounded set size (Qu et al., 2023). Instead of using this approach, our work builds on the Lyapunov drift framework to obtain Chi-square convergence, and we argue that our work is another approach to handling the bounded set issue that is typically faced when using drift arguments to obtain mixing results.
We summarize the literature and our results in Table 1. Our work covers the gap in the literature in the Light Traffic regime (where the arrival rate $\lambda_{n}$ is chosen such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=0$), the Sub-Halfin-Whitt regime (i.e. $\alpha\in(0,1/2)$), the Super-Halfin-Whitt regime (i.e. $\alpha\in(1/2,\infty)$) and the Heavy Traffic regime (i.e. $\alpha\in[1,\infty)$), completing the mixing literature for $M/M/n$. Additionally, note that our mixing results also match the asymptotic and the steady-state behavior of the system, that is the system behaves like $M/M/1$ for $\alpha\in(1/2,\infty)$ and $M/M/\infty$ for $\alpha\in(0,1/2)$.
1.4. Notations
To help the readers better understand our work, we define some notations used in our work as follows. For a countable set $\mathcal{S}$, denote the space of functions $f:\mathcal{S}\rightarrow\mathbb{R}$ that are square-integrable by $\ell_{2,\pi}$. Now, define the inner product
$\langle f,g\rangle_{\pi}:=\sum_{x\in\mathcal{S}}\pi(x)f(x)g(x)$
for $f,g\in\ell_{2,\pi}$. From the definition of the inner product, we have $\|f\|^{2}_{2,\pi}:=\langle f,f\rangle_{\pi}$. The distribution $\pi$ is said to be reversible for the operator $\mathcal{L}:\ell_{2,\pi}\rightarrow\ell_{2,\pi}$ if $\mathcal{L}$ is self-adjoint on $\ell_{2,\pi}$, i.e.,
$\langle f,\mathcal{L}g\rangle_{\pi}=\langle\mathcal{L}f,g\rangle_{\pi},\quad\forall f,g\in\ell_{2,\pi}.$
[…]
2. Model and Main Results
2.1. Model
Consider the $M/M/n$ queueing system, also known as the Erlang-C system, described as follows. Customers arrive at a central queue following a Poisson process with mean $\lambda$. There are $n$ homogeneous servers that serve the waiting customers in a first-come-first-serve (FCFS) manner. The service time of each customer is independently and exponentially distributed with mean $1/\mu$. Without the loss of generality, we assume $\mu=1$ unless specified otherwise. The number of customers $\{q(t):t\geq 0\}$ in the queue evolves as a continuous time Markov chain (CTMC) with state space $\mathbb{Z}_{\geq 0}$. This CTMC is a birth-and-death process with the birth rate $\lambda$ and the death rate $\min\{n,q\}$ in the state $q(t)=q$. We denote $\mathcal{L}$ as the infinitesimal generator of the CTMC and $P_{t}$ as the corresponding Markov semigroup where $P_{t}=e^{t\mathcal{L}}$. For a given $n$, we denote the stationary distribution of $M/M/n$ queue by $\nu_{n}$ and we omit the subscript of $n$, whenever it is clear from the context.
In addition, we define $M/M/\infty$ as the birth-and-death process with the birth rate $\lambda$ and the death rate $q$ in the state $q(t)=q$. This system behaves as if there are an infinite number of servers and whenever a customer arrives, it is immediately allocated an idle server to get served.
2.2. Chi-square distance
To measure the distance to stationarity, we use a notion of distance called the Chi-square distance, which is formally defined as follows:
Definition 1.
Given two distributions $P,Q$ whose supports are the subset of the countable set $\mathcal{S}$ and $P$ is absolutely continuous w.r.t $Q$ , the Chi-square distance of $P$ with respect to $Q$ is given by
(1) $\chi^{2}\left(P,Q\right)=\sum_{x\in\mathcal{S}}\frac{(P(x)-Q(x))^{2}}{Q(x)}.$
As $\mathbb{E}_{Q}\left[\frac{P}{Q}\right]=1$, the Chi-squared distance $\chi^{2}(P,Q)$ can also be rewritten as
(2) $\chi^{2}(P,Q)=\sum_{x\in\mathcal{S}}\frac{(P(x)-Q(x))^{2}}{Q(x)}=\sum_{x\in\mathcal{S}}Q(x)\left(\frac{P(x)^{2}}{Q(x)^{2}}-1\right)=\operatorname{Var}_{Q}\left(\frac{P}{Q}\right).$
The reason that we are interested in the Chi-square distance is that transient mean queue length and tail bounds can be derived from the Chi-square distance to stationarity. Indeed, we have the following Theorem on the variational form of the Chi-square distance.
Proposition 1.
Given two distributions $P,Q$ with countable supports and let $\mathcal{G}_{Q}$ be a collection of functions $g:\mathbb{R}\rightarrow\mathbb{R}$ such that $\mathbb{E}_{Q}\left[g(X)^{2}\right]<\infty\}$, we have
(3) $\chi^{2}(P,Q)=\sup_{g\in\mathcal{G}_{Q}}\frac{\left(\mathbb{E}_{P}[g(X)]-\mathbb{E}_{Q}[g(X)]\right)^{2}}{\operatorname{Var}_{Q}[g(X)]}$
The proof of Proposition 1 is well-known and is presented in Appendix A.3. Additionally, Proposition 7.15 in (Polyanskiy and Wu, 2024) states that the Chi-square distance can be used to give an upper bound on the TV distance as well. From Proposition 1, we can substitute the function $g$ of our interests to obtain specific bounds. In particular, we obtain the following bound on the distance of the moments of the distributions by substituting $g(x)=x^{k}$.
Corollary 0.
Given two distributions $P,Q$ with countable supports. In addition, let $k\in\mathbb{Z}^{+}$ and assume that $\mathbb{E}_{Q}[X^{2k}]<\infty$, then
(4) $\left|\mathbb{E}_{P}\left[X^{k}\right]-\mathbb{E}_{Q}\left[X^{k}\right]\right|\leq\chi(P,Q)\sqrt{\mathbb{E}_{Q}\left[X^{2k}\right]-\mathbb{E}_{Q}\left[X^{k}\right]^{2}}.$
We can substitute $g(X)=e^{\theta X}$ for an appropriately chosen $\theta\in\mathbb{R}$ to establish a bound on the difference of the Moment Generating Functions (MGF) as stated in the following corollary.
Corollary 0.
Given two distributions $P,Q$ with countable supports and $\theta\in\mathbb{R}^{+}$ such that $\mathbb{E}_{Q}[e^{2\theta X}]<\infty$, then
(5) $\left|\mathbb{E}_{P}\left[e^{\theta X}\right]-\mathbb{E}_{Q}\left[e^{\theta X}\right]\right|\leq\chi\left(P,Q\right)\sqrt{\mathbb{E}_{Q}\left[e^{2\theta X}\right]-\mathbb{E}_{Q}\left[e^{\theta X}\right]^{2}}.$
The MGF bounds are particularly useful for obtaining exponential tail bounds. With the proper scaling, many classical queueing systems such as $M/M/n$ and JSQ possess an MGF (Hurtado-Lange and Maguluri, 2020; Jhunjhunwala et al., 2023), and thus one can derive exponential tail bounds for such systems.
These results will be the bread and butter to obtain the finite-time statistics of our queueing systems of interest. Indeed, let $\pi_{n,t}$ be the queue length distribution at time $t$ and $\nu_{n}$ be the stationary distribution of the $M/M/n$ system, if we can obtain a bound on $\chi(\pi_{n,t},\nu_{n})$ then we can apply these results to obtain finite-time statistics such as mean queue length bounds and tail bounds. To this end, we develop machinery to establish the distance to stationarity bounds for $\pi_{n,t}$, in turn, resulting in finite-time statistics.
2.3. Main Results
Consider an $M/M/n$ queue with arrival rate $\lambda_{n}=n-n^{1-\alpha}$ for some $\alpha\in\mathbb{R}^{+}$ and service rate $\mu=1$ (unless stated otherwise). Such a parametrization of $\lambda$ is popularly known as the many-server-heavy-traffic regime and queueing systems are generally analyzed for $t,n\rightarrow\infty$. On the contrary, here we focus on a finite $t$ and a finite $n$. We obtain a bound on $\chi(\pi_{n,t},\nu_{n})$ of the form $e^{-\square t}$, thus, establishing the convergence of the finite-time queue length distribution $\pi_{n,t}$ to the stationary distribution $\nu_{n}$ as $t\rightarrow\infty$. Moreover, we provide a tight bound on the mixing rate, denoted by $\square$. This result is established in the following theorem.
Theorem 3.
Let $\pi_{n,t}$ be the queue length distribution at time $t$ of the continuous-time $M/M/n$ system with the arrival rate $\lambda_{n}=n-n^{1-\alpha}$ and service rate $1$ and let the stationary distribution be $\nu_{n}$. For $\alpha\geq 1$, we have that:
(6) $\chi(\pi_{n,t},\nu_{n})\leq e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0.$
For $\alpha\in(1/2,1)$ and $n\geq 65$, we have that:
(7) $\chi(\pi_{n,t},\nu_{n})\leq e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0,$
for some $C_{n}>0$ such that $\lim_{n\rightarrow\infty}C_{n}=1$.
For $\alpha\in(0,1/2)$ and a sufficiently large $n$ such that $\lambda_{n}\geq 3$, we have that:
(8) $\chi(\pi_{n,t},\nu_{n})\leq e^{-D_{n}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0,$
for some $D_{n}>0$ such that $\lim_{n\rightarrow\infty}D_{n}=1/25$.
Finally, for $\alpha=1/2$ and $n\geq 65$, we have
(9) $\chi(\pi_{n,t},\nu_{n})\leq e^{-\frac{t}{1387444804}}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0.$
The proof of this theorem and the exact expressions of $C_{n},D_{n}$ are described in Appendix B, and its proof sketch is summarized in Section 4. Note that we assume $n\geq 65$ for $\alpha\in[1/2,1)$ only to obtain an explicit mixing rate for $\alpha=1/2$ and to get reasonable constants in the proof. Obtaining an explicit bound for all $n\geq 1$ is possible at the cost of worse universal constants in our mixing rate bounds.
It is well-known that there is a phase transition at $\alpha=1/2$ (Halfin and Whitt, 1981), and the behavior in the regimes $\alpha\in(0,1/2)$ and $\alpha\in(1/2,\infty)$ is similar to $M/M/\infty$ and $M/M/1$ respectively. Our Theorem 3 captures such a phase transition, as evident by the mixing-time bounds for the $M/M/1$ and $M/M/\infty$ queues stated below:
Proposition 2.
Let $\pi_{t,1}$ be the queue length distribution at time $t\in\mathbb{R}^{+}$ of the continuous-time $M/M/1$ system with the arrival rate $\lambda$ and service rate $\mu$ such that $0<\lambda<\mu$ and let the stationary distribution be $\nu_{1}$, we have:
$\chi(\pi_{t,1},\nu_{1})\leq e^{-(\sqrt{\mu}-\sqrt{\lambda})^{2}t}\chi(\pi_{0,1},\nu_{1})\,\forall t\geq 0.$
Proposition 3.
Let $\pi_{t,\infty}$ be the queue length distribution at time $t$ of the continuous-time $M/M/\infty$ system with the arrival rate $\lambda\in\mathbb{Z}^{+}$ and the service rate $\mu$ and let the stationary distribution be $\nu_{\infty}$, we have:
$\chi(\pi_{t,\infty},\nu_{\infty})\leq e^{-\mu t}\chi(\pi_{0,\infty},\nu_{\infty})\,\forall t\geq 0.$
While the above results for $M/M/1$ and $M/M/\infty$ are previously established in (Chafaï, 2006; Morse, 1955; Robert, 2003), it is worth noting that our Lyapunov-Poincaré methodology provides a unified approach to obtain mixing results for $M/M/1$, $M/M/n$, and $M/M/\infty$ queues. The proof of Proposition 2, presented in Appendix B.1.4, is instructive to understand the case of $\alpha\in(1/2,\infty]$ of Theorem 3. Similarly, the proof of Proposition 3, while it is only applicable for $\lambda\in\mathbb{Z}^{+}$, in Appendix B.5 is helpful to understand the case of $\alpha\in(0,1/2)$. Observe that the mixing rate of the $M/M/n$ queue for $\alpha\in(1/2,\infty]$ and the $M/M/1$ queue is $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}\approx n^{1-2\alpha}\rightarrow 0$ as $n\rightarrow\infty$. On the other hand, the mixing rate of the $M/M/n$ queue for $\alpha\in(0,1/2)$ and the $M/M/\infty$ queue is $\Theta(1)$ as $n\rightarrow\infty$. Thus, our theorem characterizes the phase transition of the mixing rate from 00 to $\Theta(1)$ at $\alpha=1/2$. Finally, consider the case of $\alpha=1/2$, similar to the case of $\alpha\in(0,1/2)$, our mixing rate is $\Theta(1)$ but is off by a constant. For this case, a precise characterization of the mixing rate in the limit as $t\rightarrow\infty$, and $n\rightarrow\infty$ was provided in (Gamarnik and Goldberg, 2013), and obtaining a finite time behavior for a finite-sized system that exactly matches these limiting results is an interesting future direction.
While we establish a $\Theta(1)$ bound on the mixing rate for $\alpha\in(0,1/2)$, it is off by a constant compared to the mixing rate of $M/M/\infty$ queue, indicating that our mixing rate characterization is loose in this regime. Nonetheless, we believe our characterization of the mixing rate is state-of-the-art. For example, the result of (Zeifman, 1991) implies a mixing rate of $1-\frac{\lambda_{n}}{n-1}=O\left(n^{-\alpha}\right)\rightarrow 0$ as $n\rightarrow\infty$. Moreover, by imposing additional technical assumptions, we are able to improve our results to obtain tight mixing rates in this regime. To this end, we present two results assuming that either $\lambda_{n}$ is very small (light traffic, i.e. $\lambda_{n}/n\rightarrow 0$) or $\lambda_{n}\in\mathbb{Z}^{+}$, where the latter is an artifact of our methodology as discussed in Section 3.
Proposition 4.
Let $\pi_{n,t},\nu_{n}$ be the queue length distribution at time $t$ and the steady state distribution of the continuous-time $M/M/n$ system with unit service rate respectively and let $\lambda_{n}$ be a sequence of arrival rate such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c$ where $c\in[0,1)$. We have
$\chi(\pi_{n,t},\nu_{n})\leq e^{-L_{n}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0$
for some positive parameter $L_{n}$ such that $\lim_{n\rightarrow\infty}L_{n}=1-\sqrt{c}$.
Since $\lambda_{n}\in(0,n)$, there exists a unique positive number $\alpha_{n}$ such that $\lambda_{n}=n-n^{1-\alpha_{n}}$, and so the assumption $\lim_{n\rightarrow\infty}\lambda_{n}/n=c$ means that $\lim_{n\rightarrow\infty}\alpha_{n}\log n=-\log(1-c)$. When $c=0$, we have the mixing rate of the system matches the limiting behavior at the asymptotic (that is, mixing rate approaches $1$ as $n\rightarrow\infty$). On the other hand, for $c\in(0,1)$, Proposition 4 shows that the system admits a mixing rate that is bounded below by a constant that is independent of $n$. Moreover, when $c<\frac{576}{625}$, we get a better constant than the $\frac{1}{25}$ constant in the $\alpha\in(0,1/2)$ regime. One can find the proof of Proposition 4 in Appendix B.4. However, since we know that the phase transition happens at the Halfin-Whitt regime, this suggests that we could do better by obtaining an analysis that matches the limiting behavior for all $\alpha\in(0,1/2)$, possibly with some additional conditions. And so, we consider the case when the arrival rate is an integer, i.e. there is a sequence $\{\lambda_{k}\}_{k\geq\mathbb{Z}^{+}}$ such that $\lambda_{k}\in\mathbb{Z}\,\forall k\in\mathbb{Z}^{+}$. In this case, we can show that as $\lambda_{k}\rightarrow\infty$, we have the mixing rate approaches $1$ as follows.
Proposition 5.
Let $\pi_{n,t},\nu_{n}$ be the queue length distribution at time $t$ and the steady state distribution of the continuous-time $M/M/n$ system with unit service rate respectively and let $\{\lambda_{n}\}$ be a sequence of integer arrival rates ($\lambda_{n}\in\mathbb{Z}$) such that $\lambda_{n}\geq 0,n-\lambda_{n}\geq 1$ and
(10) $\lim_{n\rightarrow\infty}\frac{\log\left(n-\lambda_{n}\right)}{\log n}=1-\alpha$
where $\alpha<1/2$. Then, we have
(11) $\chi(\pi_{n,t},\nu_{n})\leq e^{-\overline{D}_{n}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0$
such that $\overline{D}_{n}>0$ and $\lim_{n\rightarrow\infty}\overline{D}_{n}=1$.
The proof of Proposition 5 is presented in Appendix B.3.3. With additional technical assumptions on $\lambda$, our mixing rate now matches with that of the $M/M/\infty$ queue asymptotically for $\alpha\in(0,1/2)$. Now, when both of these conditions are not satisfied, we can still show mixing but the mixing rate is only tight up to a constant, as shown in Theorem 3. Nevertheless, we believe that these two conditions are rather artificial and we conjecture that these conditions can be lifted to get a mixing rate approaching $1$, matching the mixing rate of the $M/M/\infty$ at the asymptotics.
2.3.1. Finite-time statistics
From the established mixing bounds in Theorem 3, we obtain finite-time statistics like the mean and tail. We start with the mean queue length below.
Corollary 0.
Let $\pi_{n,t}$ be the queue length distribution at time $t$ of the continuous-time $M/M/n$ system with the arrival rate $\lambda_{n}=n-n^{1-\alpha}$ and a service rate $1$ whose stationary distribution be $\nu_{n}$. For $\alpha\geq 1$, we have that
(12) $\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\sqrt{2}\left(n+n^{\alpha}\right)\chi(\pi_{n,0},\nu_{n}).$
For $\alpha\in(1/2,1)$, we have
(13) $\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\sqrt{2}\left(n+n^{\alpha}\right)\chi(\pi_{n,0},\nu_{n})$
for some $C_{n}>0$ such that $\lim_{n\rightarrow\infty}C_{n}=1$.
For $\alpha\in(0,1/2)$, we have
(14) $\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-D_{n}t}\sqrt{2}\left(n+n^{\alpha}\right)\chi(\pi_{n,0},\nu_{n})$
for some $D_{n}>0$ such that $\lim_{n\rightarrow\infty}D_{n}=1/25$.
And finally, if $\lambda_{n}$ is a sequence of arrival rates such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=0$, then we have
(15) $\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-L_{n}t}n\chi(\pi_{n,0},\nu_{n})$
for some $L_{n}>0$ such that $\lim_{n\rightarrow\infty}L_{n}=1$.
One can prove Corollary 4 by combining the Theorem 3 and Corollary 1 for $k=1$. We defer the details of the proof to Appendix C.1. Similar to the above corollary, we can obtain bounds on any moment $k\in\mathbb{Z}^{+}$ by Corollary 1 combined with a $2k$-moment bound on the steady-state distribution $\nu_{n}$. Moreover, one can go beyond the moments to also obtain tail bounds for the finite-time distribution which we present below.
Corollary 0.
Let $\pi_{n,t},\nu_{n}$ be the queue length distribution at time $t$ and the stationary distribution of the $M/M/n$ system with arrival rate $\lambda_{n}=n-n^{1-\alpha_{n}}$ and unit service rate respectively, and let $\varepsilon=1-\frac{\lambda_{n}}{n}$ and $q$ be the random variable denoting the queue length. For $n\geq n_{0},\delta\in(0,+\infty)$ and $n_{0}$ is a constant dependent on the regime, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]\leq\left(1+\chi(\pi_{n,t},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}$
In particular, for $\alpha_{n}=\alpha\in[1,\infty)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}.$
For $\alpha_{n}=\alpha\in(1/2,1)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}$
where $C_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}C_{n}=1$.
For $\alpha_{n}=\alpha\in(0,1/2)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-D_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{e^{-\frac{4n^{1-2\alpha}}{7(1+\delta)}}+\left(10+\frac{2}{\delta}\right)n^{2\alpha-1}}\times e^{-\frac{x}{2(1+\delta)}}$
where $D_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}D_{n}=1/25$.
And finally, if $\lambda_{n}$ is a sequence of arrival rates such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=0$ then we have
$\mathbb{P}_{\pi_{n,t}}\left[q-n>x\right]$ $\leq\left(1+e^{-L_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{e^{-\frac{n^{1-\alpha_{n}}}{2}}+4n^{\alpha_{n}-1}+\frac{8n^{2\alpha_{n}-1}}{e}}\times e^{-\frac{x}{2}}$
where $\lim_{n\rightarrow\infty}\alpha_{n}=0$ and $L_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}L_{n}=1$.
The proof of Corollary 5 is deferred to Appendix C.2. Note that if $q(t)\geq n$, then $q(t)-n$ is the number of waiting customers at time $t$. The results in Corollary 5 tell us what is the tail distribution of the number of waiting customers at time $t$. In the Sub-Halfin-Whitt regime, we have the pre-exponent term approaches 00 as $n\rightarrow\infty$, which suggests a faster decay rate than exponential. Moreover, we can also estimate the probability of having an idle server at time $t$ as follows.
Corollary 0.
Let $\pi_{n,t},\nu_{n}$ be the queue length distribution at time $t$ and the stationary distribution of the $M/M/n$ system with arrival rate $\lambda_{n}=n-n^{1-\alpha_{n}}$ and unit service rate respectively. Denote $r_{n}$ be a random variable denoting the number of idle servers, that is $r_{n}=\left[n-q\right]^{+}$ where $q$ is the queue length random variable. For $\alpha_{n}=\alpha\in(1/2,\infty)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[r_{n}>0\right]\leq 4e\pi n^{\frac{1}{2}-\alpha}+2\sqrt{e\pi}n^{\frac{1}{4}-\frac{\alpha}{2}}e^{-C_{n}(\sqrt{n}-\sqrt{\lambda})^{2}t}\chi(\pi_{n,0},\nu_{n})$
where $\lim_{n\rightarrow\infty}C_{n}=1$ for $\alpha\in(1/2,1)$ and $C_{n}=1$ for $\alpha\in[1,\infty)$.
For $\alpha_{n}=\alpha\in(0,1/2)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[r_{n}>0\right]\geq 1-\kappa n^{\alpha-\frac{1}{2}}e^{-n^{\frac{1}{2}-\alpha}}-e^{-D_{n}t}\chi\left(\pi_{n,0},\nu_{n}\right)$
where $\lim_{n\rightarrow\infty}D_{n}=1/25$.
The proof of Corollary 6 can be found in Appendix C.3. These results are the finite-time version of the steady-state results established in (Jhunjhunwala et al., 2023) and thus can be seen as the generalization of such results.
3. The Lyapunov-Poincaré method
To establish the mixing results stated in Subsection 2.3, we dedicate the following Section to introduce the so-called Lyapunov-Poincaré machinery. All the definitions and results in this section are valid for any CTMC with a countable state space $\mathcal{S}$, generator $\mathcal{L}$, and Markov semigroup $P_{t}$.
Assumption 1.
The given CTMC is irreducible, aperiodic, positive recurrent, and reversible.
Assumption 1 implies that there exists a unique stationary distribution $\nu$ such that $\nu(x)\mathcal{L}(x,y)=\nu(y)\mathcal{L}(y,x)$ for all $x,y\in\mathcal{S}$.
Definition 2.
A CTMC admits a Poincaré constant $C_{P}$ if for all test function $f\in\ell_{2,\nu}$, we have
(16) $\operatorname{Var}_{\nu}(f)\leq C_{P}\mathcal{E}(f,f).$
Here, the Dirichlet form $\mathcal{E}(f,f)$ is defined as:
(17) $\mathcal{E}(f,f)=\langle f,-\mathcal{L}f\rangle_{\nu}=-\sum_{x,y\in\mathcal{S}}\nu(x)\mathcal{L}(x,y)f(x)f(y).$
It is known that Poincaré inequalities immediately imply exponential ergodicity of the CTMC (Bakry et al., 2013). We formally state it in the following proposition for completeness.
Proposition 6.
Under Assumption 1, if the system admits the Poincaré inequality (16) with constant $C_{P}$, then
(18) $\chi^{2}\left(\pi_{t},\nu\right)\leq e^{-\frac{2t}{C_{P}}}\chi^{2}(\pi_{0},\nu)$
where $\pi_{t}$ is distribution at time $t$ and $\nu$ is the stationary distribution.
The proof of Proposition 6 is well-known and can be found in (Bakry et al., 2013) and in Appendix A.1. Proposition 6 shows that Poincaré inequality with constant $C_{P}$ implies exponential convergence in the chi-square distance with mixing rate $1/C_{P}$. Thus, for the rest of the section, our focus is on establishing the Poincaré constant.
Building on the recent developments on Markov chain mixing (Taghvaei and Mehta, 2022; Andrieu et al., 2022), we now present the Lyapunov-Poincaré methodology below, which has two main ingredients. The first one is a Foster-Lyapunov like assumption:
Assumption 2.
There exists a Lyapunov function $V:\mathcal{S}\rightarrow[1,\infty)$ along with $b:\mathcal{S}\rightarrow[0,\infty)$ and $\gamma>0$ such that
(19) $\mathcal{L}V(q)\leq-\gamma V(q)+b(q)\,\forall q\in\mathcal{S}.$
For the special case of $b=B1_{K}$ for some finite $K\subseteq S$ and $B>0$, the above is same as the classical Foster-Lyapunov condition. More generally, let $K=\{q:b(q)>0\}$, then in words, the above assumption ensures a negative drift proportional to $V(q)$ outside $K$. On the other hand, inside $K$, our drift might be positive but it is controlled by the $b(q)$ for $q\in K$. If $K$ is a singleton then we have a slightly modified assumption as follows.
Assumption 3.
There exists a single point $x^{*}\in\mathcal{S}$, a Lyapunov function $V:\mathcal{S}\rightarrow[0,\infty)$ satisfying $V(q)>0\,\forall q\neq x^{*}$ and constants $\gamma,B>0$ such that
(20) $\mathcal{L}V(q)\leq-\gamma V(q)+B1_{\{x^{*}\}}\,\forall q\in\mathcal{S}.$
Unlike Assumption 2, Assumption 3 allows us to have $V(x^{*})<1$. In many settings to be discussed below, this modification is necessary to obtain the tight mixing rate in the case we have a negative drift outside of a singleton. From here, we have the following proposition.
Proposition 7.
Assume that Assumptions 1 and 3 holds where $\{b(q)>0\}\subseteq\mathcal{S}$ is a singleton, then the CTMC admits the Poincaré constant $C_{P}=\frac{1}{\gamma}$.
While a similar convergence result for TV distance was previously established (Lund et al., 1996), this result allows us to establish Chi-square convergence, and we present the proof in Appendix A.2.3. Combining the above result with Proposition 6, one can obtain the Chi-square convergence as in (18). More generally, the finite set $K$ may not be a singleton for most of the queueing systems (including the $M/M/n$ queue), and thus, the Foster-Lyapunov assumption alone fails to provide a Poincaré inequality since we do not know how the system would mix inside $K$. So we need to establish a local mixing result inside $K$, which can be done using a local Poincaré inequality. Such a result is formalized in the following assumption.
Assumption 4.
(Weighted Poincaré) Given the stationary distribution $\nu$, a function $b:\mathcal{S}\rightarrow[0,\infty)$ such that $\sum_{q\in\mathcal{S}}b(q)\nu(q)$ is finite and let $\tau(q)=\frac{b(q)\nu(q)}{\sum_{q\in\mathcal{S}}b(q)\nu(q)}\forall q\in\mathcal{S}$ be a $(b,\nu)$-weighted distribution. Denote $K=\{q\in\mathcal{S}:b(q)>0\}$, a CTMC is said to admit a $b$-weighted Poincaré inequality if there exists a non-negative constant $C_{b}$ such that
(21) $\operatorname{Var}_{\tau}(f)=\left\lVert f-\mathbb{E}_{\tau}[f]\right\rVert_{2,\tau}^{2}\leq\frac{C_{b}}{\nu(K)}\langle f,-\mathcal{L}f\rangle_{\nu}\,\forall f\in\ell_{2,\nu}.$
Now, we combine Assumption 4 along with the drift assumption (Assumption 2) to obtain the Poincaré constant for the entire system, which is stated in the following Stitching Theorem.
Theorem 1.
(Stitching Theorem) Denote $K=\{q\in\mathcal{S}:b(q)>0\}$ and $\nu_{K}(q)=\frac{\nu(q)}{\nu(K)}\,\forall q\in K$. Under Assumptions 1, 2 and 4, the following inequality holds
(22) $\operatorname{Var}_{\nu}(f)\leq\frac{1+\left(\sum_{q\in\mathcal{S}}b(q)\nu_{K}(q)\right)C_{b}}{\gamma}\langle f,-\mathcal{L}f\rangle_{\nu}\,\forall f\in\ell_{2,\nu},$
where $C_{b}$ is the weighted local Poincaré constant in Assumption 4, $\nu$ is the stationary distribution of the CTMC and $b:\mathcal{S}\rightarrow[0,\infty)$ is the positive drift term in Assumption 2.
The Stitching Theorem (Theorem 1) is applicable for irreducible, aperiodic, positive recurrent and reversible CTMCs. For the purpose of analyzing queueing systems, we state the Stitching Theorem for the countable state space Markov chains, but this theorem can be easily generalized to the continuous state space (Taghvaei and Mehta, 2022) and can potentially be applied to other areas such as Learning and Sampling (Durmus and Moulines, 2014; Raginsky et al., 2017; Vempala and Wibisono, 2022). The proof of Theorem 1 can be found in Appendix A.2.1. For the special case of $b=B1_{K}$ for some finite $K\subseteq S$ and $B>0$, Assumption 4 gives the local Poincaré inequality
(23) $\operatorname{Var}_{\nu_{K}}(f)\leq\frac{C_{L}}{\nu(K)}\langle f,-\mathcal{L}f\rangle_{\nu},$
where $\nu_{K}(x)=\frac{\nu(x)}{\sum_{x\in K}\nu(x)}\forall x\in K$. And so, this allows us to recover the result of (Taghvaei and Mehta, 2022) as stated below.
Corollary 0.
Under Assumptions 1 and 2, assume that $b=B1_{K}$ where $B>0$ and $K\subseteq S$ is some finite set. Also, assume that the system satisfies Assumption 4 with $b$ and constant $C_{L}>0$. Then, the CTMC admits the Poincaré constant $\frac{1+BC_{L}}{\gamma}$.
A version of Corollary 2 is known in (Taghvaei and Mehta, 2022; Rosenthal, 1995) for reversible Markov chains and (Dominique et al., 2008; Raginsky et al., 2017) for Markov processes. The readers can find the proof of Corollary 2 in Appendix A.2.2, which is by applying the result of Theorem 1 for $b=B1_{K}$. The constant $C_{L}$ is called the local Poincaré constant w.r.t. the finite set $K$ of the CTMC. Intuitively, the CTMC mixes fast outside the finite set $K$ due to the negative drift, in addition to the rapid mixing within $K$ ensured by the “local” Poincaré inequality. In many applications, establishing a tight characterization of the local mixing behavior is the key to obtaining the tight mixing rate, and previously, the local Poincaré constant is predominantly obtained using the minorization condition (Rosenthal, 1995; Down et al., 1995; Taghvaei and Mehta, 2022). However, it is known that the approach using minorization condition often yields sub-optimal bounds, and the minorization condition itself is usually hard to obtain as well (Qin and Hobert, 2020; Anderson et al., 2020). Moreover, whenever the bounded set is not a singleton, there will be a ”stitching” error when combining the mixing behavior inside the finite set with the mixing behavior outside the finite set together. Therefore, our Stitching Theorem (Theorem 1) is a necessary generalization to achieve a tight characterization of the local mixing behavior inside the finite set, whereas a naive application of Corollary 2 would not yield such a result. For a more detailed explanation, we refer the readers to the discussion around Equation (29).
4. Proof sketch of mixing results
Recall that the $M/M/n$ queue experiences a phase transition at the Halfin-Whitt regime, i.e. $\alpha=\frac{1}{2}$. And so, to prove our main results, we will take two different approaches: a Lyapunov-Poincaré approach for the Super-Halfin-Whitt regime and a Lyapunov coupling approach for the Sub-Halfin-Whitt regime. The intuitions behind our approaches and the reasons why these regimes require two different approaches will become apparent as we go through the subsections below.
4.1. Regime 1: Super-Halfin-Whitt Regime
Let $K=\{q\in S:b(q)>0\}$, recall that the Lyapunov-Poincaré machinery requires the following:
  • Step 1: Obtain the correct Lyapunov drift outside of some finite set $K$, which ensures rapid mixing outside $K$.
  • Step 2: Obtain local mixing guarantees inside the finite set $K$ by exploiting the properties of $K$ and $b$.
In Step 1, we wish to choose an appropriate Lyapunov function to correctly capture the drift. To gain some intuition, let’s start with the $M/M/1$ queue. The $M/M/1$ queue has the negative drift $\lambda-\mu=-\varepsilon$ everywhere except when the queue is empty. And thus, to obtain exponential convergence, we choose the Lyapunov function $V(q)=e^{\theta q}$. Intuitively, the fluid analog of the $M/M/1$ queue would then satisfy $\dot{V}\approx-\varepsilon\theta V$ providing the negative drift as required. More precisely, we obtain
(24) $\mathcal{L}V\leq-\left(\sqrt{n}-\sqrt{\lambda}\right)^{2}V+b1_{\{0\}}$
where $b$ is some universal constant. The rate of $\left(\sqrt{n}-\sqrt{\lambda}\right)^{2}$ in the above equation matches previously established mixing results (Van Doorn, 1985; Robert, 2003) for the $M/M/1$ queue suggesting that $V$ is indeed the correct Lyapunov function. As seen in (24), we have a negative drift everywhere except the singleton $\{0\}$, so we simply apply Proposition 7 to obtain the Poincaré constant, which gives us the desired mixing rate.
For the $M/M/n$ queue, the dynamic is more complicated since there is another upward drift to $\lambda$ when $q<\lambda$ in addition to the downward drift when $q>\lambda$. So, we appropriately adjust our Lyapunov function and define it as follows:
(25) $V(q)=e^{\theta\left[q-(n-1)\right]^{+}+\theta\left[(n-1)-q\right]^{+}}$
with appropriately chosen $\theta\geq 0$ to accurately capture the negative drift above $\lambda$ and the positive drift below $\lambda$. While $\lambda$ is the center of the drift, we consider the distance of the queue length to $n-1$ in the exponent because of algebraic convenience in the proof. From this choice of Lyapunov function, we obtain the following drift lemma for the Super-Halfin-Whitt regime.
Lemma 0.
Let $V$ be defined as in (25) and $\mathcal{L}$ be the generator of the $M/M/n$ queue in the super Halfin-Whitt regime (i.e. $\alpha>1/2$) with arrival rate $\lambda_{n}=n-n^{1-\alpha}$. Then, there exists $\gamma_{n}\geq\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}$ and a function $b:\mathbb{Z}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ such that:
(26) $\mathcal{L}V(q)\leq-\gamma_{n}V(q)+b(q)1_{K}\,\forall q\in\mathbb{Z}_{\geq 0}$
where
(27) $K=\begin{cases}\{0,1,...,n-1\}&\text{ if }1\leq n\leq 7\\ \{\lceil 2\lambda_{n}\rceil-n,...,n-1\}&\text{ if }n>7.\end{cases}$
Furthermore, we have $b(q)\leq L\,\forall q\in K$ for some constant $L>0$ and $2\leq n\leq 7$ and for $n>7$, we have
(28) $b(q)=\begin{cases}2e^{2}n^{1-2\alpha}\,\,&\forall q\in K-\{n-1\}\\ 3n^{1-\alpha}&\text{if }q=n-1\\ 0\,\,&\forall q\not\in K.\end{cases}$
The above lemma completes Step 1 of our methodology by establishing a negative drift with rate $\gamma_{n}$ outside the finite set $K$. Unlike (Taghvaei and Mehta, 2022), the function $b$ in our drift condition (26) is not a constant function, which allows a more refined analysis and precipitates the use of our Stitching Theorem 1. To understand why the chosen Lyapunov function works for the $M/M/n$ queue, consider the following ODE that is a fluid analog of $M/M/n$ queue: $\dot{q}=\lambda_{n}-\min\{q,n\}$. When $q\geq\lambda_{n}+n^{1-\alpha}$ and $\theta_{1}\approx n^{-\alpha}$, we have $\dot{q}=\lambda_{n}-n=-n^{1-\alpha}$ which gives $\dot{V}\approx-n^{1-2\alpha}V\leq-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V$, which is the desired negative drift. Similarly, when $q\leq\lambda_{n}-n^{1-\alpha}$, we also have $\dot{V}\leq-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V$. This allows us to establish a negative drift with rate $-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ outside of a set $K$ with cardinality at most $\lceil 2n^{1-\alpha}-1\rceil$.
Observe that for $\alpha\geq 1$ or $n=1$, $K$ is a singleton as $2n^{1-\alpha}-1\leq 1$. In such a case, we simply apply Proposition 7 to obtain the desired mixing rate $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ with no constant factor loss. Otherwise, if $\alpha\in(1/2,1)$ then $K$ is no longer a singleton. And so, we come to Step 2, where we wish to show that the system mixes quickly in the finite set $K$ by showing that the system admits a local Poincaré inequality inside the finite set. To this end, we apply the local canonical path method (described in Appendix B.1.3) to establish the local Poincaré constant $C_{L}=\Theta(n^{1-2\alpha})$ in Lemma 2, which means that $C_{L}$ goes to zero as $n$ goes to infinity and allows us to show rapid mixing in $K$.
However, note that from (28), if we choose $B=\max_{q\in\mathbb{Z}_{\geq 0}}b(q)$ and then naively apply Theorem 1 in (Taghvaei and Mehta, 2022) (also Corollary 2 in our work), we would get a Poincaré constant
(29) $C_{P}=\frac{1+O\left(n^{2-3\alpha}\right)}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}$
which matches the limiting behavior as $n\rightarrow\infty$ only when $\alpha\in(2/3,\infty)$ rather than for the entire Super-Halfin-Whitt regime $\alpha\in(1/2,\infty)$. Thus, naively applying previously proposed ”stitching theorems” would not yield the desired bound. To address this gap, we need a more refined local mixing analysis, by treating the $b$ term carefully as follows.
Lemma 0.
(Weighted Poincaré for $M/M/n$ in the Super-Halfin-Whitt regime) Under Assumptions 1, 2, let $b$ be given by (28), $n\geq 65$ and $K=\{\lceil 2\lambda_{n}\rceil-n,...,n-1\}$. In addition, denote $\nu_{n}$ as the stationary distribution , we have Assumption 4 holds with constant $C_{b}=\Delta_{1}n^{2-4\alpha}+\Delta_{2}n^{3-6\alpha}$ holds for $\alpha\in[1/2,1)$ and for all test function $f\in\ell_{2,\nu_{n}}$ and for some positive constants $\Delta_{1},\Delta_{2}$.
Note that we assume $n\geq 65$ only to obtain reasonable universal constants in the weighted-Poincaré constant. Obtaining an explicit bound for all $n\geq 1$ is possible, but doing so would result in unreasonably large universal constants. One can further improve such universal constants by increasing the threshold depending on the applications. The proof and the precise characterization of the constant $C_{b}$ is reserved in Appendix 5.1, which is done by a clever application of the local canonical path method. In the final step, we combine the results in Lemma 1 and Lemma 2 and apply Theorem 1 to obtain the results in Equation (6) and Equation (7) in Theorem 3 which gives us a convergence bound for $\alpha\in(1/2,\infty)$ that matches the limiting behavior at the asymptotics. The full detail of the proof is deferred to Appendix B.1.4.
Remark: Previously, several works have attempted different methods to establish the local Poincaré inequality. Previous works (Taghvaei and Mehta, 2022; Rosenthal, 1995) rely on the minorization assumption to establish some form of local mixing inside $K$, i.e. a local Poincaré inequality as in (23). However, such an approach does not directly work in our setting since the assumption requires every state in the state space to be within reach of any other state, which clearly does not hold for a birth-and-death process. Such an assumption can be satisfied if one looks at an $m-$step transition matrix for a large enough $m>0$. However, such an approach usually results in a poor minorization constant. In addition, a few other works have obtained local Poincaré inequalities using the diameter of the finite set, such as (Raginsky et al., 2017; Divol et al., 2024), but generally, the local Poincaré constant bounds in these works are loose. Thus, as our goal is to obtain a tight characterization of the mixing rate, we take a different approach and propose a local version of the canonical path method that exploits the birth-and-death structure of the transition matrix. In fact, the non-local version has been used to bound the mixing time of random walks on different graph topologies (Berestycki, 2020).
The mixing rate $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ in the regime $\alpha\geq 1$ is tight and matches the result in (Zeifman, 1991). In addition, the mixing rate for the regime $\alpha\in(1/2,1)$ is asymptotically tight, i.e., it matches with $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ for large $n$. For $\alpha=1/2$, the Poincaré constant inside the finite set $K$ does not vanish. We show that it is bounded above by a constant implying that the mixing rate is bounded below by a universal constant.
4.2. Regime 2: Sub-Halfin-Whitt Regime
As we saw before in Proposition 3, the $M/M/n$ queue with $\alpha\in(0,1/2)$ behaves similarly to the $M/M/\infty$ queue. We use this connection to motivate our Lyapunov function. In the rest of the section, we first discuss the prove strategy for the $M/M/\infty$ queue, followed by that of the $M/M/n$ queue.
4.2.1. $M/M/\infty$
Consider the queue length $q$ of the $M/M/\infty$ queue with arrival rate $\lambda$ and service rate $\mu=1$. Unlike the $M/M/1$ queue, the dynamics of the diffusion analog of $(q-\lambda)/\sqrt{\lambda}$ is governed by the standard Ornstein-Uhlenbeck (OU) process. The Poincaré constant, and thus, the mixing rate of the OU process is equal to $1$ as a result of the Gaussian-Poincaré inequality (Bakry et al., 2013). One can show that $f(x)=x$ satisfy the drift equation (19) with exact equality for the OU-process with $\gamma=1$ and $b=0$, and so, it is a good choice for a Lyapunov function. One can also view $f(x)=x$ as the eigenfunction corresponding to the largest non-trivial eigenvalue $(=1)$ of the generator of the OU process, which is also the first non-constant Hermite polynomial. Hence, with the appropriate centering, we choose $V(q)=|q-\lambda|$ as our choice of the Lyapunov function to analyze the $M/M/\infty$ queue. In particular, analyzing the drift of $V(q)=|q-\lambda|$ allows us to prove Proposition 3. This result recovers the previous known mixing rate for the $M/M/\infty$ queue, see e.g.: (Chafaï, 2006). However, we present proof of this result using our Lypaunov-Poincaré methodology, which then motivates the proof for the $M/M/n$ queue in the sub-Halfin-Whitt regime.
4.2.2. $M/M/n$ in the Sub Halfin-Whitt regime
Observe that for an $M/M/n$ queue, the system behaves like an $M/M/\infty$ queue when $q<n$. Similarly, for $q>n$, the system behaves like an $M/M/1$ queue. And so, we stitch an exponential function for $q>n$ along with $|q-\lambda_{n}|$ for $q<n$ to obtain the following drift lemma.
Lemma 0.
For a sufficiently large $n$ such that $\lambda_{n}=n-n^{1-\alpha}\geq 3$, set $V(q)=\zeta e^{\theta(q-\lambda_{n})}\forall q>n$. If $\lambda_{n}\in\mathbb{Z}_{\geq 0}$, set $V(q)=|q-\lambda_{n}|\forall q\leq n$. Otherwise, set
(30) $V(q)=\begin{cases}|q-\lambda_{n}|\,&\forall q\leq n,q\not\in\{\lfloor\lambda_{n}\rfloor,\lceil\lambda_{n}\rceil\}\\ |\lfloor\lambda_{n}\rfloor-\lambda_{n}-1|&\textit{if }q=\lfloor\lambda_{n}\rfloor\\ |\lceil\lambda_{n}\rceil-\lambda_{n}+1|&\textit{if }q=\lceil\lambda_{n}\rceil\end{cases}$
where $\zeta$ is chosen such that $|n-\lambda_{n}|=\zeta e^{\theta(n-\lambda_{n})}$ and $\theta>0$ is a parameter to be chosen. Then, there exists $\gamma_{n},b$ with $\gamma_{n}\rightarrow 1$, such that $\mathcal{L}V\leq-\gamma_{n}V+b1_{K}$ where $b\leq 2\lambda_{n}$ and $K=\{\lambda_{n}\}$ if $\lambda_{n}\in\mathbb{Z}_{\geq 0}$ and $K=\{\lfloor\lambda_{n}\rfloor-1,\lfloor\lambda_{n}\rfloor,\lceil\lambda_{n}\rceil,\lceil\lambda_{n}\rceil+1\}$ otherwise.
The proof of this Lemma is deferred to Appendix B.3.1. Using the above lemma, the finite set is a singleton when $\lambda_{n}\in\mathbb{Z}_{\geq 0}$, so we directly use Proposition 7 to infer the Poincaré constant, which in turn gives us the mixing rate. Observe that as $\gamma_{n}\rightarrow 1$ in the above lemma, we asymptotically match the mixing rate of the $M/M/\infty$ queue.
However, when $\lambda_{n}\notin\mathbb{Z}_{\geq 0}$, the finite set ceases to be a singleton. Nonetheless, $|K|=4$ is independent of $n$, and so we obtain a $\Omega(1/n)$ local Poincaré constant using the canonical path method since there are $\Theta(n)$ transitions per time unit. Although this is not enough to overcome the stitching error to get a mixing rate that approaches $1$, combining this local Poincaré constant with our drift lemma above using the Stitching Theorem (Theorem 1) is sufficient to obtain an $\Omega(1)$ mixing rate. We defer the full proof to Appendix B.3.
4.3. Regime 3: Mean Field regime
In this subsection, we outline the proof of Proposition 4, which establishes a tight mixing bound for $\lambda_{n}\notin\mathbb{Z}_{\geq 0}$, but with an additional assumption that $\lambda_{n}\ll n$. As the result in the previous subsection did not yield a tight mixing rate for $\lambda_{n}\notin\mathbb{Z}_{\geq 0}$, we propose the following alternative Lyapunov function: $V(q)=e^{\theta\left[q-n\right]^{+}}$ with $\theta\approx\varepsilon$. This choice does not align with the intuition to engineer $V(q)\approx|q-\lambda_{n}|$ for $q<n$ to mimic the $M/M/\infty$ behavior. Indeed, analyzing the drift of $V(q)=e^{\theta\left[q-n\right]^{+}}$ results in a finite set $K=\{0,1,\ldots,n\}$, which is much larger than what we previously had. However, $V(q)$ allows us to exploit the strong drift for $q>n$. Indeed, we establish the drift inequality (19) with $\gamma=\Theta(n^{1-2\alpha})\rightarrow\infty$ as $n\rightarrow\infty$. We then exploit the connection to the $M/M/\infty$ queue to obtain a local Poincaré inequality corresponding to the set $K=\{0,1,\ldots,n\}$ using the following lemma:
Lemma 0.
Let $K=\{m,...,M\}$ be a connected subset of the state space $\mathcal{S}\subset\mathbb{Z}$ of the birth-and-death process, $\nu$ be the stationary distribution of the CTMC and $\nu_{K}(x)=\nu(x)/\sum_{x\in K}\nu(x)\,\forall x\in\mathcal{S}$. Furthermore, let $\mathcal{L}$ be the generator of the birth-and-death process and assume that $\operatorname{Var}_{\nu}(f)\leq C_{P}\langle f,-\mathcal{L}f\rangle\nu\,\forall f\in\ell_{2,\nu}$, we have
(31) $\operatorname{Var}_{\nu_{K}}(f)\leq C_{P}\sum_{x:x,x+1\in K}Q_{K}(x,x+1)\left(f(x)-f(x+1)\right)^{2}\,\forall f\in\ell_{2,\nu_{K}}$
where $Q_{K}(x,y)=\nu_{K}(x)\mathcal{L}(x,y)\,\forall x,y\in\mathcal{S}$.
In words, the above lemma allows us to obtain local mixing results if we know the mixing behavior of another system that subsumes the dynamic inside the finite set. The $M/M/\infty$ queue indeed subsumes the dynamics inside the finite set $K\in\{0,1,\ldots,n\}$. The above lemma then allows us to establish the local Poincaré constant of $\approx 1$. We call this approach the “truncation argument” to obtain a local Poincaré inequality. We then apply our Stitching Theorem (Theorem 1) and use the fact $\lambda_{n}\ll n$ to conclude that the Poincaré constant for the $M/M/n$ queue is $\approx 1$ which completes the proof. The proof details of Proposition 4 and Lemma 4 are deferred to Appendix B.4 and B.4.2 respectively.
5. Proof Details for the Super-Halfin-Whitt regime
5.1. Proof of Lemma 2
In this subsection, we will establish the $b$-weighted Poincaré results where $b$ is the positive drift term in Lemma 1. To this end, we first need to develop a toolkit to analyze the local mixing property of the birth-and-death process in a bounded state space.
Lemma 0.
(Local canonical path method) For any birth-and-death process with the generator $\mathcal{L}$, a subset $K$ of the state space $\mathbb{Z}_{\geq 0}$ and the stationary distribution $\nu$, we have the following inequality holds:
(32) $\operatorname{Var}_{\nu_{K}}(f)\leq C_{L}\sum_{k:k,k+1\in K}Q_{K}(k,k+1)(f(k+1)-f(k))^{2}$
where the constant $C_{L}$ is defined as:
(33) $C_{L}=\max_{k:k,k+1\in K}\frac{\sum_{x,y\in K:x\leq k\leq y-1}|x-y|\nu_{K}(x)\nu_{K}(y)}{Q_{K}(k,k+1)}\,\forall f\in\ell_{2,\nu_{K}}$
where $\nu_{K}(x)=\frac{\nu(x)}{\sum_{x\in K}\nu(x)}\,\forall x\in K$ and $\nu_{K}(x)=0$ otherwise and $Q_{K}(u,v)=\nu_{K}(u)\mathcal{L}(u,v)=\nu_{K}(v)\mathcal{L}(v,u)\forall u,v\in\mathbb{Z}_{\geq 0}$.
Here, $C_{L}$ can also be understood as the congestion ratio, which is lower bounded by the inverse mixing rate (Levin and Peres, 2017). When $\nu_{K}$ is roughly uniform and the transitions are roughly symmetric, the constant $C_{L}$ is proportion to the sum of the path length of the path passes through an edge $e=(k,k+1)$ that maximizes the constant. As the name suggests, we prove this lemma by analyzing the path structure as follows.
Proof.
We emulate the proof for the canonical path method as in (Berestycki, 2020) to handle the continuous-time Markov chain. Denote $\gamma_{x,y}$ as one of the paths connecting $x,y$, note that $\mathcal{L}(x,y)>0$ if and only if $|x-y|=1$, thus all path in $K$ is a sequence of adjacent states. For any test function $f\in\ell_{2,\nu_{K}}$ and let $Q_{K}(x,y)=\nu_{K}(x)\mathcal{L}(x,y)\forall x\neq y$, we have:
$\operatorname{Var}_{\nu_{K}}(f)$ $=\sum_{x<y\in K}\nu_{K}(x)\nu_{K}(y)\left(f(x)-f(y)\right)^{2}$
$\overset{(a)}{\leq}\sum_{x<y\in K}|x-y|\nu_{K}(x)\nu_{K}(y)\left(\sum_{x\leq k\leq y-1}(f(k)-f(k+1))^{2}\right)$
$=\sum_{x<y\in K}|x-y|\nu_{K}(x)\nu_{K}(y)\left(\sum_{x\leq k\leq y-1}\frac{(f(k)-f(k+1))^{2}Q_{K}(k,k+1)}{Q_{K}(k,k+1)}\right)$
$\overset{(b)}{\leq}\sum_{k:k,k+1\in K}\left(\frac{\sum_{x,y\in K:x\leq k\leq y-1}|x-y|\nu_{K}(x)\nu_{K}(y)}{Q_{K}(k,k+1)}\right)\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1).$
Here, (a) follows from Cauchy-Schwarz inequality and (b) follows from the interchange of sums. Hence proved. ∎
In the next step, we show that $\nu_{K}$ is a roughly uniform distribution whose support is the set $K$. This lemma lays the groundwork for us to apply the local canonical path method.
Lemma 0.
Let $\lambda_{n}=n-n^{1-\alpha},K=\{\lceil 2\lambda_{n}\rceil-n,...,n-1\}$ and $\nu_{K}(x)=\frac{\nu(x)}{\sum_{x\in K}\nu(x)}\forall x\in K$ and $\nu_{K}(x)=0$ otherwise. For $\alpha\in(1/2,1)$ and $n\geq 65$, we have
(34) $\frac{13\sqrt{3}e^{\frac{-1}{12}}n^{\alpha-1}}{320}\leq\nu_{K}(x)\leq\frac{160e^{\frac{1}{12}}}{13\sqrt{3}}n^{\alpha-1}\,\forall x\in K.$
The proof of the above lemma is deferred to Appendix B.1.2. With these results in hand, we are now ready to prove Lemma 2.
Proof of Lemma 2.
Let $\tau$ be a measure such that $\tau(q)=b(q)\nu_{K}(q)\,\forall q\in\mathbb{Z}_{\geq 0}$, where $b$ is defined as in (28). Note that from Lemma 1 and Lemma 2:
(35) $\tau(q)$ $\leq\frac{480\sqrt{3}}{13}e^{\frac{1}{12}}\text{ for }q=n-1,$
(36) $\tau(q)$ $\leq\frac{320e^{2}}{13\sqrt{3}}e^{\frac{1}{12}}n^{-\alpha}\text{ for }q\in K-\{n-1\}.$
In addition, we also have
(37) $\tau(q)$ $\geq 3n^{1-\alpha}\frac{13}{80}\frac{e^{\frac{-1}{12}}\sqrt{3}n^{\alpha-1}}{4}=\frac{39\sqrt{3}e^{\frac{-1}{12}}}{320}\text{ for }q=n-1,$
(38) $\tau(q)$ $\geq 2e^{2}n^{1-2\alpha}\times\frac{13}{80}\frac{e^{\frac{-1}{12}}\sqrt{3}n^{\alpha-1}}{4}=\frac{13\sqrt{3}e^{\frac{23}{12}}}{160}n^{-\alpha}\text{ for }q\in K-\{n-1\}.$
From Lemma 1 and Lemma 2, we have
$\operatorname{Var}_{\tau}(f)$ $=\frac{1}{2}\sum_{x,y\in K}\tau(x)\tau(y)(f(x)-f(y))^{2}$
$=\underbrace{\sum_{x\in K}\tau(n-1)\tau(x)\left(f(n-1)-f(x)\right)^{2}}_{T_{1}}+\underbrace{\frac{1}{2}\sum_{x,y\in K-\{n-1\}}\tau(x)\tau(y)(f(x)-f(y))^{2}}_{T_{2}}.$
The terms $T_{1},T_{2}$ can be bounded by the following claims.
Claim 1.
$T_{1}\leq 560229\times n^{2-4\alpha}\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1)$.
Next, using the local canonical path method, we have the following claim:
Claim 2.
$T_{2}\leq 459951\times n^{3-6\alpha}\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1).$
Finally, we show that the total mass of $\tau$ is no more than a constant factor of the total mass of $\nu_{K}$.
Claim 3.
$\sum_{q\in K}\tau(q)\leq 340$ for $K=\{\lceil 2\lambda_{n}\rceil-n,...,n-1\}$.
The proofs of these claims are deferred to Appendix B.1.3. Denote $\Delta_{1}=560229,\Delta_{2}=459951$, combining the claims gives us for all $f\in\ell_{2,\nu_{n}}$ that
$\operatorname{Var}_{\tau}(f)=T_{1}+T_{2}\leq\left(\Delta_{1}n^{2-4\alpha}+\Delta_{2}n^{3-6\alpha}\right)\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1)$
$=\frac{\Delta_{1}n^{2-4\alpha}+\Delta_{2}n^{3-6\alpha}}{\nu_{n}(K)}\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}\nu_{n}(k)\mathcal{L}(k,k+1)$
$\leq\frac{\Delta_{1}n^{2-4\alpha}+\Delta_{2}n^{3-6\alpha}}{\nu_{n}(K)}\sum_{k=0}^{\infty}\left(f(k)-f(k+1)\right)^{2}\nu_{n}(k)\mathcal{L}(k,k+1)=\frac{\Delta_{1}n^{2-4\alpha}+\Delta_{2}n^{3-6\alpha}}{\nu_{n}(K)}\langle f,-\mathcal{L}f\rangle_{\nu_{n}}.$
Hence proved. ∎
6. Conclusion and Future Work
In this paper, we establish bounds on the Chi-squared distance between the finite-time and the steady-state queue length distribution for the $M/M/n$ queue. We demonstrate that the mixing rate is tight (possibly up to a constant) in the many-server-heavy-traffic regimes. We also obtain bounds on the mean, moments, and tail of the queue length for a finite time and finite $n$. To prove these results, we build a Lyapunov-Poincaré framework and also propose two different ways to obtain the local Poincaré inequality which are of independent interest. In particular, our proposed local canonical path method and the truncation method have the potential to provide stronger bounds than the classical drift and minorization method (Down et al., 1995; Rosenthal, 1995).
While we focus on analyzing a birth-and-death chain in this paper, note that our Lyapunov - Poincaré methodology is flexible enough to be applied to any reversible CTMC. Generalizing our results to other reversible Markov chains would pose the main challenge of constructing a suitable Lyapunov function resulting in suitable negative drift outside a suitable finite set. Nonetheless, the results in Section 3 are versatile, and one can hope to employ this methodology to obtain mixing results beyond birth-and-death chains. Beyond queueing theory, our finite set approaches can be further applied to other areas such as Learning, Generative AI and Sampling (Raginsky et al., 2017; Vempala and Wibisono, 2022; Durmus and Moulines, 2014; Chervonyi et al., 2025; Chen et al., 2024).
Acknowledgements.
This work was partially supported by NSF grants EPCN-2144316 and CMMI-2140534 and Georgia Tech ARC-ACO Fellowship. The authors thank Professor Srikant Rayadurgam and Professor Debankur Mukherjee for suggesting relevant references for the project. The authors also kindly thank the anonymous reviewers for their helpful reviews and feedback.
Appendix A Proofs of foundational results
In the following Section, we will provide the missing proofs for the key results in the main text.
A.1. Proof of Proposition 6
Proposition 0.
Under Assumption 1 and if the system admits the Poincaré inequality (16) with constant $C_{P}$ then
(39) $\chi^{2}\left(\pi_{t},\nu\right)\leq e^{-\frac{2t}{C_{P}}}\chi^{2}(\pi_{0},\nu)$
where $\pi_{t}$ is distribution at time $t$ and $\nu$ is the stationary distribution.
Proof.
Let $\pi_{t}=\pi_{0}P_{t}$ and note that $\chi^{2}(\pi_{t},\nu)=\operatorname{Var}_{\nu}\left(\frac{\pi_{t}}{\nu}\right)=\operatorname{Var}_{\nu}\left(\frac{\pi_{0}P_{t}}{\nu}\right)$. Then, we have
$\frac{d}{dt}\chi^{2}(\pi_{t},\nu)=\frac{d}{dt}\operatorname{Var}_{\nu}\left(\frac{\pi_{0}P_{t}}{\nu}\right)=\frac{d}{dt}\left\lVert\frac{\pi_{0}P_{t}}{\nu}-1\right\rVert_{2,\nu}^{2}=2\left\langle\frac{d}{dt}\frac{\pi_{0}P_{t}}{\nu},\frac{\pi_{0}P_{t}}{\nu}-1\right\rangle_{\nu}.$
Since we have $\frac{d}{dt}\frac{\pi_{0}P_{t}}{\nu}=\lim_{\delta\rightarrow 0}\frac{\pi_{0}}{\nu}\left(\frac{P_{t+\delta}-P_{t}}{\delta}\right)=\frac{\pi_{0}P_{t}}{\nu}\lim_{\delta\rightarrow 0}\left(\frac{P_{\delta}-1}{\delta}\right)=\frac{\pi_{0}P_{t}\mathcal{L}}{\nu}$, this gives us
$\frac{d}{dt}\chi^{2}(\pi_{t},\nu)$ $=2\left\langle\frac{\pi_{0}P_{t}\mathcal{L}}{\nu},\frac{\pi_{0}P_{t}}{\nu}\right\rangle_{\nu}=2\left\langle\frac{\pi_{0}P_{t}}{\nu},\mathcal{L}\frac{\pi_{0}P_{t}}{\nu}\right\rangle_{\nu}\text{ from reversibility via the adjoint property,}$,
$=-2\mathcal{E}\left(\frac{\pi_{0}P_{t}}{\nu},\frac{\pi_{0}P_{t}}{\nu}\right)$
[…]
$=-\frac{2}{C_{P}}\chi^{2}(\pi_{t},\nu).$
where the inequality comes from the Poincaré inequality (16). Using Gronwall’s inequality, we have
$\chi^{2}\left(\pi_{t},\nu\right)\leq e^{-\frac{2t}{C_{P}}}\chi^{2}(\pi_{0},\nu).$
Hence proved.
Remark on the convergence result: To the best of our knowledge, we believe that we are the first to establish a Chi-square convergence result for countable state-space CTMCs. Our approach to handle the Markov semigroup and the Poincaré inequality follows similarly from the arguments in (van Handel, 2016; Bakry et al., 2013, 2008). In addition, (Chewi et al., 2020) previously established Chi-square convergence from the Poincaré inequality in the context of Langevin dynamics. We combine these ideas and reestablish the convergence result in the context of CTMCs with a countable state space.
Comparison with TV distance convergence: Aside from Chi-square convergence, we also note that several works obtained TV distance convergence for queueing systems and countable state space Markov chains (Robert, 2003; Baxendale, 2005; Luczak and McDiarmid, 2006; Brightwell et al., 2018). Most notably, the load balancing results (Luczak and McDiarmid, 2006; Brightwell et al., 2018) does not give convergence to 00 as $t\rightarrow\infty$ as we have to take into account potential bad initial states. To handle these bad initial states, one has to either analyze the system conditioning on the fact that the system stays in the good states (as in (Luczak and McDiarmid, 2006; Brightwell et al., 2018)) or somehow incorporate the bad initial states in the initial distance. For the latter, this would mean that the pre-exponent term will be very bad since the TV distance is upper bounded by $1$. On the other hand, our Chi-square convergence result can capture all possible initial states using the initial Chi-square distance, which allows us to obtain convergence as $t\rightarrow\infty$ even for very bad initial states.
A.2. Proof of Proposition 7 and Theorem 1
To help explain the intuitions of the proof, we will first present the proof for Theorem 1 and then prove its singleton variant.
A.2.1. Proof of Theorem 1
Lemma 0.
For any continuous-time reversible Markov chain with the state space $\mathcal{S}$, the generator $\mathcal{L}$ and the stationary distribution $\pi$, the following inequality holds for any test function $f\in\ell_{2,\pi},m\in\mathbb{R}$ and the Lyapunov function $V$ such that $V(q)>0\,\forall q\in\mathcal{S}$:
(40) $\langle(f-m\mathbf{1})^{2}/V,-\mathcal{L}V\rangle_{\pi}$ $\leq\langle f,-\mathcal{L}f\rangle_{\pi}$
Proof.
We will replicate the proof in (Taghvaei and Mehta, 2022) to obtain a result for the continuous-time Markov chain. Denote $g=f-m\mathbf{1}$ where $\mathbf{1}$ is the vector of $1$s (see Subsection 1.4). Since it is known that $\mathcal{L}\mathbf{1}=0$, observe that
(41) $\langle f,\mathcal{L}f\rangle_{\pi}=\langle f,\mathcal{L}f\rangle_{\pi}-m\underbrace{\langle\mathcal{L}\mathbf{1},f\rangle_{\pi}}_{=0}\overset{(a)}{=}\langle f,\mathcal{L}f\rangle_{\pi}-m\langle\mathbf{1},\mathcal{L}f\rangle_{\pi}=\langle f-m\mathbf{1},\mathcal{L}f\rangle_{\pi}=\langle g,\mathcal{L}f\rangle_{\pi}$
where (a) follows from the self-adjoint property. Similarly, we also have
(42) $\langle g,\mathcal{L}f\rangle_{\pi}=\langle g,\mathcal{L}(f-m\mathbf{1}+m\mathbf{1})\rangle_{\pi}=\langle g,\mathcal{L}g+m\underbrace{\mathcal{L}\mathbf{1}}_{=0}\rangle_{\pi}=\langle g,\mathcal{L}g\rangle_{\pi}.$
And so, from (41) and (42), we have
(43) $\langle f,-\mathcal{L}f\rangle_{\pi}\leq\langle g,-\mathcal{L}g\rangle_{\pi}.$
Which means that showing (40) is equivalent to showing:
(44) $\left\langle\frac{g^{2}}{V},-\mathcal{L}V\right\rangle_{\pi}$ $=\langle g,-\mathcal{L}g\rangle_{\pi}.$
Note that $\mathcal{L}(x,y)\geq 0$ whenever $x\neq y$, so we have
00 0 $\leq\sum_{x\in\mathcal{S}}\sum_{y\in\mathcal{S}}V(x)V(y)\left(\frac{g(y)}{V(y)}-\frac{g(x)}{V(x)}\right)^{2}\pi(x)\mathcal{L}(x,y)$
$=\sum_{x\in\mathcal{S}}\sum_{y\in\mathcal{S}}V(x)V(y)\left(\frac{g(x)^{2}}{V(x)^{2}}+\frac{g(y)^{2}}{V(y)^{2}}-\frac{2g(x)g(y)}{V(x)V(y)}\right)\pi(x)\mathcal{L}(x,y)$
$=\left\langle\frac{g^{2}}{V},\mathcal{L}V\right\rangle_{\pi}+\left\langle V,\mathcal{L}\frac{g^{2}}{V}\right\rangle_{\pi}-2\langle g,\mathcal{L}g\rangle_{\pi}.$
$\Leftrightarrow$ $\left\langle\frac{g^{2}}{V},-\mathcal{L}V\right\rangle_{\pi}+\left\langle V,-\mathcal{L}\frac{g^{2}}{V}\right\rangle_{\pi}\leq 2\langle g,-\mathcal{L}g\rangle_{\pi}$
Observe that the generator $\mathcal{L}$ possesses the self-adjoint property from the reversibility of the system, it follows that $\langle\frac{g^{2}}{V},-\mathcal{L}V\rangle_{\pi}=\langle V,-\mathcal{L}\frac{g^{2}}{V}\rangle_{\pi}$. This gives (44). And since we have $g=f-m\mathbf{1}$ and (43), we have that
$\langle(f-m\mathbf{1})^{2}/V,-\mathcal{L}V\rangle_{\pi}$ $\leq\langle f,-\mathcal{L}f\rangle_{\pi}.$
Hence proved. ∎
Theorem 0 (Restatement of Theorem 1).
Denote $K=\{q\in\mathcal{S}:b(q)>0\}$ and $\nu_{K}(q)=\frac{\nu(q)}{\nu(K)}\,\forall q\in K$. Under Assumptions 1, 2 and 4, the following inequality holds
(45) $\operatorname{Var}_{\nu}(f)\leq\frac{1+\left(\sum_{q\in\mathcal{S}}b(q)\nu_{K}(q)\right)C_{b}}{\gamma}\langle f,-\mathcal{L}f\rangle_{\nu}\,\forall f\in\ell_{2,\nu},$
where $C_{b}$ is the weighted local Poincaré constant in Assumption 4, $\nu$ is the stationary distribution of the CTMC and $b:\mathcal{S}\rightarrow[0,\infty)$ is the positive drift term in Assumption 2.
Proof.
It is well-known that $\left\lVert f-\mathbb{E}_{\nu}[f]\mathbf{1}\right\rVert_{2,\nu}\leq\left\lVert f-m\mathbf{1}\right\rVert_{2,\nu}$ for all constants $m\in\mathbb{R}$. Therefore, in order to prove the Poincaré inequality, it suffices to show that:
(46) $\left\lVert f-m\mathbf{1}\right\rVert^{2}_{2,\nu}\leq C_{P}\langle f,-\mathcal{L}f\rangle_{\nu},\quad\forall f\in\ell_{2,\nu},$
for the designated $C_{P}$ and for some constant $m$ to be chosen later. Consider the general drift condition (19)
$\mathcal{L}V(q)\leq-\gamma V(q)+b(q)\,\forall q\in\mathcal{S}.$
Multiply both sides by $\frac{(f(q)-m)^{2}}{V(q)}$ for some $f\in\ell_{2,\nu}$ to obtain
$\frac{(f(q)-m)^{2}}{V(q)}\mathcal{L}V(q)$ $\leq-\gamma(f(q)-m)^{2}+\frac{b(q)}{V(q)}(f(q)-m)^{2}$
$\leq-\gamma(f(q)-m)^{2}+b(q)(f(q)-m)^{2}\text{ since }V(q)\geq 1,$
where the second inequality follows the fact that we have assumed $V(q)\geq 1$. Rearranging the terms to
$\gamma(f(q)-m)^{2}\leq-\frac{(f(q)-m)^{2}}{V(q)}\mathcal{L}V(q)+b(q)(f(q)-m)^{2},$
multiplying both sides by $\nu(q)$ and summing with respect to $q\in S$, we have:
(47) $\gamma\left\lVert f-m\mathbf{1}\right\rVert^{2}_{2,\nu}\leq\langle\left(f-m\mathbf{1}\right)^{2}/V,-\mathcal{L}V\rangle_{\nu}+\sum_{q\in\mathcal{S}}b(q)\nu(q)(f(q)-m)^{2}.$
Denote $K=\{q:b(q)>0\}$ and $\tau(q)=\frac{b(q)\nu(q)}{\sum_{q\in K}b(q)\nu(q)}=\frac{b(q)\nu_{K}(q)}{\sum_{q\in K}b(q)\nu_{K}(q)}$ where $\nu_{K}(q)=\frac{\nu(q)}{\sum_{x\in K}\nu(x)}$. From Assumption 4 and choose $m=\frac{\sum_{q\in K}b(q)\nu_{K}(q)f(q)}{\sum_{q\in K}b(q)\nu_{K}(q)}$, we have:
(48) $\operatorname{Var}_{\tau}(f)$ $\leq\frac{C_{b}}{\nu(K)}\langle f,-\mathcal{L}f\rangle_{\nu}$
(49) $\Leftrightarrow\sum_{q\in K}b(q)\nu(q)(f(q)-m)^{2}$ $\leq\left(\sum_{q\in K}b(q)\nu_{K}(q)\right)C_{b}\langle f,-\mathcal{L}f\rangle_{\nu}.$
From Lemma 1 and (47), (49), we have:
(50) $\operatorname{Var}_{\nu}(f)\leq\left\lVert f-m\mathbf{1}\right\rVert^{2}_{2,\nu}\leq\frac{1+\left(\sum_{q\in\mathcal{S}}b(q)\nu_{K}(q)\right)C_{b}}{\gamma}\langle f,-\mathcal{L}f\rangle_{\nu}\,\forall f\in\ell_{2,\nu},$
and hence, the system admits a Poincaré constant $C_{P}=\frac{1+\left(\sum_{q\in\mathcal{S}}b(q)\nu_{K}(q)\right)C_{b}}{\gamma}$.
A.2.2. Proof of Corollary 2
Proof.
Let $b(q)=B1_{K}(q)$ and $m=\frac{\sum_{q\in K}\nu(q)f(q)}{\nu(K)}$, we have
(51) $\tau(q)=\frac{\nu(q)1_{K}(q)}{\sum_{x\in K}\nu(x)}=\frac{\nu(q)1_{K}(q)}{\nu(K)}=\nu_{K}(q).$
We have Assumption 4 satisfied for constant $C_{L}$ means that
(52) $\operatorname{Var}_{\nu_{K}}(f)\leq\frac{C_{L}}{\nu(K)}\langle f,-\mathcal{L}f\rangle_{\nu}\,\forall f\in\ell_{2,\nu}$
which gives
(53) $\sum_{q\in\mathcal{S}}b(q)\nu(q)(f(q)-m)^{2}=B\nu(K)\operatorname{Var}_{\nu_{K}}(f)\leq BC_{L}\langle f,-\mathcal{L}f\rangle_{\nu}.$
Substitute this into equation (47) with our choice of $m$ and apply Lemma 1, we have
(54) $\operatorname{Var}_{\nu}(f)\leq\frac{1+BC_{L}}{\gamma}\langle f,-\mathcal{L}f\rangle_{\nu}\,\forall f\in\ell_{2,\nu},$
as desired. ∎
A.2.3. Proof of Proposition 7
Observe that if we have the finite set $K$ is a singleton, then we can show the local Poincaré constant of the singleton set is 00. Indeed, let $K=\{x^{*}\}$ and the local measure corresponds to the finite set $K$ be $\nu_{K}$, then we have $E_{\nu_{K}}(f)=f(x^{*})$ and so $\operatorname{Var}_{\nu_{K}}(f)=\mathbb{E}_{\nu_{K}}\left[(f-\mathbb{E}_{\nu_{K}}(f))^{2}\right]=0$. And so, we can perform a similar analysis to the proof of Theorem 1 but slightly modify it so that we cab ignore the positive drift term inside the finite set (which happens to be a singleton as well) and bypass the requirement that $V\geq 1$ everywhere.
Proposition 0.
Assume that Assumptions 1 and 3 holds where the set $\{b(q)>0\}\subseteq\mathcal{S}$ is a singleton, then the CTMC admits the Poincaré constant $C_{P}=\frac{1}{\gamma}$.
Proof.
Let $K=\{x^{*}\}$, we will follow a similar approach to the proof of Theorem 1 but with a slight modification. From Assumption 2, we have
(55) $\mathcal{L}V(q)\leq-\gamma V(q)\,\forall q\neq x^{*}.$
This gives
(56) $\frac{(f(q)-f(x^{*}))^{2}}{V(q)}\mathcal{L}V(q)\leq-\gamma(f(q)-f(x^{*}))^{2}\quad\forall q\neq x^{*}.$
Summing this over $q\in\mathcal{S}-\{x^{*}\}$ with weight $\pi(q)$, we have
$\sum_{q\neq x^{*}}\pi(q)\frac{(f(q)-f(x^{*}))^{2}}{V(q)}\mathcal{L}V(q)$ $\leq-\gamma\left\lVert f-f(x^{*})\right\rVert_{2,\pi}^{2}$
(57) $\implies\gamma\left\lVert f-f(x^{*})\right\rVert_{2,\pi}^{2}$ $\leq\sum_{q\neq x^{*}}-\frac{\pi(q)(f(q)-f(x^{*}))^{2}}{V(q)}\mathcal{L}V(q).$
Now, we want to upper bound the RHS of (57). Denote $g=f-f(x^{*})$ We have
$\sum_{q\neq x^{*}}-\frac{\pi(q)(f(q)-f(x^{*}))^{2}}{V(q)}\mathcal{L}V(q)\overset{(a)}{=}-\sum_{q\neq x^{*}}\frac{\pi(q)g(q)^{2}}{V(q)}\sum_{q^{\prime}\in\mathcal{S}}\mathcal{L}(q,q^{\prime})V(q^{\prime})\qquad\qquad\qquad\qquad\qquad\qquad$
$=-\sum_{q\in\mathcal{S}-\{x^{*}\}}\frac{\pi(q)g(q)^{2}}{V(q)}\left[\mathcal{L}(q,x^{*})V(x^{*})+\sum_{q^{\prime}\in\mathcal{S}-\{x^{*}\}}\mathcal{L}(q,q^{\prime})V(q^{\prime})\right]$
$\overset{(b)}{\leq}-\sum_{q,q^{\prime}\in\mathcal{S}-\{x^{*}\}}\pi(q)g(q)^{2}\mathcal{L}(q,q^{\prime})\frac{V(q^{\prime})}{V(q)}$
$=-\frac{1}{2}\sum_{q,q^{\prime}\in\mathcal{S}-\{x^{*}\}}V(q)V(q^{\prime})Q(q,q^{\prime})\left[\frac{g(q)^{2}}{V(q)^{2}}+\frac{g(q^{\prime})^{2}}{V(q^{\prime})^{2}}\right]$
$\overset{(c)}{\leq}-\sum_{q,q^{\prime}\in\mathcal{S}-\{x^{*}\}}Q(q,q^{\prime})g(q)g(q^{\prime})$
$=-\sum_{q,q^{\prime}\in\mathcal{S}}\pi(q)\mathcal{L}(q,q^{\prime})g(q)g(q^{\prime})$
$=\langle g,-\mathcal{L}g\rangle_{\pi}$
(58) =$\overset{(d)}{=}\langle f,-\mathcal{L}f\rangle_{\pi}$
where $Q(x,y)=\pi(x)\mathcal{L}(x,y)=Q(y,x)$ from the reversibility of the Markov chain. $(a)$ follows from the definition of generator, which is
$\mathcal{L}V(q)=\sum_{q^{\prime}\in\mathcal{S}}V(q,q^{\prime})V(q^{\prime}).$
(b) follows from the property of generators that $\mathcal{L}(x,y)\geq 0\,\forall x\neq y$ and the fact that $V(q)\geq 0$ for all $q\in\mathcal{S}$. (c) follows from the AM-GM inequality and noting that $V(q)>0$ for all $q\neq x^{*}$ and (d) is from (43). Apply the upper bound (58) to (57) and notice that $\operatorname{Var}_{\pi}(f)\leq\left\lVert f-f(x^{*})\right\rVert_{2,\pi}^{2}$, we have
$\operatorname{Var}_{\pi}(f)\leq\left\lVert f-f(x^{*})\right\rVert_{2,\pi}^{2}\leq\frac{1}{\gamma}\langle f,\mathcal{L}f\rangle_{\pi}.$
Which implies the system admits the Poincaré constant $\frac{1}{\gamma}$. ∎
Remark: By avoiding $x^{*}$ in the proof, we can avoid the division by 00 issue even in the case $V(x^{*})=0$ and lifts the restriction $V(q)\geq 1\,\forall q\in\mathcal{S}$ as in the general case. Thus, the advantage of having $K$ as a singleton would simplify the Lyapunov-Poincaré proof by a wide margin. Furthermore, this allows us to lift the stringent $V\geq 1$ assumption that is commonly used in many Markov chain mixing works (Taghvaei and Mehta, 2022; Meyn and Tweedie, 1994), which somewhat restricts our Lyapunov choice. It turns out that allowing $V(x^{*})=0$ would allow us to obtain the tight mixing rate, which we will see in the $M/M/\infty$ setting and the Sub-Halfin-Whitt case. In addition, since the final Poincaré constant does not depend on $b$ nor the local Poincaré constant $C_{L}$ with respect to the set $K$, we will not suffer any loss in terms of the mixing rate and the obtained Poincaré constant would be tight when we have $K$ being a singleton given that we have the right Lyapunov drift.
A.3. Proof of Proposition 1
Proposition 0 (Restatement of Proposition 1).
Given two distributions $P,Q$ with countable supports $\mathcal{D}$ and let $\mathcal{G}_{Q}$ be a collection of functions $g:\mathcal{D}\rightarrow\mathbb{R}$ such that $\mathbb{E}_{Q}\left[g(X)^{2}\right]<\infty\}$, we have
(59) $\chi^{2}(P,Q)=\sup_{g\in\mathcal{G}_{Q}}\frac{\left(\mathbb{E}_{P}[g(X)]-\mathbb{E}_{Q}[g(X)]\right)^{2}}{\operatorname{Var}_{Q}[g(X)]}$
The following proof follows from Example 7.4 in (Polyanskiy and Wu, 2024).
Proof.
Let $f(x)=(x-1)^{2}$, we have the conjugate $f^{*}_{ext}(y)=y+\frac{y^{2}}{4}$. From Theorem 7.26 in (Polyanskiy and Wu, 2024), we have
(60) $\chi^{2}(P,Q)=D_{f}(P||Q)$ $=\sup_{h:\mathbb{R}\rightarrow\mathbb{R}}\mathbb{E}_{P}\left[h(X)\right]-\mathbb{E}_{Q}\left[f^{*}_{ext}(h(X))\right]$
(61) $=\sup_{h:\mathbb{R}\rightarrow\mathbb{R}}\mathbb{E}_{P}\left[h(X)\right]-\mathbb{E}_{Q}\left[h(X)+\frac{h(X)^{2}}{4}\right]\text{ from }f^{*}_{ext}(y)=y+\frac{y^{2}}{4}$
(62) $=\sup_{h:\mathbb{R}\rightarrow\mathbb{R}}2\mathbb{E}_{P}\left[h(X)\right]-\mathbb{E}_{Q}\left[h(X)^{2}\right]-1$
where the last step we perform a change of variable $h\leftarrow\frac{h}{2}-1$. Let $h=\lambda g$, we have
(63) $\chi^{2}(P,Q)$ $=\sup_{\lambda\in\mathbb{R}}\sup_{g:\mathbb{R}\rightarrow\mathbb{R}}2\lambda\mathbb{E}_{P}\left[g(X)\right]-\lambda^{2}\mathbb{E}_{Q}\left[g(X)^{2}\right]-1$
(64) $=\sup_{g:\mathbb{R}\rightarrow\mathbb{R}}\frac{(\mathbb{E}_{P}[g(X)]-\mathbb{E}_{Q}[g(X)])^{2}}{\operatorname{Var}_{Q}(g(X))}$
after optimizing over $\lambda$. ∎
Appendix B Proof of Theorem 3
In this Section, we present the proofs of the main mixing Theorem (Theorem 3). Since there is a phase transition at $\alpha=1/2$, we split the Theorem into two parts: the Super-Halfin-Whitt regime (when $\alpha\in(1/2,\infty)$) and the Sub-Halfin-Whitt regime (when $\alpha\in(0,1/2)$).
In particular, we present the complete proof for the mixing results in the Super-Halfin-Whitt regime Subsection B.1. We first obtain the drift analysis in Subsection B.1.1 and then obtain the local Poincaré inequality for the finite set via the local canonical path method in Appendix B.1.3. Then, as the final step, we put them together using the Stitching Theorem 1 in Appendix B.1.4.
In Subsection B.3, we present the proof for the mixing results in the Sub-Halfin-Whitt regime. We first perform the drift analysis in Appendix B.3.1 and then obtain the mixing time bound in Appendix B.3.2. Additionally, we investigate the case when $\lambda_{n}$ is an integer and show that the mixing rate approaches $1$ at the asymptotic in Appendix B.3.3.
B.1. Proof of Proposition 2 and Theorem 3 for the Super-Halfin-Whitt regime
In the following Subsection, we provide the proof of the mixing time results for the $M/M/1$ system and the $M/M/n$ system in the Super-Halfin-Whitt regime, that is $\alpha\in(1/2,\infty)$. For the final step of the proof, please refer to Appendix B.1.4. For the drift lemma (Lemma 1), we refer the reader to Appendix B.1.1. To show Lemma 2, that is to show that $\nu_{K}$ is roughly uniform, we refer the reader to Appendix B.1.2. Finally, to prove the local canonical path Lemma 1 and Claim 1 and Claim 2, we refer the readers to Appendix B.1.3.
B.1.1. Proof of Lemma 1
Proof.
Recall that $V(q)=e^{\theta[q-(n-1)]^{+}+\theta[(n-1)-q]^{+}}$, where we fix $\theta=\log\sqrt{\frac{n}{\lambda_{n}}}$. Now, when $q\geq n$, we have $V(q)=e^{\theta(q-(n-1))}$. We then have
(65) $\mathcal{L}V(q)$ $=\left(\lambda_{n}e^{\theta}+\frac{n}{e^{\theta}}-(\lambda_{n}+n)\right)V(q)=\left(e^{\theta}-1\right)\left(\lambda_{n}-\frac{n}{e^{\theta}}\right)V(q)=-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V(q).$
The choice of $\theta$ is evident in the third equality as it minimizes $\left(e^{\theta}-1\right)\left(\lambda_{n}-\frac{n}{e^{\theta}}\right)$. For $q<n$, we have $V(q)=e^{\theta(n-1-q)}$ and now consider two cases: $n=1\leq n\leq 7$, and $n\geq 8$.
Case 1 $(1\leq n\leq 7)$: We have $K=\{0,1,...,n-1\}$. From (65), we have $\mathcal{L}V\leq-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V\,\,\forall x\not\in K$. If $x\in K$ then $q\leq n-1$ and since $n\leq 7$ we have that there is only a finite number of pairs $(n,q)$. Thus, we have
$\mathcal{L}V(q)\leq-\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}V(q)+L1_{K}$
for some constant $L$, where it suffices to consider the maximum value of $\mathcal{L}V(q)$ taken over all choice of $q\in K$ and $1\leq n\leq 7$. Hence, the case $1\leq n\leq 7$ is done.
Case 2 $(n>7)$: We have $K=\{\lceil 2\lambda_{n}\rceil-n,\ldots,n\}$. This gives $V(q)=e^{\theta(n-1-q)}$ for $q\leq n-1$. So, for $q<n-1$, we have
$\mathcal{L}V(q)$ $=\left(e^{\theta}-1\right)\left(q-\frac{\lambda_{n}}{e^{\theta}}\right)V(q)=\left(e^{\theta}-1\right)\bigg{(}q-(2\lambda_{n}-n)+\underbrace{2\lambda_{n}-n-\frac{\lambda_{n}}{e^{\theta}}}_{\leq\lambda_{n}-\sqrt{\lambda_{n}n}}\bigg{)}V(q)$
$\leq-\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}V(q)+\left(e^{\theta}-1\right)\left(q-(2\lambda_{n}-n)\right)V(q)$
For $q\leq\lceil 2\lambda_{n}\rceil-n-1<2\lambda_{n}-n$, we have $q\not\in K$ and
(66) $\mathcal{L}V(q)\leq-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V(q).$
For $22\lambda_{n}-n\leq q\leq n-$ 2, we have $V(q)\leq e^{n^{-\alpha}\times 2n^{1-\alpha}}=e^{2n^{1-2\alpha}}\leq e^{2}=O(1)$ for $\alpha>1/2$. Furthermore
(67) $\theta=\sqrt{\frac{n}{\lambda_{n}}}-1=\frac{\frac{n-\lambda_{n}}{\lambda_{n}}}{\sqrt{\frac{n}{\lambda_{n}}}+1}=\frac{n^{1-\alpha}}{n\left(\frac{\lambda_{n}}{n}+\sqrt{\frac{\lambda_{n}}{n}}\right)}\leq n^{-\alpha}$
since we have $\lambda_{n}\geq 1\geq n^{1-2\alpha}$. Thus
$\mathcal{L}V(q)$ $\leq-\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}V(q)+b1_{K-\{n-1\}}$
where $e^{\theta}-1\leq n^{-\alpha}$ from (67). And so for $q\in K-\{n-1\}$
$\underbrace{\left(e^{\theta}-1\right)}_{O(n^{-\alpha})}\underbrace{\left(q-(2\lambda_{n}-n)\right)}_{O(n^{1-\alpha})}\underbrace{V(q)}_{O(1)}\leq n^{-\alpha}\times 2n^{1-\alpha}\times e^{2}\leq 2e^{2}n^{1-2\alpha}=b(q)$
For $q=n-1$, we have
$\mathcal{L}V(n-1)$ $=(\lambda_{n}+n-1)\left(e^{\theta}-1\right)\leq 2n^{1-\alpha}=-\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}V(n-1)+\bigg{(}2n^{1-\alpha}+\underbrace{\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}}_{\leq n^{1-\alpha}}\bigg{)}$
$\leq-\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}V(n-1)+3n^{1-\alpha}.$
Hence we have for $q\leq n$:
$\mathcal{L}V(q)\leq-\gamma_{n}V(q)+b(q)1_{K}.$
with $\gamma_{n}=\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}$ and $b$ as defined in (28), which completes the proof. ∎
B.1.2. Proof of Lemma 2
Proof of Lemma 2.
Let $n_{0}=65$, we have $\lambda_{n}=n-n^{1-\alpha}\geq\frac{7n}{8}$ and $\lceil 2\lambda_{n}\rceil-n\geq\frac{3n}{4}$ for $\alpha\in(1/2,\infty)$, we want to show $\nu_{K}$ roughly approximates the uniform distribution whose support is the finite set $K$. In particular, we want to prove $\nu_{K}(x)=\Theta\left(n^{\alpha-1}\right)\,\forall x\in K$. From Stirling’s approximation, we have
$\sqrt{2\pi n}\left(\frac{n}{e}\right)^{n}\leq n!\leq\sqrt{2\pi n}\left(\frac{n}{e}\right)^{n}e^{\frac{1}{12n}}.$
Thus, we have
$\frac{e^{-\lambda_{n}}\lambda_{n}^{x}}{x!}$ $\leq e^{-\lambda_{n}}\frac{\lambda_{n}^{x}}{\sqrt{2\pi x}\left(\frac{x}{e}\right)^{x}}=\frac{e^{x-\lambda_{n}}}{\sqrt{2\pi x}}\left(1+\frac{\lambda_{n}-x}{x}\right)^{x}$
$\leq\frac{e^{x-\lambda_{n}}}{\sqrt{2\pi x}}e^{\lambda_{n}-x}=\frac{1}{\sqrt{2\pi x}}\text{ from }1+t\leq e^{t}$
(68) $\leq\sqrt{\frac{2}{3\pi n}}\quad\forall x\in K$
since we have $x\geq\lceil 2\lambda_{n}\rceil-n\geq\frac{3n}{4}$ for $n\geq n_{0}$. We also have from Stirling’s approximation:
(69) $\frac{e^{-\lambda_{n}}\lambda_{n}^{x}}{x!}$ $\geq\frac{e^{x-\lambda_{n}}}{\sqrt{2\pi x}}\left(1+\frac{\lambda_{n}-x}{x}\right)^{x}e^{-\frac{1}{12x}}.$
Note that for $n\geq n_{0}$, we have that $\lceil 2\lambda_{n}\rceil-n\geq 20$, thus we have $0\not\in K$ and $20\leq x\leq n\,\forall x\in K$. We have:
$\frac{e^{-\lambda_{n}}\lambda_{n}^{x}}{x!}$ $\geq\frac{e^{x-\lambda_{n}}}{\sqrt{2\pi x}}\left(1+\frac{\lambda_{n}-x}{x}\right)^{x}e^{-\frac{1}{12x}}\geq\frac{e^{x-\lambda_{n}}}{\sqrt{2\pi n}}\left(1+\frac{\lambda_{n}-x}{x}\right)^{x}e^{-\frac{1}{12}}$
Let $t=\frac{\lambda_{n}}{x}-1$, we have for $n\geq n_{0}$ that $|t|\leq\frac{4}{3}n^{-\alpha}\,\forall x\in K$ and $1+t\leq e^{t}\leq 1+t+t^{2}e^{t}$ from Lemma 1. This gives
$\left(\frac{1+t}{e^{t}}\right)^{x}=\left(1+\frac{1+t-e^{t}}{e^{t}}\right)^{x}\geq\left(1-t^{2}\right)^{x}\geq\left(1-\frac{16}{9n^{2\alpha}}\right)^{\frac{9n^{2\alpha}}{16}\times\frac{16}{9}n^{1-2\alpha}}\geq\delta^{\frac{16}{9}n^{1-2\alpha}}\geq\delta^{\frac{16}{9}}\,\,\forall n\geq n_{0}$
where $\delta=\left(1-\frac{16}{9}n_{0}^{-2\alpha}\right)^{9n_{0}^{2\alpha}/16}\in\left(\left(1-\frac{1}{36}\right)^{36},1\right)\,\forall\alpha>1/2$. From here, we can further lower bound $\delta^{\frac{16}{9}}\geq\frac{13}{80}$. Putting this back to (69), we obtain
(70) $\frac{e^{-\lambda_{n}}\lambda_{n}^{x}}{x!}\geq\frac{13}{80}\frac{e^{-\frac{1}{12}}}{\sqrt{2\pi x}}\geq\frac{13}{80}\frac{e^{-\frac{1}{12}}}{\sqrt{2\pi n}}\,\forall x\in K,n\geq n_{0}.$
This implies that:
[…]
Hence, we have shown that $\sum_{x\in K}e^{-\lambda_{n}}\lambda_{n}^{x}/x!=\Theta\left(n^{1/2-\alpha}\right)$, which gives
(71) $\nu_{K}(x)$ $=\frac{e^{-\lambda_{n}}\lambda_{n}^{x}/x!}{\sum_{x\in K}e^{-\lambda_{n}}\lambda_{n}^{x}/x!}\geq\frac{13}{80}\frac{e^{\frac{-1}{12}}\sqrt{3}n^{\alpha-1}}{4}=\frac{13\sqrt{3}e^{-\frac{1}{12}}}{320}n^{\alpha-1}\text{ and }$
(72) $\nu_{K}(x)$ $=\frac{e^{-\lambda_{n}}\lambda_{n}^{x}/x!}{\sum_{x\in K}e^{-\lambda_{n}}\lambda_{n}^{x}/x!}\leq\frac{80}{13}e^{\frac{1}{12}}\frac{2}{\sqrt{3}}n^{\alpha-1}=\frac{160e^{\frac{1}{12}}}{13\sqrt{3}}n^{\alpha-1}.$
And so, we have shown that $\nu_{K}(x)=\Theta\left(n^{\alpha-1}\right)=\Theta\left(|K|^{-1}\right)$, which implies that $\nu_{K}$ is a roughly uniform distribution. ∎
B.1.3. Proof of Lemma 2
Before moving to the proofs of the claims, observe that $f\in\ell_{2,\nu_{n}}$ trivially implies $f\in\ell_{2,\nu_{K}}$ and $f\in\ell_{2,\nu_{K}}$ is equivalent to $f\in\ell_{2,\tau}$ for $\tau(q)=b(q)\nu_{K}(q)\,\forall q\in\mathcal{S}$, given that $b$ is sufficiently well-behaved. This essentially tells us that if a test function $f$ is valid for the global distribution then it also works for the weighted distribution and the local distribution. Now, we will prove the claims as follows.
Proof of Claim 1.
Denote $\tau(q)=b(q)\nu_{K}(q)\,\forall q\in\mathbb{Z}_{\geq 0}$, we have
$T_{1}$ $=\sum_{x\in K}\tau(n-1)\tau(x)\left(f(n-1)-f(x)\right)^{2}$
[…]
$=\sum_{x\in K}\left(\frac{160\sqrt{6}}{13}e^{\frac{13}{12}}\right)^{2}n^{-\alpha}\left(f(n-1)-f(x)\right)^{2}$
$\leq 2\left(\frac{160\sqrt{6}}{13}e^{\frac{13}{12}}\right)^{2}n^{1-2\alpha}\max_{x\in K}\left(f(n-1)-f(x)\right)^{2}\text{ since the cardinality of $K$ is at most }2n^{1-\alpha}$
$\leq 2\left(\frac{160\sqrt{6}}{13}e^{\frac{13}{12}}\right)^{2}n^{1-2\alpha}\times 2n^{1-\alpha}\sum_{x\in K^{\prime}}\left(f(x+1)-f(x)\right)^{2}$
$=\left(\frac{320\sqrt{6}}{13}e^{\frac{13}{12}}\right)^{2}n^{2-3\alpha}\sum_{x\in K^{\prime}}\left(f(x+1)-f(x)\right)^{2}.$
Here, the last inequality is obtained from applying Cauchy-Schwarz along the path $n-1$ to $x$. Recall that $K^{\prime}=K-\{n-1\}\Rightarrow K=K^{\prime}\cup\{n-1\}$. For $e=(x,x+1)$ where $x,x+1\in K^{\prime}=K-\{n-1\}$, we have that $Q_{K}(e)=Q_{K}(x,x+1)=\nu_{K}(x)\mathcal{L}(x,x+1)=\lambda_{n}\nu_{K}(x)$. Again, recall that $n\geq n_{0}$ and so $\frac{7n}{8}\leq\lambda_{n}\leq n$. From Lemma 2, this gives
(73) $\frac{91\sqrt{3}e^{-\frac{1}{12}}}{2560}n^{\alpha}=\frac{7n}{8}\times\frac{13\sqrt{3}e^{-\frac{1}{12}}}{320}n^{\alpha-1}\leq Q_{K}(x,x+1)=\lambda_{n}\nu_{n}(x)$
for all $x\in K^{\prime}=K-\{n-1\}$. And so we can bound $T_{1}$ as
$T_{1}$ $\leq\left(\frac{320\sqrt{6}}{13}e^{\frac{13}{12}}\right)^{2}n^{2-3\alpha}\sum_{x\in K^{\prime}}\left(f(x+1)-f(x)\right)^{2}$
$=\left(\frac{320\sqrt{6}}{13}e^{\frac{13}{12}}\right)^{2}n^{2-4\alpha}\frac{1}{\frac{91\sqrt{3}e^{-\frac{1}{12}}}{2560}}\times\frac{91\sqrt{3}e^{-\frac{1}{12}}}{2560}n^{\alpha}\sum_{k:k,k+1\in K}\left(f(k)-f(k+1)\right)^{2}$
[…]
$\leq 560229n^{2-4\alpha}\sum_{k:k,k+1\in K}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1).$
This completes the proof. ∎
In summary, the proof of Claim 1 is simply performing Cauchy-Schwarz along the path joining $n-1$ and $x$. For Claim 2, we will need to be more involved.
Proof of Claim 2.
The second term $T_{2}$ can be bounded using the local canonical path method. First, we bound the number of paths passing through $e$. For $e=(k,k+1)$ where $n-1\geq k\geq\lceil 2\lambda_{n}\rceil-n$, note that there are $n+k-\lceil 2\lambda_{n}\rceil$ ways to choose $x\leq k$ and $n-k-1$ ways to choose $y\geq k+1$. Hence, there are at most $\left(n+k-\lceil 2\lambda_{n}\rceil\right)\times\left(n-k-1\right)\leq\left(\frac{2n-\lceil 2\lambda_{n}\rceil-1}{2}\right)^{2}\leq\left(n-\lambda_{n}\right)^{2}=n^{2-2\alpha}$ paths containing the edge $e$, each with length at most $2(n-\lambda_{n})=2n^{1-\alpha}$. Let $K^{\prime}=K-\{n-1\}$, observe that $\tau(q)=b(q)\nu_{K}(q)=2e^{2}n^{1-2\alpha}\nu_{K}(q)\forall q\in K^{\prime}$, i.e. $\tau(q)$ is proportional to $\nu_{K}(q)$ whenever $q\in K^{\prime}$. Thus, we can bound $T_{2}$ using Lemma 1 with respect to the distribution $\tau$ as follows
$T_{2}=\frac{1}{2}\sum_{x,y\in K^{\prime}}\tau(x)\tau(y)(f(x)-f(y))^{2}$
$=4e^{4}n^{2-4\alpha}\times\frac{1}{2}\sum_{x,y\in K^{\prime}}\nu_{K}(x)\nu_{K}(y)(f(x)-f(y))^{2}$
$\overset{(a)}{\leq}4e^{4}n^{2-4\alpha}\max_{k:k,k+1\in K^{\prime}}\frac{\sum_{x,y\in K^{\prime}:x\leq k\leq y-1}|x-y|\nu_{K}(x)\nu_{K}(y)}{Q_{K}(k,k+1)}\sum_{k:k,k+1\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1)$
$\overset{(b)}{\leq}4e^{4}n^{2-4\alpha}\times\frac{n^{2-2\alpha}\times 2n^{1-\alpha}\times\frac{160e^{\frac{1}{12}}}{13\sqrt{3}}n^{\alpha-1}\times\frac{160e^{\frac{1}{12}}}{13\sqrt{3}}n^{\alpha-1}}{\frac{91\sqrt{3}e^{-\frac{1}{12}}}{2560}n^{\alpha}}\times\sum_{k:k,k+1\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1)$
$=\frac{2^{22}e^{\frac{17}{4}}}{7}\left(\frac{5\sqrt{3}}{39}\right)^{3}n^{3-6\alpha}\sum_{k:k,k+1\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1)$
$\leq 459950.915\times n^{3-6\alpha}\sum_{k:k,k+1\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1)$
$\leq 459951\times n^{3-6\alpha}\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1).$
Here, we have (a) follows from Lemma 1 and (b) follows from the number of paths upper bound and the bounds on $\nu_{K}$ (from Lemma 2) and the lower bound of $Q_{K}$ (73). ∎
Proof of Claim 3.
We will show that the total mass of $\tau$ is only a constant factor different from the total mass of $\nu_{K}$. Indeed, from (35) and (36), we have
$\sum_{q\in K}b(q)\nu_{K}(q)=\sum_{q\in K}\tau(q)$ $=\tau(n-1)+\sum_{q\in K-\{n-1\}}\tau(q)$
$\leq\frac{480\sqrt{3}}{13}e^{\frac{1}{12}}+2n^{1-\alpha}\times\frac{320e^{2}}{13\sqrt{3}}e^{\frac{1}{12}}n^{-\alpha}$
$\leq\frac{480\sqrt{3}}{13}e^{\frac{1}{12}}+\frac{640e^{2}}{13\sqrt{3}}e^{\frac{1}{12}}n^{1-2\alpha}$
(74) $\leq\frac{480\sqrt{3}}{13}e^{\frac{1}{12}}+\frac{640e^{2}}{13\sqrt{3}}e^{\frac{1}{12}}\leq 339.1841\leq 340\,\forall n\geq 1,\alpha\geq 1/2.$
So the total mass of $\tau$ is at most a constant factor more than the total mass of $\nu_{K}$. ∎
B.1.4. Final step: Proof of Theorem 3 and Proposition 2
Since we have obtained all necessary ingredients, we shall provide the proof of Theorem 3 in the Super-Halfin-Whitt regime below for completeness. Additionally, the proof of Proposition 2 is also included here.
Proof.
We will consider two separate cases: the case where $\alpha\geq 1$ or $n=1$ and the case where $\alpha\in(1/2,1)$.
Case 1 ($\alpha\geq 1$ or $n=1$): From Proposition 7, Lemma 1 and observe that in this case, we have $K=\{n-1\}$, we have that the system admits the Poincaré constant
$C_{P}=\frac{1}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}.$
And so, from Proposition 6, we have
(75) $\chi(\pi_{n,t},\nu_{n})\leq e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n}).$
Case 2 ($\alpha\in(1/2,1)$): From Lemma 1, Lemma 2, Claim 3 and Theorem 1, we obtain the Poincaré constant to be:
$C_{P}$ $=\frac{1+\left(\sum_{q\in K}b(q)\nu_{K}(q)\right)C_{b}}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}\leq\frac{1+339.1841\times\left(\Delta_{1}n^{2-4\alpha}+\Delta_{2}n^{3-6\alpha}\right)}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}$
Let
(76) $C_{n}=\frac{1}{1+340\times\left(\Delta_{1}n^{2-4\alpha}+\Delta_{2}n^{3-6\alpha}\right)},$
we have that
$\lim_{n\rightarrow\infty}C_{n}=1$
since $\alpha>1/2$. And so
$\lim_{n\rightarrow\infty}C_{n}=\lim_{n\rightarrow\infty}\frac{1}{1+C_{b}}=1.$
From Proposition 6, we have
(77) $\chi(\pi_{n,t},\nu_{n})\leq e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})$
where $\lim_{n\rightarrow\infty}C_{n}=1$. ∎
B.2. Proof of Theorem 3 for the Halfin-Whitt regime
For the $\alpha=1/2$ regime, we are not able to obtain a tight finite-time bound and characterize a phase transition as in (Gamarnik and Goldberg, 2013). Yet, we can still show that the mixing is bounded below by some universal constant by performing a similar analysis as in Appendix B.1 as follows.
Theorem 0 (Restatement of Theorem 3 for $\alpha=1/2$).
Let $\pi_{n,t}$ be the queue length distribution at time $t$ of the continuous-time $M/M/n$ system with the arrival rate $\lambda=n-n^{1-\alpha}$, the service rate $1$ and the stationary distribution be $\nu_{n}$. For $\alpha=1/2$, we have
(78) $\chi(\pi_{n,t},\nu_{n})\leq e^{-\frac{t}{1387444804}}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0.$
Proof.
Let $K=\{\lceil 2\lambda\rceil-n,...,n-1\}$, choosing the same Lyapunov function $V$ as in Lemma 1, we have that
(79) $\mathcal{L}V(q)\leq-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V(q)+b(q)$
where $b(n-1)\leq 3\sqrt{n}$ and $b(q)\leq L\,\forall q\in K-\{n-1\}$ for some constant $L$. Next, from Lemma 2 for $\alpha=1/2$, we have $\Delta_{1}=560229,\Delta_{2}=459951$, and so
(80) $C_{b}=\Delta_{1}+\Delta_{2}=560229+459951=1020180.$
Applying the Stitching Theorem 1, we obtain the Poincaré constant bound
(81) $C_{P}=\frac{1+\left(\sum_{q\in K}b(q)\nu_{K}(q)\right)C_{b}}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}.$
Since $\lambda_{n}=n-\sqrt{n}$, we have for $n\geq 2$:
(82) $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}=\left(\frac{n-\lambda_{n}}{\sqrt{n}+\sqrt{n-\sqrt{n}}}\right)^{2}=\frac{1}{\left(1+\sqrt{1-1/\sqrt{n}}\right)^{2}}\geq\frac{1}{4}$
Which implies that $C_{P}\leq 4\times(1+340\times 1020180)\leq 1387444804$. And so, from Proposition 6, we have
(83) $\chi(\pi_{n,t},\nu_{n})\leq e^{-\frac{t}{1387444804}}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0.$
The proof is done. ∎
Remark on the constant: It is noteworthy that our goal here is to show that the mixing rate is bounded below by some universal constant, which is $\frac{1}{1387444804}$ in this case. Since we are not able to obtain a convergence rate that matches the limiting behavior, we have not made any effort to optimize this constant. Based on the spectral gap characterization at Halfin-Whitt in (Gamarnik and Goldberg, 2013), it seems that characterizing the phase transition in this regime is rather non-trivial.
B.3. Proof of Theorem 3 in the Sub-Halfin-Whitt regime
To prove the Theorem 3 for $\alpha\in(0,1/2)$ (Equation (8) in Theorem 3), we will first perform a drift analysis and then apply the Lyapunov-Poincaré method. Similar to the Super-Halfin-Whitt case, we will first establish the negative drift result in Appendix B.3.1 and prove Equation (8) in Appendix B.3.2.
B.3.1. Proof of Lemma 3
Lemma 0 (Restatement of Lemma 3).
For a sufficiently large $n$ such that $\lambda_{n}=n-n^{1-\alpha}\geq 3$, set $V(q)=\zeta e^{\theta(q-\lambda_{n})}\forall q>n$. If $\lambda_{n}\in\mathbb{Z}_{\geq 0}$, set $V(q)=|q-\lambda_{n}|\forall q\leq n$. Otherwise, set
(84) $V(q)=\begin{cases}|q-\lambda_{n}|\,&\forall q\leq n,q\not\in\{\lfloor\lambda_{n}\rfloor,\lceil\lambda_{n}\rceil\}\\ |\lfloor\lambda_{n}\rfloor-\lambda_{n}-1|&\textit{if }q=\lfloor\lambda_{n}\rfloor\\ |\lceil\lambda_{n}\rceil-\lambda_{n}+1|&\textit{if }q=\lceil\lambda_{n}\rceil\end{cases}$
where $\zeta$ is chosen such that $|n-\lambda_{n}|=\zeta e^{\theta(n-\lambda_{n})}$ and $\theta>0$ is a parameter to be chosen. Then, there exists $\gamma_{n},b$ with $\gamma_{n}\rightarrow 1$, such that
$\mathcal{L}V\leq-\gamma_{n}V+b1_{K}$
where $b\leq 2\lambda_{n}$ and $K=\{\lambda_{n}\}$ if $\lambda_{n}\in\mathbb{Z}_{\geq 0}$ and $K=\{\lfloor\lambda_{n}\rfloor-1,\lfloor\lambda_{n}\rfloor,\lceil\lambda_{n}\rceil,\lceil\lambda_{n}\rceil+1\}$ otherwise.
Proof.
Due to the discrete nature of the state space, we will consider two cases: $\lambda_{n}\in\mathbb{Z}$ and $\lambda_{n}\not\in\mathbb{Z}$. In both cases, we choose
$V(q)=\zeta e^{\theta(q-\lambda_{n})}\forall q>n$
where $\zeta$ is chosen such that $\zeta e^{\theta n^{1-\alpha}}=\zeta e^{\theta(n-\lambda_{n})}=|n-\lambda_{n}|=n^{1-\alpha}$. For $\theta$, we choose
$\theta=\log\left(1+n^{\alpha-1}\right)>0$
for some $\delta>0$. Since $\theta>0$, we have that $e^{\theta}>1$ and so $V(q)\geq V(n)\geq 1\forall q\geq n$. For $q\leq n$, our choice of $V$ will be slightly different between the two cases, which we will discuss in detail below.
Case 1: Assume that $\lambda_{n}\in\mathbb{Z}$, we choose $V(q)=|q-\lambda_{n}|\,\forall q\leq n$. We have that $V(q)\geq 1\forall q\leq n,q\neq\lambda_{n}$. In this case, our analysis is relatively simple since we will have a negative drift outside of the singleton $\{\lambda_{n}\}$.
For $q<n$ and $q\neq\lambda_{n}$, we have:
$\mathcal{L}V$ $=\lambda_{n}|q+1-\lambda_{n}|+q|q-1-\lambda_{n}|-(\lambda_{n}+q)|q-\lambda_{n}|$
$=-|q-\lambda_{n}|=-V.$
For $q=\lambda_{n}$, we have $V(\lambda_{n})=0,V(\lambda_{n}-1)=V(\lambda_{n}+1)=1$. And so
$\mathcal{L}V(\lambda_{n})$ $=\lambda_{n}+\lambda_{n}=2\lambda_{n}-V(\lambda_{n})$
For $q>n$, we have:
$\mathcal{L}V$ $=\left(e^{\theta}-1\right)\left(\lambda_{n}-\frac{n}{e^{\theta}}\right)V$
$=n^{\alpha-1}\left(\lambda_{n}-\frac{n}{1+n^{\alpha-1}}\right)V$
$=n^{\alpha-1}\left(\frac{n^{\alpha}}{1+n^{\alpha-1}}-n^{1-\alpha}\right)V$
$=\left(\frac{n^{2\alpha-1}}{1+n^{\alpha-1}}-1\right)V$
for $\alpha\in(0,1/2)$. Note that since $\frac{n^{2\alpha-1}}{1+n^{\alpha-1}}<1\forall n\geq 1$ and for $\alpha\in(0,1/2)$, we have this is a negative drift.
For $q=n$, we have:
$\mathcal{L}V(n)$ $=\lambda_{n}\zeta e^{\theta(n+1-\lambda_{n})}+n|n-1-\lambda_{n}|-(\lambda_{n}+n)|n-\lambda_{n}|$
$=\lambda_{n}\left(e^{\theta}|n-\lambda_{n}|-|n+1-\lambda_{n}|\right)-|n-\lambda_{n}|$
$=\left(\lambda_{n}\left(e^{\theta}-1\right)-\frac{\lambda_{n}}{n-\lambda_{n}}-1\right)V(n)$
$=\left((n-n^{1-\alpha})n^{\alpha-1}-\frac{n-n^{1-\alpha}}{n^{1-\alpha}}-1\right)V(n)$
$=\left(n^{\alpha}-1-n^{\alpha}+1-1\right)V(n)$
$=-V(n).$
This gives us the drift condition for $\lambda_{n}\in\mathbb{Z}$ and some constant $b\geq 0$ (in this case, we don’t have to compute $b$ since $K$ is a singleton):
$\mathcal{L}V$ $\leq-\min\left\{1-\frac{n^{2\alpha-1}}{1+n^{\alpha-1}},1\right\}V+b1_{\{\lambda_{n}\}}=-\left(1-\frac{n^{2\alpha-1}}{1+n^{\alpha-1}}\right)V+b1_{\{\lambda_{n}\}}.$
Note that the rate of the negative drift $\gamma_{n}=1-\frac{n^{2\alpha-1}}{1+n^{\alpha-1}}$ is positive since we are consider $n\geq\lambda_{n}\geq 3$.
Case 2: Assume that $\lambda_{n}\not\in\mathbb{Z}$, we denote $r=\lambda_{n}-\lfloor\lambda_{n}\rfloor$ to be the fractional part of $\lambda_{n}$. Note that in this case, however, choosing the same Lyapunov function as in Case 1 will no longer give us the finite set $K$ as a singleton. And so, we will get a different finite set $K$, and we need to slightly modify the Lyapunov function $V$ inside $K$.
From our choice of Lyapunov function (30), we have that $V(q)\geq 1\,\forall q\in\mathbb{Z}_{\geq 0}$. Perform a similar drift analysis to the case $\lambda_{n}\in\mathbb{Z}_{\geq 0}$, we have that for $q\not\in K$:
$\mathcal{L}V(q)\leq-\left(1-\frac{n^{2\alpha-1}}{1+n^{\alpha-1}}\right)V(q).$
As $1\leq V(q)\leq 2$ for $q\in K$ and either $V(q)=V(q+1)$ or $V(q)=V(q-1)$ for $q\in K$, we have for all $q\in K-\{\lceil\lambda_{n}\rceil+1\}$
$\mathcal{L}V(q)=\lambda_{n}V(q+1)+qV(q-1)-(\lambda_{n}+q)V(q)\leq\max\{\lambda_{n},q\}\leq\lceil\lambda_{n}\rceil\leq\lceil\lambda_{n}\rceil+2-V(q).$
Similarly, for $q=\lceil\lambda_{n}\rceil+1$, as $V(q-1)=V(q)$, we get
$\mathcal{L}V(q)=\lambda_{n}V(q+1)+qV(q-1)-(\lambda_{n}+q)V(q)\leq\lambda_{n}\leq\lceil\lambda_{n}\rceil+2-V(q).$
Thus, we have $b\leq\lceil\lambda_{n}\rceil+2$. Since $\lambda_{n}\geq 3$, we have $b\leq\max\{\lceil\lambda_{n}\rceil+2,2\lambda_{n}\}\leq 2\lambda_{n}$.
As we can see in the drift analysis, the main difference between the two cases is a different finite set $K$, which would affect the tightness of our mixing rate bound since the local mixing bound will not be strong enough to cancel the $b$ term (we refer the readers to Appendix B.3.2 for the full proof). Furthermore, a crucial observation here is that $\lim_{n\rightarrow\infty}\gamma_{n}=1$. This will be the foundation to show that the mixing rate of the system in the Sub-Halfin-Whitt regime is bounded below by some universal constant. Moreover, we can only obtain $\lim_{n\rightarrow\infty}\gamma_{n}=1$ when $\alpha\in(0,1/2)$, and so this drift analysis is only applicable in this regime. For the regime $\alpha\in[1/2,\infty)$, we will need to redo the drift analysis in order to obtain a good mixing bound, as done in Appendix B.1.
B.3.2. Proof of Theorem 3 for $\alpha\in(0,1/2)$
Before we go to the proof of Theorem 3 for $\alpha\in(0,1/2)$, we need to establish a result that is analogous to Lemma 2 for the Sub-Halfin-Whitt regime as follows.
Lemma 0.
Consider the continuous-time $M/M/n$ system with the arrival rate $\lambda_{n}=n-n^{1-\alpha}$, the service rate $1$ and the stationary distribution be $\nu_{n}$. For $\alpha\in(0,1/2)$ and $n$ sufficiently large such that $\lambda_{n}\geq 3$ and $n-\lambda_{n}\geq 1$, we have that:
(85) $\frac{\nu(q)}{\sum_{q\in K}\nu(q)}\geq\frac{\lambda_{n}^{2}}{4(\lambda_{n}+1)^{2}}\,\forall q\in K.$
Proof.
For convenience, denote $s=\lfloor\lambda_{n}\rfloor-1,t=\lfloor\lambda_{n}\rfloor,u=\lceil\lambda_{n}\rceil,v=\lceil\lambda_{n}\rceil+1$. Observe that as $\lambda_{n}\geq 3$ and $n-\lambda_{n}\geq 1$ by assumption, we have $s+2\geq\lambda_{n}\geq 3\Rightarrow s\geq 1$ and we have $K=\{\lfloor\lambda_{n}\rfloor-1,\lfloor\lambda_{n}\rfloor,\lceil\lambda_{n}\rceil,\lceil\lambda_{n}\rceil+1\}\subset\{0,...,n\}$. And so, we have
(86) $\nu_{n}(q)=\nu_{n}(0)\frac{\lambda_{n}^{q}}{q!}.$
From here, we will bound $\nu_{n}(q)$ with respect to $\sum_{q\in K}\nu_{n}(q)$ as follows. From $\lambda_{n}-1\leq s+1\leq\lambda_{n}$ and $\lambda_{n}\geq 3$, we have
(87) $\nu_{n}(s)\leq\nu_{n}(t)$ $=\nu_{n}(s)\frac{\lambda_{n}}{s+1}\leq\nu_{n}(s)\frac{\lambda_{n}}{\lambda_{n}-1}.$
Similarly, we have from $\lambda_{n}\geq 3,\lambda_{n}-1\geq s=\lfloor\lambda_{n}\rfloor-1\geq\lambda_{n}-2$ that
(88) $\frac{\lambda_{n}}{\lambda_{n}+1}\nu_{n}(s)\leq\nu_{n}(u)$ $=\nu_{n}(s)\frac{\lambda_{n}^{2}}{(s+1)(s+2)}\leq\nu_{n}(s)\frac{\lambda_{n}}{\lambda_{n}-1}$
and since $\lambda_{n}-1\leq s+1=\lfloor\lambda_{n}\rfloor\leq\lambda_{n}$, we have
(89) $\frac{\lambda_{n}^{2}}{(\lambda_{n}+1)(\lambda_{n}+2)}\nu_{n}(s)\leq\nu_{n}(v)$ $=\frac{\lambda_{n}^{3}}{(s+1)(s+2)(s+3)}\nu_{n}(s)\leq\frac{\lambda_{n}^{2}}{(\lambda_{n}-1)(\lambda_{n}+1)}\nu_{n}(s)$
and since $\lfloor\lambda_{n}\rfloor+1\geq\lambda_{n}$, we have
(90) $\frac{\lambda_{n}}{\lambda_{n}+1}\nu_{n}(t)\leq\nu_{n}(u)$ $=\frac{\lambda_{n}}{t+1}\nu_{n}(t)\leq\nu_{n}(t)$
and since $\lceil\lambda_{n}\rceil+1\geq\lambda_{n}$, we have
(91) $\frac{\lambda_{n}}{\lambda_{n}+2}\nu_{n}(u)$ $\leq\nu_{n}(v)=\frac{\lambda_{n}}{u+1}\nu_{n}(u)\leq\nu_{n}(u).$
Finally, we have from $\lambda_{n}+1\geq t+1=\lfloor\lambda_{n}\rfloor+1\geq\lambda_{n}$ that
(92) $\frac{\lambda_{n}^{2}}{(\lambda_{n}+1)(\lambda_{n}+2)}\nu_{n}(t)$ $\leq\nu_{n}(v)=\frac{\lambda_{n}^{2}}{(t+1)(t+2)}\nu_{n}(t)\leq\nu_{n}(t)$
From (87), (88), (89), (90), (91), (92), we have
(93) $\sum_{q\in K}\nu_{n}(q)$ $\leq\left(1+\frac{\lambda_{n}}{\lambda_{n}-1}+\frac{\lambda_{n}}{\lambda_{n}-1}+\frac{\lambda_{n}^{2}}{\lambda_{n}^{2}-1}\right)\nu_{n}(s)=\frac{4\lambda_{n}^{2}+2\lambda_{n}-1}{\lambda_{n}^{2}-1}\nu_{n}(s)$
(94) $\sum_{q\in K}\nu_{n}(q)$ $\leq\left(1+1+1+1\right)\nu_{n}(t)=4\nu_{n}(t)$
(95) $\sum_{q\in K}\nu_{n}(q)$ $\leq\left(\frac{\lambda_{n}+1}{\lambda_{n}}+\frac{\lambda_{n}+1}{\lambda_{n}}+1+1\right)\nu_{n}(u)=\frac{4\lambda_{n}+2}{\lambda_{n}}\nu_{n}(u)$
(96) $\sum_{q\in K}\nu_{n}(q)$ $\leq\left(\frac{(\lambda_{n}+1)(\lambda_{n}+2)}{\lambda_{n}^{2}}+\frac{(\lambda_{n}+1)(\lambda_{n}+2)}{\lambda_{n}^{2}}+\frac{\lambda_{n}+2}{\lambda_{n}}+1\right)\nu_{n}(v)=\frac{4(\lambda_{n}+1)^{2}}{\lambda_{n}^{2}}\nu_{n}(v)$
Since $\frac{4(\lambda_{n}+1)^{2}}{\lambda_{n}^{2}}=\max\left\{4,\frac{4\lambda_{n}^{2}+2\lambda_{n}-1}{\lambda_{n}^{2}-1},\frac{4\lambda_{n}+2}{\lambda_{n}},\frac{4(\lambda_{n}+1)^{2}}{\lambda_{n}^{2}}\right\}$ for $\lambda_{n}\geq 3$, we have
(97) $\frac{\nu(q)}{\sum_{q\in K}\nu(q)}\geq\frac{\lambda_{n}^{2}}{4(\lambda_{n}+1)^{2}}\,\forall q\in K.$
Hence proved. ∎
Now that we have obtained all necessary ingredients, we proceed to prove the final piece of Theorem 3 as follows.
Theorem 0 (Restatement of Theorem 3 for $\alpha\in(0,1/2)$).
Let $\pi_{n,t}$ be the queue length distribution at time $t$ of the continuous-time $M/M/n$ system with the arrival rate $\lambda_{n}=n-n^{1-\alpha}$, the service rate $1$ and the stationary distribution be $\nu_{n}$. For $\alpha\in(0,1/2)$ and $n$ sufficiently large such that $\lambda_{n}\geq 3$, we have that:
$\chi\left(\pi_{n,t},\nu_{n}\right)\leq e^{-D_{n}t}\chi\left(\pi_{n,0},\nu_{n}\right)\,\forall t\geq 0$
where $D_{n}$ is some positive parameter such that $\lim_{n\rightarrow\infty}D_{n}=1/25$.
Proof.
Our proof consists of two parts: Proving $D_{n}$ is bounded below and showing $\lim_{n\rightarrow\infty}D_{n}=1/25$. We will show that the Poincaré constant is bounded above by some universal constant, which will imply the mixing rate of the $M/M/n$ in the Sub-Halfin-Whitt regime is bounded below by some constant as well. Since $n\geq 2,\gamma_{n}>0$ and $\lim_{n\rightarrow\infty}\gamma_{n}=1$ from Lemma 3, we have that $\gamma_{n}\geq d_{\gamma}>0$ for some positive constant $d_{\gamma}$ dependent on the system parameter $\alpha$ for all positive integer $n\geq 2$. If we have $\lambda_{n}\in\mathbb{Z}$ then $D_{n}=\gamma_{n}\geq d_{\gamma}$ from Proposition 7 where $d_{\gamma}$ is some constant and we are done.
Otherwise, if $\lambda_{n}\not\in\mathbb{Z}$ then we have to handle the positive drift term via the weighted-Poincaré inequality. Recall that from Lemma 3, we have $b\leq 2\lambda_{n}$. From here, we have for a sufficiently large $n$ that
$\sum_{q\in\mathcal{S}}b\nu_{n}(q)(f(q)-m)^{2}$ $\leq 2\lambda_{n}\sum_{q\in K}\nu_{n}(q)(f(q)-m)^{2}.$
Let $K=\{\nu_{n}(\lfloor\lambda_{n}\rfloor-1,\nu_{n}(\lfloor\lambda_{n}\rfloor,\lceil\lambda_{n}\rceil,\lceil\lambda_{n}\rceil+1\}$ and choose $m=\frac{\sum_{q\in K}\nu_{n}(q)f(q)}{\sum_{q\in K}\nu_{n}(q)}$, we have
(98) $\sum_{q\in K}\nu_{n}(q)(f(q)-m)^{2}=\left(\sum_{q\in K}\nu_{n}(q)\right)\frac{\sum_{q\in K}\nu_{n}(q)(f(q)-m)^{2}}{\sum_{q\in K}\nu_{n}(q)}$
and
$\frac{\sum_{q\in K}\nu_{n}(q)(f(q)-m)^{2}}{\sum_{q\in K}\nu_{n}(q)}$ $=\frac{\sum_{q_{1}\in K}\nu_{n}(q_{1})\left(\sum_{q_{2}\in K}(f(q_{1})-f(q_{2}))\sqrt{\nu_{n}(q_{2})}\sqrt{\nu_{n}(q_{2})}\right)^{2}}{\left(\sum_{q\in K}\nu_{n}(q)\right)^{3}}$
$\leq\frac{\sum_{q_{1},q_{2}\in K}\nu_{n}(q_{1})\nu_{n}(q_{2})(f(q_{1})-f(q_{2}))^{2}}{\left(\sum_{q\in K}\nu_{n}(q)\right)^{2}}$
$\leq 3\frac{\sum_{q_{1},q_{2}\in K}\nu_{n}(q_{1})\nu_{n}(q_{2})\sum_{k=\lfloor\lambda_{n}\rfloor-1}^{\lceil\lambda_{n}\rceil}(f(k+1)-f(k))^{2}}{\left(\sum_{q\in K}\nu_{n}(q)\right)^{2}}$
$=3\sum_{k=\lfloor\lambda_{n}\rfloor-1}^{\lceil\lambda_{n}\rceil}(f(k+1)-f(k))^{2}.$
This gives us
$\sum_{q\in K}\nu_{n}(q)(f(q)-m)^{2}$ $\leq\frac{3}{\lambda_{n}}\left(\sum_{q\in K}\nu_{n}(q)\right)\sum_{k=\lfloor\lambda_{n}\rfloor-1}^{\lceil\lambda_{n}\rceil}\lambda_{n}(f(k)-f(k+1))^{2}.$
From Lemma 1, we have
$\sum_{q\in K}\nu_{n}(q)(f(q)-m)^{2}$ $\leq\frac{3}{\lambda_{n}}\times\frac{4(\lambda_{n}+1)^{2}}{\lambda_{n}^{2}}\sum_{k=\lfloor\lambda_{n}\rfloor-1}^{\lceil\lambda_{n}\rceil}\lambda_{n}\nu_{n}(k)(f(k+1)-f(k))^{2}$
$\leq\frac{12(\lambda_{n}+1)^{2}}{\lambda_{n}^{3}}\sum_{k=0}^{\infty}\lambda_{n}\nu_{n}(k)(f(k)-f(k+1))^{2}$
(99) $=\frac{12(\lambda_{n}+1)^{2}}{\lambda_{n}^{3}}\mathcal{E}(f,f),$
where the last inequality follows from $\lambda_{n}\nu_{n}(k)(f(k)-f(k+1))^{2}\geq 0\forall k\geq 0$. And so, this implies $C_{L}=\frac{12(\lambda_{n}+1)^{2}}{\lambda_{n}^{3}}$. From here, Lemma 3 and the fact that $\lambda_{n}\geq 3$ and there exists a constant $d_{\gamma}$ such that $\gamma_{n}\geq d_{\gamma}\forall n$, we can apply Corollary 2 and obtain that the system admits Poincaré constant $C_{P}(n)=\frac{1+\left(2\lambda_{n}\right)\frac{12(\lambda_{n}+1)^{2}}{\lambda_{n}^{3}}}{\gamma_{n}}=\frac{1+\frac{24(\lambda_{n}+1)^{2}}{\lambda_{n}^{2}}}{\gamma_{n}}\leq\frac{131}{3\gamma_{n}}\leq\frac{131}{3d_{\gamma}}\,\forall n$ such that $\lambda_{n}\geq 3$.
Thus, from Proposition 6, we have in both cases that:
(100) $\chi(\pi_{t},\pi)\leq e^{-D_{n}t}\chi(\pi_{0},\pi)\,\forall t\geq 0$
for some constant $D_{n}\geq d=\frac{3}{131}\times d_{\gamma}$ where $d_{\gamma}=\min_{n}\gamma_{n}$. Moreover, since $\lim_{n\rightarrow\infty}\gamma_{n}=\lim_{n\rightarrow\infty}1-\frac{n^{2\alpha-1}}{1+n^{\alpha-1}}=1$ and $\lim_{n\rightarrow\infty}\frac{(\lambda_{n}+1)^{2}}{\lambda_{n}^{2}}=1$, we also have
(101) $\lim_{n\rightarrow\infty}C_{P}(n)=\frac{1+\frac{24(\lambda_{n}+1)^{2}}{\lambda_{n}^{2}}}{\gamma_{n}}=25.$
Which gives $\lim_{n\rightarrow\infty}D_{n}=\frac{1}{25}$. Hence, we are done.
Remark on the approach: A natural question to ask is what convergence rate we would obtain if we use the same drift analysis and the same approach as in the Super-Halfin-Whitt regime. A back of envelope calculation shows that we would get a mixing rate of $\frac{O\left(n^{1-2\alpha}\right)}{1+O\left(n^{2-4\alpha}\right)+O\left(n^{3-6\alpha}\right)}$ which would vanish to 00 as $n$ goes to infinity for $\alpha\in(0,1/2)$. And so, such an approach would not work for this regime.
Remark on the convergence results: Despite our best efforts, the reason that we cannot achieve constant $1$ is possibly because the Lyapunov-Poincaré framework does not give a tight enough bound when the set $K$ is not a singleton and we do not always get a singleton as the set that does not have a negative drift, especially when there is discreteness in the jump. Moreover, note that the eigenfunction of the generator of the $M/M/\infty$ system is the function $V(q)=q-\lambda$, which suggests that this should have been our choice of Lyapunov function. However, this function does not satisfy the $V\geq 1$ condition in Theorem 1 or Corollary 2, rendering this function ineligible for the Lyapunov-Poincaré method. Thus, we have to choose a suboptimal Lyapunov function and so it is expected to get a suboptimal convergence rate.
B.3.3. Proof of Proposition 5
To prove this Proposition, we first re-establish the negative drift result with a slightly loosened condition on the traffic but with the arrival rates being integers as follows.
Lemma 0.
Let $\{\lambda_{k}\}$ be a sequence of integral arrival rate such that $\lambda_{n}\in\mathbb{Z},n-\lambda_{n}\geq 1$ and
(102) $\lim_{n\rightarrow\infty}\frac{\log\left(n-\lambda_{n}\right)}{\log n}=1-\alpha$
where $\alpha\in(0,1/2)$ and let $V(q)$ be the Lyapunov function of the $M/M/n$ system in the sub-Halfin-Whitt regime such that
(103) $V(q)=\begin{cases}|q-\lambda_{n}|\,&\forall q\leq n,\\ \zeta e^{\theta\left(q-\lambda_{n}\right)}\,&\forall q>n,\end{cases}$
where $\zeta$ is chosen such that $|n-\lambda_{n}|=\zeta e^{\theta\left(n-\lambda_{n}\right)}$. Then, there exists a positive constant $\gamma_{n},b$ such that:
$\mathcal{L}V\leq-\gamma_{n}V+b1_{\{\lambda_{n}\}}$
where $b\leq 2\lambda_{n}$ and $\gamma_{n}\rightarrow 1$.
Proof.
Here, we follow a similar analysis as in the $\lambda_{n}\in\mathbb{Z}$ case of the proof of Lemma 3. Denote $\alpha_{n}=1-\frac{\log\left(n-\lambda_{n}\right)}{\log n}$, from Equation (102), we have $n-\lambda_{n}=n^{1-\alpha_{n}}$ where
(104) $\lim_{n\rightarrow\infty}\alpha_{n}=\alpha\in(0,1/2).$
We have that $V(q)\geq 1\,\forall q\leq n,q\neq\lambda_{n}$. For $q>n$, we choose $V(q)=\zeta e^{\theta\left(q-\lambda_{n}\right)}$ where $\zeta$ is a constant satisfies $|n-\lambda_{n}|=\zeta e^{\theta\left(n-\lambda_{n}\right)}$. Now, choose $\theta=\log\left(1+n^{\alpha_{n}-1}\right)$ such that $\delta>0$. Here, the term $\delta>0$ is added to ensure that the drift rate approaches to $1$ as in the proof of Lemma 3. Since $\theta>0$, we have that $V(q)\geq V(n)\geq 1\forall q\geq n$ and $V(q)\geq 1\forall q\leq n,q\neq\lambda_{n}$.
For $q<n$ and $q\neq\lambda_{n}$, we have:
$\mathcal{L}V$ $=\lambda_{n}|q+1-\lambda_{n}|+q|q-1-\lambda_{n}|-(\lambda_{n}+q)|q-\lambda_{n}|$
$=-|q-\lambda_{n}|=-V.$
For $q=\lambda_{n}$, we have $V(\lambda_{n})=V(\lambda_{n}-1)=V(\lambda_{n}+1)=1$. And so
$\mathcal{L}V$ $=\lambda_{n}+\lambda_{n}=2\lambda_{n}-V$
For $q>n$, we have:
$\mathcal{L}V$ $=\left(e^{\theta}-1\right)\left(\lambda_{n}-\frac{n}{e^{\theta}}\right)V$
$=n^{\alpha_{n}-1}\left(\lambda_{n}-\frac{n}{1+n^{\alpha_{n}-1}}\right)V$
$=\left(\frac{n^{2\alpha_{n}-1}}{1+n^{\alpha_{n}-1}}-1\right)V$
For $q=n$, we have:
$\mathcal{L}V(n)$ $=\lambda_{n}\zeta e^{\theta(n+1-\lambda_{n})}+n|n-1-\lambda_{n}|-(\lambda_{n}+n)|n-\lambda_{n}|$
$=\lambda_{n}\left(e^{\theta}|n-\lambda_{n}|-|n+1-\lambda_{n}|\right)-|n-\lambda_{n}|$
$=\left(\lambda_{n}\left(e^{\theta}-1\right)-\frac{\lambda_{n}}{n-\lambda_{n}}-1\right)V(n)$
$=-V(n).$
And so, we have that:
$\mathcal{L}V\leq-\gamma_{n}V+b1_{\{\lambda_{n}\}}$
where $b\leq 2\lambda_{n}$ and $\gamma_{n}=1-\frac{n^{2\alpha_{n}-1}}{1+n^{\alpha_{n}-1}}$. And so, we have
(105) $\lim_{n\rightarrow\infty}\gamma_{n}=\lim_{n\rightarrow\infty}1-\frac{n^{2\alpha_{n}-1}}{1+n^{\alpha_{n}-1}}=1$
since $\lim_{n\rightarrow\infty}\alpha_{n}=\alpha\in(0,1/2)$. ∎
Proposition 0 (Restatement of Proposition 5).
Let $\pi_{n,t},\nu_{n}$ be the queue length distribution at time $t$ and the steady state distribution of the continuous-time $M/M/n$ system with unit service rate respectively and let $\{\lambda_{n}\}$ be a sequence of integer arrival rates ($\lambda_{n}\in\mathbb{Z}$) such that $\lambda_{n}\geq 0,n-\lambda_{n}\geq 1$ and
(106) $\lim_{n\rightarrow\infty}\frac{\log\left(n-\lambda_{n}\right)}{\log n}=1-\alpha$
where $\alpha<1/2$. Then, we have
(107) $\chi(\pi_{n,t},\nu_{n})\leq e^{-\overline{D}_{n}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0$
such that $\overline{D}_{n}>0$ and $\lim_{n\rightarrow\infty}\overline{D}_{n}=1$.
Proof.
From Lemma 2, we have that there exists a Lyapunov function $V$ such that
$\mathcal{L}V\leq-\gamma_{n}+b1_{\{\lambda_{n}\}}$
where $\gamma_{n}=1-\frac{n^{2\alpha_{n}-1}}{1+n^{\alpha_{n}-1}}$. From Proposition 7, we have that the system admits the Poincaré constant $\frac{1}{\gamma_{n}}$. On the other hand, note that $\lim_{n\rightarrow\infty}\gamma_{n}=1$ from Lemma 2. And so, from Proposition 6, we have that
$\chi(\pi_{n,t},\nu_{n})\leq e^{-\overline{D}_{n}t}\chi(\pi_{n,0},\nu_{n})$
such that $\overline{D}_{n}>0$ and $\lim_{n\rightarrow\infty}\overline{D}_{n}=1$. ∎
B.4. Proof of Mean Field results
Here, we will provide a detail analysis of the Mean Field regime (when $\lambda_{n}/n\rightarrow 0$, we call this the Light Traffic regime). First, we will redo the drift analysis in Appendix B.4.1. Next, since we redo the drift analysis, we have a different finite set and so we do another local mixing analysis in Appendix B.4.2 using a different method called the truncation method. Finally, we put these results together in Appendix B.4.3.
B.4.1. Mean Field regime drift analysis
It is evident that our previous attempt in the Sub-Halfin-Whitt regime that choosing $V(q)=|q-\lambda_{n}|$ for $q\leq n$ will be problematic whenever $\lambda_{n}\not\in\mathbb{Z}$ since the bounded set that does not have the negative drift is no longer a singleton. And so, we redo the drift lemma as follows.
Lemma 0.
Let $V(q)=e^{\theta(q-n)}$ for $q\geq n$ and $V(q)=1$ for $q<n$ and $\theta\in\left(0,\log\frac{n}{\lambda_{n}}\right)$ and let $\mathcal{L}$ be the generator of the $M/M/n$ system with arrival rate $\lambda_{n}$ and unit service rate, we have
(108) $\mathcal{L}V(q)\leq-\gamma V(q)+B1_{K}$
where $\gamma=\left(e^{\theta}-1\right)\left(\frac{n}{e^{\theta}}-\lambda_{n}\right),B=\gamma+\lambda_{n}\left(e^{\theta}-1\right),K=\{0,1,...,n\}$.
Proof.
For $q>n$, we have
$\mathcal{L}V(q)=-\underbrace{\left(e^{\theta}-1\right)\left(\frac{n}{e^{\theta}}-\lambda_{n}\right)}_{=\gamma}V(q)=-\gamma V(q).$
For $q<n$, we have
$\mathcal{L}V(q)=0=-\gamma\underbrace{V(q)}_{=1}+\gamma.$
For $q=n$, we have
$\mathcal{L}V(n)=\lambda_{n}\left(e^{\theta}-1\right)=-\gamma V(n)+\gamma+\lambda_{n}\left(e^{\theta}-1\right).$
And so, we have
(109) $\mathcal{L}V(q)\leq-\gamma V(q)+B1_{K}$
where $B=\gamma+\lambda_{n}\left(e^{\theta}-1\right)$. ∎
Remark: This lemma provides a more general drift analysis for $V(q)=e^{\theta(q-n)}$, where we have the drift rate is
$\gamma=\left(e^{\theta}-1\right)\left(\frac{n}{e^{\theta}}-\lambda_{n}\right).$
Since $\theta\in\left(0,\log\frac{n}{\lambda_{n}}\right)$, we have $\gamma>0$, and so we have a negative drift outside of $K$ for this choice of $\theta$.
B.4.2. Local mixing analysis
Observe that the stationary distribution of $M/M/n$ truncated at $n$ is a truncated Poisson distribution. And so, one can show that if the original distribution admits some Poincaré constant $C_{P}$ then the truncated distribution also admits that Poincaré constant. We prove this in the following Lemma 4.
Lemma 0 (Restatement of Lemma 4).
Let $K=\{m,...,M\}$ be a connected subset of the state space $\mathcal{S}\subset\mathbb{Z}$ of the birth-and-death process, $\nu$ be the stationary distribution of the CTMC and $\nu_{K}(x)=\nu(x)/\sum_{x\in K}\nu(x)\,\forall x\in\mathcal{S}$. Furthermore, let $\mathcal{L}$ be the generator of the birth-and-death process and assume that $\operatorname{Var}_{\nu}(f)\leq C_{P}\langle f,-\mathcal{L}f\rangle_{\nu}\,\forall f\in\ell_{2,\nu}$, we have
(110) $\operatorname{Var}_{\nu_{K}}(f)\leq C_{P}\sum_{x:x,x+1\in K}Q_{K}(x,x+1)\left(f(x)-f(x+1)\right)^{2}\,\forall f\in\ell_{2,\nu_{K}}$
where $Q_{K}(x,y)=\nu_{K}(x)\mathcal{L}(x,y)\,\forall x,y\in\mathcal{S}$.
The proof of this lemma is adapted from (Roustant et al., 2016) where it is proved for continuous state space. Here, we prove it for the countable space CTMC.
Proof of Lemma 4.
Since $\mathcal{L}$ is the generator of a birth-and-death process, we have
(111) $\mathcal{L}(x,y)=0\,\forall x,y\in\mathcal{S}\text{ such that }|x-y|>1.$
Let $f\in\ell_{2,\nu_{K}}$ and let $\overline{f}$ be the extension of $f$ such that $\overline{f}(x)=f(m)\forall x\leq m$ and $\overline{f}(x)=f(M)\forall x\geq M$, we have $\overline{f}\in\ell_{2,\nu}$. Denote $Q_{K}(x,y)=\nu_{K}(x)\mathcal{L}(x,y)$, observe that
$\operatorname{Var}_{\nu_{K}}(f)$ $\leq\frac{\operatorname{Var}_{\nu}(\overline{f})}{\nu(K)}\text{ by direct algebraic manipulation}$
[…]
$=C_{P}\sum_{x\in\mathcal{S}}Q_{K}(x,x+1)\left(\overline{f}(x)-\overline{f}(x+1)\right)^{2}$
$=C_{P}\sum_{x:x,x+1\in K}Q_{K}(x,x+1)\left(f(x)-f(x+1)\right)^{2}$
[…]
Now that we have established that the truncated distribution also admits the same Poincaré constant as the original distribution (but this is not necessarily the best Poincaré constant), we can easily establish the following Poincaré inequality result for the truncated Poisson distribution, which corresponds to the stationarity distribution of the $M/M/n$ system restricted to $K=\{0,1,...,n\}$.
Lemma 0.
(Truncated $M/M/n$) Let $\mathcal{L}_{n},\mathcal{L}_{\infty}$ be the generator of the $M/M/n$ and $M/M/\infty$ system, each with unit service rate, arrival rate $\lambda_{n},\lambda$ and stationary distribution $\nu_{n},\nu$ respectively. Furthermore, consider $K=\{0,1,...,n\}$ and let $\nu_{K}(x)=\nu_{n}(x)/(\sum_{x\in K}\nu_{n}(x))\sim\frac{e^{-\lambda}\lambda^{k}}{k!}$ is the steady state of the $M/M/n$ queue restricted to $K$, we have
$\operatorname{Var}_{\nu_{K}}(f)\leq\sum_{x:x,x+1\in K}\nu_{K}(x)\mathcal{L}_{n}(x,x+1)\left(f(x)-f(x+1)\right)^{2}.$
Proof.
Let $\nu$ be the Poisson distribution with mean $\lambda$, we have that $\nu$ is also the steady state distribution of the $M/M/\infty$ system with arrival rate $\lambda$ and unit service rate. From the Poisson Poincaré inequality (Chafaï, 2006; Klaassen, 1985; Kontoyiannis et al., 2005), we have
(112) $\operatorname{Var}_{\nu}(f)\leq\langle f,\mathcal{L}_{\infty}f\rangle_{\nu}.$
Applying Lemma 4 to the set $K=\{0,1,...,n\}$, we have
(113) $\operatorname{Var}_{\nu_{K}}(f)\leq\sum_{x:x,x+1\in K}\nu_{K}(x)\mathcal{L}_{\infty}(x,x+1)\left(f(x)-f(x+1)\right)^{2}.$
Now, note that the generator $\mathcal{L}_{n}$ of the $M/M/n$ system is the same as the generator $\mathcal{L}_{\infty}$ of the $M/M/\infty$ system with arrival rate $\lambda$ and unit service rate when the queue length is no more than $n$. And so, from (112) and (113), we have
(114) $\operatorname{Var}_{\nu_{K}}(f)\leq\sum_{x:x,x+1\in K}\nu_{K}(x)\mathcal{L}_{n}(x,x+1)\left(f(x)-f(x+1)\right)^{2}.$
which implies that the local Poincaré of the $M/M/n$ system on the truncated Poisson distribution $\nu_{K}$ also admits Poincaré constant $1$. Hence proved. ∎
Remark: It is well-known that the Poisson distribution admits Poincaré constant $1$ (Klaassen, 1985; Kontoyiannis et al., 2005), so this result also implies the truncated Poisson distribution also admits this Poincaré constant. While this is not necessarily the best constant, (Jhunjhunwala et al., 2023) shows that the probability of having the queue length exceeding $n$ goes to 00 as $n$ goes to infinity, so this is a good mixing approximation of the truncated $M/M/n$ system as $n$ becomes larger.
B.4.3. Proof of Proposition 4
Proposition 0 (Restatement of Proposition 4).
Let $\pi_{n,t},\nu_{n}$ be the queue length distribution at time $t$ and the steady state distribution of the continuous-time $M/M/n$ system with unit service rate respectively and let $\lambda_{n}$ be a sequence of arrival rate such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c$ where $c\in[0,1)$. We have
$\chi(\pi_{n,t},\nu_{n})\leq e^{-L_{n}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0$
for some positive parameter $L_{n}$ such that $\lim_{n\rightarrow\infty}L_{n}=1-\sqrt{c}$.
Proof.
Choose $\theta=\log\sqrt{\frac{n}{\lambda_{n}}}$ and denote $\nu_{K}(x)=\frac{\nu_{n}(x)}{\sum_{q\in K}\nu_{n}(q)}$, we have from Lemma 3 that the system admits the drift
(115) $\mathcal{L}V\leq-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V+b1_{K}$
where $b=(\sqrt{n}-\sqrt{\lambda_{n}})^{2}+\lambda_{n}\left(\sqrt{\frac{n}{\lambda_{n}}}-1\right)$ and $K=\{0,1,...,n\}$. From Lemma 4, we have that
(116) $\operatorname{Var}_{\nu_{K}}(f)$ $\leq\sum_{x:x,x+1\in K}\nu_{K}(x)\mathcal{L}_{n}(x,x+1)\left(f(x)-f(x+1)\right)^{2}$
(117) $=\frac{1}{\nu_{n}(K)}\sum_{x:x,x+1\in K}\nu_{n}(x)\mathcal{L}_{n}(x,x+1)\left(f(x)-f(x+1)\right)^{2}$
(118) $\leq\frac{1}{\nu_{n}(K)}\sum_{x=0}^{\infty}\nu_{n}(x)\mathcal{L}_{n}(x,x+1)\left(f(x)-f(x+1)\right)^{2}$
(119) $=\frac{1}{\nu_{n}(K)}\langle f,-\mathcal{L}f\rangle_{\nu_{n}}$
where $m=\frac{\sum_{q\in K}\nu_{n}(q)f(q)}{\nu_{n}(K)}$. Now that we have established the drift in Lemma 3 and showed that Assumption 4 holds for constant $B=(\sqrt{n}-\sqrt{\lambda_{n}})^{2}+\lambda_{n}\left(\sqrt{\frac{n}{\lambda_{n}}}-1\right)$ and $C_{L}=1$, from Corollary 2, we have the system admits the Poincaré inequality
(120) $\operatorname{Var}_{\nu_{n}}\leq C_{P}(n)\langle f,-\mathcal{L}f\rangle_{\nu_{n}}$
where
(121) $C_{P}(n)=\frac{1+B}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}\leq\frac{1+(\sqrt{n}-\sqrt{\lambda_{n}})^{2}+\lambda_{n}\left(\sqrt{\frac{n}{\lambda_{n}}}-1\right)}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}=1+\frac{1}{\sqrt{\frac{n}{\lambda_{n}}}-1}+\frac{1}{n\left(1-\sqrt{\frac{\lambda_{n}}{n}}\right)^{2}}.$
Note that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c\in[0,1)$, and so
(122) $\lim_{n\rightarrow\infty}C_{P}(n)=\lim_{n\rightarrow\infty}1+\frac{1}{\underbrace{\sqrt{\frac{n}{\lambda_{n}}}}_{\rightarrow\sqrt{\frac{1}{c}}}-1}+\frac{1}{n\left(1-\underbrace{\sqrt{\frac{\lambda_{n}}{n}}}_{\rightarrow\sqrt{c}}\right)^{2}}=\frac{1}{1-\sqrt{c}}.$
Let $L_{n}=\frac{1}{C_{P}(n)}$, we have from Proposition 6 that
(123) $\chi(\pi_{n,t},\nu_{n})\leq e^{-L_{n}t}\chi(\pi_{n,0},\nu_{n})$
such that $L_{n}>0$ and $\lim_{n\rightarrow\infty}L_{n}=1-\sqrt{c}$. ∎
B.5. Proof of Proposition 3
Here, we present the proof of $M/M/\infty$ convergence bound. As mentioned in the proof sketch (see Subsection 4.2.1), we want to choose $V(q)=|q-\lambda|$ as the Lyapunov function. Unfortunately, a direct application can only yield the right mixing bound when $\lambda\in\mathbb{Z}$, which will be shown as follows.
Proposition 0 (Restatement of Proposition 3).
Let $\pi_{t,\infty}$ be the queue length distribution at time $t$ of the continuous-time $M/M/\infty$ system with the arrival rate $\lambda\in\mathbb{Z}^{+}$ and the service rate $\mu$ such that $\lambda<\mu$ and let the stationary distribution be $\nu_{\infty}$, we have:
$\chi(\pi_{t,\infty},\nu_{\infty})\leq e^{-\mu t}\chi(\pi_{0,\infty},\nu_{\infty})\,\forall t\geq 0.$
Proof.
Without the loss of generality, we assume that $\mu=1$. We will show that the Poincaré constant of the $M/M/\infty$ system is $1$. To this end, we will first analyze the drift of the system. Considering the Lyapunov function $V(q)=|q-\lambda|$, we want to show that $\mathcal{L}V\leq-V$ whenever $V$ takes some value outside of the singleton set $K=\{\lambda\}$ since $\lambda\in\mathbb{Z}^{+}$.
Indeed, when $q\not\in K$, we have that $|q-\lambda|\geq 1$ which means that $q-\lambda+1,q-\lambda,q-\lambda-1$ all have the same signs. We have:
$\mathcal{L}V(q)$ $=\lambda|q+1-\lambda|+q|q-1-\lambda|-(\lambda+q)|q-\lambda|$
$=-|q-\lambda|=-V(q).$
For $q\in K=\{\lambda\}$, we have:
$\mathcal{L}V(\lambda)$ $=\lambda|\lambda+1-\lambda|+\lambda|\lambda-1-\lambda|-(\lambda+\lambda)|\lambda-\lambda|$
$=2\lambda.$
Thus, we have shown that $M/M/\infty$ satisfies a Foster-Lyapunov condition with rate $-1$ when $\lambda$ is an integer. And so, we have that whenever $\lambda\in\mathbb{Z}$, we have a negative drift with rate $1$ outside the singleton set $K=\{\lambda\}$. From Proposition 6 and Proposition 7, the following holds for $\lambda\in\mathbb{Z}$
$\chi(\pi_{t,\infty},\nu_{\infty})\leq e^{-t}\chi(\pi_{0,\infty},\nu_{\infty}).$
Remark: The $e^{-\mu t}$ convergence rate matches the transient solution of $M/M/\infty$ for $\mu$ is the service rate of the system. Previously, there are multiple mixing proofs of $M/M/\infty$, including an entropic functional inequality proof (Chafaï, 2006). Our proof is the first Lyapunov-Poincaré proof for $M/M/\infty$. However, it is unfortunate that this proof only works for $\lambda\in\mathbb{Z}$, as otherwise we will not be able to obtain the mixing rate $1$ for $M/M/\infty$ due to the finite set in this case is no longer a singleton.
Appendix C Proof of finite-time statistics results
C.1. Proof of Corollary 4
In this Subsection, we will provide the full details of the proofs of Corollary 4. From Proposition 1, we need to establish an upper bound on the variance of the stationary distribution. In particular, we establish the variance bound for the stationary distribution of $M/M/n$ as follows.
Lemma 0.
Let $\nu_{n}$ be the stationary distribution of the $M/M/n$ queue with arrival rate $\lambda=n-n^{1-\alpha}$ and unit service rate, and let $X$ be a random variable that admits $\nu_{n}$ as the distribution, we have
$\operatorname{Var}_{\nu_{n}}\left[X\right]\leq 2(n^{\alpha}+n)^{2}.$
Proof.
To prove this result, we will do algebraic manipulation on the LHS and establish an upper bound on the second moment. We have
$\operatorname{Var}_{\nu_{n}}\left[X\right]$ $=\mathbb{E}_{\nu_{n}}\left[X^{2}\right]-\mathbb{E}_{\nu_{n}}\left[X\right]^{2}\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad$
$\leq\sum_{q=0}^{n-1}\nu_{n}(0)\frac{\lambda^{q}}{q!}q^{2}+\sum_{q=n}^{\infty}\nu_{n}(0)\frac{\lambda^{n}}{n!}\left(\frac{\lambda}{n}\right)^{q-n}q^{2}$
$=\sum_{q=0}^{n-1}\nu_{n}(0)\frac{\lambda^{q}}{q!}q^{2}+\sum_{q=0}^{\infty}\nu_{n}(0)\frac{\lambda^{n}}{n!}\left(\frac{\lambda}{n}\right)^{q}(q+n)^{2}\text{ by letting $q\rightarrow q-n$ for the second term}$
$=\sum_{q=0}^{n-1}\nu_{n}(0)\frac{\lambda^{q}}{q!}q^{2}+\sum_{q=0}^{\infty}\nu_{n}(0)\frac{\lambda^{n}}{n!}\left(\frac{\lambda}{n}\right)^{q}(q^{2}+2qn+n^{2})$
where the inequality follows from the fact that $\mathbb{E}_{\nu_{n}}[X]^{2}=\nu_{n}(0)^{2}\left(\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}q+\frac{\lambda^{n}}{n!}n^{\alpha}\left(n^{\alpha}+n\right)\right)^{2}\geq 0$. Moreover, we have
(124) $\operatorname{Var}_{\nu_{n}}\left[X\right]$ $\leq\sum_{q=0}^{n-1}\nu_{n}(0)\frac{\lambda^{q}}{q!}q^{2}+\nu_{n}(0)\frac{\lambda^{n}}{n!}\left(\frac{\frac{\lambda}{n}\left(\frac{\lambda}{n}+1\right)}{\left(1-\frac{\lambda}{n}\right)^{3}}+\frac{2n}{\left(1-\frac{\lambda}{n}\right)^{2}}+\frac{n^{2}}{1-\frac{\lambda}{n}}\right)\qquad\qquad\qquad\qquad\qquad\qquad$
(125) $=\sum_{q=0}^{n-1}\nu_{n}(0)\frac{\lambda^{q}}{q!}q^{2}+\nu_{n}(0)\frac{\lambda^{n}}{n!}n^{\alpha}\left(\frac{\lambda}{n}\left(\frac{\lambda}{n}+1\right)n^{2\alpha}+2n^{1+\alpha}+n^{2}\right).$
From the fact that $\nu_{n}(0)=\left[\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}q^{2}+\frac{\lambda^{n}}{n!}n^{\alpha}\right]^{-1}$, we have the bound
(126) $\sum_{q=0}^{n-1}\nu_{n}(0)\frac{\lambda^{q}}{q!}q^{2}$ $\leq\left(\sum_{q=0}^{n-1}\nu_{n}(0)\frac{\lambda^{q}}{q!}\right)(n-1)^{2}\leq\left(\underbrace{\sum_{q=0}^{n-1}\nu_{n}(0)\frac{\lambda^{q}}{q!}+\sum_{q=n}^{\infty}\nu_{n}(0)\frac{\lambda^{n}}{n!}\left(\frac{\lambda}{n}\right)^{q-n}}_{=1}\right)(n-1)^{2}<n^{2}.$
Furthermore, we have
(127) $\nu_{n}(0)\frac{\lambda^{n}}{n!}n^{\alpha}\left(\frac{\lambda}{n}\left(\frac{\lambda}{n}+1\right)n^{2\alpha}+2n^{1+\alpha}+n^{2}\right)\leq 2n^{2\alpha}+2n^{1+\alpha}+n^{2}$
from the fact that
(128) $\nu_{n}(0)\frac{\lambda^{n}}{n!}n^{\alpha}=\nu_{n}(0)\frac{\lambda^{n}}{n!}\frac{1}{1-\frac{\lambda}{n}}=\sum_{q=n}^{\infty}\nu_{n}(0)\frac{\lambda^{n}}{n!}\left(\frac{\lambda}{n}\right)^{q-n}=\sum_{q=n}^{\infty}\nu_{n}(q)\leq 1$
and $\frac{\lambda}{n}\leq 1$. From (125), (126) and (127), we have
(129) $\operatorname{Var}_{\nu_{n}}\left[X\right]\leq 2\left(n^{2\alpha}+n^{1+\alpha}+n^{2}\right)\leq 2(n^{\alpha}+n)^{2}.$
In essence, Lemma 1 tells us that $\operatorname{Var}_{\nu_{n}}\left[X\right]=O\left((n+n^{\alpha})^{2}\right)$, which is the same order of the squared mean queue length. While we may obtain a tighter bound with a more fine-grained analysis on $\mathbb{E}_{\nu_{n}}[X]^{2}$, this bound matches the desired order and is sufficient for our needs. For the Light Traffic regime (i.e. $\lambda_{n}/n\rightarrow 0$), as we have the arrival rate $\lambda_{n}$ grows at a much slower pace compared to $n$, and so we have the quantity $\operatorname{Var}_{\nu_{n}}\left[X\right]$ is much smaller. We have the following lemma.
Lemma 0.
Let $\nu_{n}$ be the stationary distribution of the $M/M/n$ queue with arrival rate $\lambda_{n}=n-n^{1-\alpha_{n}}$ such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=0$ and unit service rate, and let $X$ be a random variable that admits $\nu_{n}$ as the distribution, we have
$\operatorname{Var}_{\nu_{n}}\left[X\right]\leq n\,\forall n\geq n_{0}$
where $n_{0}$ is a sufficiently large positive integer.
Proof.
Similarly from Lemma 1, we have the bound
[…]
(130) $\leq(n-1)^{2}+\nu_{n}(0)\frac{\lambda_{n}^{n}}{n!}\left(\frac{\lambda_{n}}{n}\left(\frac{\lambda_{n}}{n}+1\right)+2n+n^{2}\right)\text{ from$\lambda_{n}\geq 0$}.$
Since $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=0$, there exists a positive integer $n_{1}$ such that $\frac{\lambda_{n}}{n}\leq\frac{1}{3}\,\forall n\geq n_{1}$. We have
(131) $\frac{\lambda_{n}^{n}}{n!}\leq\frac{\left(\frac{n}{3}\right)^{n}}{n!}\leq\frac{\left(\frac{n}{3}\right)^{n}}{\sqrt{2\pi n}\left(\frac{n}{e}\right)^{n}}=\frac{\left(\frac{e}{3}\right)^{n}}{\sqrt{2\pi n}}$
from Stirling’s approximation. From here, we have
(132) $\nu_{n}(0)\frac{\lambda_{n}^{n}}{n!}\left(\frac{\lambda_{n}}{n}\left(\frac{\lambda_{n}}{n}+1\right)+2n+n^{2}\right)$ $\leq\frac{\lambda_{n}^{n}}{n!}\left(\frac{\lambda_{n}}{n}\left(\frac{\lambda_{n}}{n}+1\right)+2n+n^{2}\right)\leq\left(\frac{e}{3}\right)^{n}\frac{1}{\sqrt{2\pi n}}\left(\frac{4}{9}+2n+n^{2}\right).$
Since $e<3$, we can choose a sufficiently large $n_{2}$ such that for $n\geq n_{2}$ such that
(133) $\left(\frac{e}{3}\right)^{n}\frac{1}{\sqrt{2\pi n}}\left(\frac{4}{9}+2n+n^{2}\right)\leq n.$
From (130), (133) and for a sufficiently large $n_{0}\geq\max\{n_{1},n_{2}\}$, we have
(134) $\operatorname{Var}_{\nu_{n}}\left[X\right]\leq(n-1)^{2}+n\leq n^{2}.$
Hence proved. ∎
From here, we can use the variance bounds in Lemma 1 and Lemma 2 to establish the finite-time mean queue length as follows.
Corollary 0 (Restatement of Corollary 4).
Let $\pi_{n,t}$ be the queue length distribution at time $t$ of the continuous-time $M/M/n$ system with the arrival rate $\lambda_{n}=n-n^{1-\alpha}$ and a service rate $1$ whose stationary distribution be $\nu_{n}$. For $\alpha\geq 1$, we have that
(135) $\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\sqrt{2}\left(n+n^{\alpha}\right)\chi(\pi_{n,0},\nu_{n}).$
For $\alpha\in(1/2,1)$, we have
(136) $\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\sqrt{2}\left(n+n^{\alpha}\right)\chi(\pi_{n,0},\nu_{n})$
for some $C_{n}>0$ such that $\lim_{n\rightarrow\infty}C_{n}=1$.
For $\alpha\in(0,1/2)$, we have
(137) $\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-D_{n}t}\sqrt{2}\left(n+n^{\alpha}\right)\chi(\pi_{n,0},\nu_{n})$
for some $D_{n}>0$ such that $\lim_{n\rightarrow\infty}D_{n}=1/25$.
And finally, if $\lambda_{n}$ is a sequence of arrival rates such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=0$, then we have
(138) $\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-L_{n}t}n\chi(\pi_{n,0},\nu_{n})$
for some $L_{n}>0$ such that $\lim_{n\rightarrow\infty}L_{n}=1$.
Proof.
Let $M_{n,\alpha}$ be the corresponding mixing rate of the system when the system has $n$ servers and the heavy-traffic parameter $\alpha$. From Theorem 3, we have
(139) $\chi^{2}(\pi_{n,t},\nu_{n})=\sum_{x=0}^{\infty}\frac{\left(\pi_{n,t}(x)-\nu_{n}(x)\right)^{2}}{\pi(x)}\leq e^{-2M_{n,\alpha}t}\chi^{2}(\pi_{n,0},\nu_{n}).$
From Corollary 1, we have that
(140) $\left|\mathbb{E}_{\pi_{n,t}}\left[X\right]-\mathbb{E}_{\nu_{n}}\left[X\right]\right|\leq\chi(\pi_{n,t},\nu_{n})\sqrt{\mathbb{E}_{\nu_{n}}\left[X^{2}\right]-\mathbb{E}_{\nu_{n}}\left[X\right]^{2}}$
From (139) and (140), we have
(141) $\big{|}\mathbb{E}_{\pi_{n,t}}\left[X\right]-\mathbb{E}_{\nu_{n}}\left[X\right]\big{|}\leq e^{-M_{n,\alpha}t}\sqrt{\operatorname{Var}_{\nu_{n}}\left[X\right]}\chi(\pi_{n,0},\nu_{n}).$
For $\alpha>0$ and from Lemma 1, we have
(142) $\big{|}\mathbb{E}_{\pi_{n,t}}\left[X\right]-\mathbb{E}_{\nu_{n}}\left[X\right]\big{|}\leq e^{-M_{n,\alpha}t}\sqrt{2}\left(n^{\alpha}+n\right)\chi(\pi_{n,0},\nu_{n}).$
From Theorem 3, we have three cases: $\alpha\geq 1,\alpha\in(1/2,1)$ and $\alpha\in(0,1/2)$. For $\alpha\geq 1$, we have $M_{n,\alpha}=\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}$ and so
$\big{|}\mathbb{E}_{\pi_{n,t}}\left[X\right]-\mathbb{E}_{\nu_{n}}\left[X\right]\big{|}\leq e^{-\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}t}\sqrt{2}\left(n^{\alpha}+n\right)\chi(\pi_{n,0},\nu_{n}).$
For $11/2<\alpha<$ 1, we have $M_{n,\alpha}=C_{n}\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}$ such that $\lim_{n\rightarrow\infty}C_{n}=1$ and this gives
$\big{|}\mathbb{E}_{\pi_{n,t}}\left[X\right]-\mathbb{E}_{\nu_{n}}\left[X\right]\big{|}\leq e^{-C_{n}\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}t}\sqrt{2}\left(n^{\alpha}+n\right)\chi(\pi_{n,0},\nu_{n}).$
Similarly, for $20<\alpha<1/$ 2, we have $M_{n,\alpha}=D_{n}$ such that $D_{n}>d$ for some positive universal constant $d$ and we have
$\big{|}\mathbb{E}_{\pi_{n,t}}\left[X\right]-\mathbb{E}_{\nu_{n}}\left[X\right]\big{|}\leq e^{-D_{n}t}\sqrt{2}\left(n^{\alpha}+n\right)\chi(\pi_{n,0},\nu_{n}).$
Finally, when $\lambda_{n}$ is a sequence of arrival rates such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=0$, we have from (141) and Lemma 2 that
(143) $\big{|}\mathbb{E}_{\pi_{n,t}}\left[X\right]-\mathbb{E}_{\nu_{n}}\left[X\right]\big{|}\leq e^{-L_{n}t}n\chi(\pi_{n,0},\nu_{n}).$
Remark on the dependence of $n$: While it is known that with a proper choice of the rescaling factor, we can show the existence of the MGF and consequently, the boundedness of the second moment, it would give a weak characterization on the dependence of $n$. By establishing the second moment bound of $\nu$, we have shown that the $n$ dependency is only polynomial in $n$, and thus, it would be negligible given the exponential convergence of the system.
C.2. Proof of Corollary 5
To establish tail bound results, we first need to obtain bounds on the MGF, which involves careful analysis and choice of parameters. Our MGF bounds are stated as follows.
Lemma 0.
Let $n\geq n_{0}$ and $\nu_{n}$ be the stationary distribution of the $M/M/n$ system with arrival rate $\lambda=n-n^{1-\alpha}$ and unit service rate, we have $\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right]<\infty$ for $\theta<\log\frac{n}{\lambda}$. Moreover, when $\varepsilon=1-\frac{\lambda}{n}$ and $\theta_{n}=\frac{1}{1+\delta}\log\left(\frac{n}{\lambda}\right)$, we have
(144) $\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon(q-n)}{1+\delta}}\right]\leq\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}+\frac{\delta+1}{\delta}\frac{\lambda^{n}}{n!}n^{\alpha}}{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}+\frac{\lambda^{n}}{n!}n^{\alpha}}\leq 1+\frac{1}{\delta}.$
Moreover, if $\alpha\in(0,1/2)$ then
$\lim_{n\rightarrow\infty}\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon(q-n)}{1+\delta}}\right]=0.$
If $\alpha\in\left(1/2,\infty\right)$ then
$\lim_{n\rightarrow\infty}\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon(q-n)}{1+\delta}}\right]$ $=1+\frac{1}{\delta}.$
Proof.
Let $n_{0}=\max\left\{65,2^{\frac{1}{\alpha}}\right\}$ and $\theta_{n}=\frac{1}{1+\delta}\log\left(\frac{n}{\lambda}\right)$ for $\delta\in(0,+\infty)$. First, we will establish the following bound:
$\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon(q-n)}{1+\delta}}\right]$ $\leq\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right].$
Let $n_{0}=\max\left\{65,2^{\frac{1}{\alpha}}\right\}$ and $\theta_{n}=\frac{1}{1+\delta}\log\left(\frac{n}{\lambda}\right)$ for $\delta\in(0,+\infty)$. Note that for all $n\geq n_{0}$, we have
(145) $\frac{\epsilon}{1+\delta}=\theta_{n}\left(n^{\alpha}\log\left(\frac{n}{\lambda}\right)\right)^{-1}\overset{(a)}{\leq}\theta_{n}\implies\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon(q-n)}{1+\delta}}\right]$ $\leq\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right],$
where the (a) follows from Lemma 2 as
(146) $n^{\alpha}\log\left(\frac{n}{\lambda}\right)$ $\geq n^{\alpha}\left(\frac{n^{1-\alpha}}{\lambda}-\frac{1}{2}\left(\frac{n^{1-\alpha}}{\lambda}\right)^{2}\right)$
(147) $=\frac{n}{\lambda}-\frac{n^{2-\alpha}}{2\lambda^{2}}$
(148) $=1+\frac{n^{1-\alpha}}{\lambda}-\frac{n^{2-\alpha}}{2\lambda^{2}}$
(149) $\geq 1$
where the last inequality holds since $\frac{\lambda}{n}\geq\frac{1}{2}$, which is true for all $n\geq n_{0}$. Next, we will bound the quantity $\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right]$ for a fixed $n\geq n_{0}$. Let $\nu_{n}$ be the stationary distribution of the $M/M/n$ system, we have
$\nu_{n}(0)=\left[\sum_{k=0}^{n-1}\frac{\lambda^{k}}{k!}+\frac{\lambda^{n}}{n!}\frac{1}{1-\rho}\right]^{-1}=\left[\sum_{k=0}^{n-1}\frac{\lambda^{k}}{k!}+\frac{\lambda^{n}}{n!}n^{\alpha}\right]^{-1}.$
Consider the MGF $\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right]$, we have
$\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right]=\sum_{q<n}\nu_{n}(0)\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}+\sum_{q\geq n}\nu_{n}(0)\frac{\lambda^{n}}{n!}\left(\frac{e^{\theta_{n}}\lambda}{n}\right)^{q-n}.$
Since $\theta_{n}=\frac{1}{1+\delta}\log\frac{n}{\lambda}$, we have $\frac{e^{\theta_{n}}\lambda}{n}<1$ and so $\sum_{q\geq n}\nu_{n}(0)\frac{\lambda^{n}}{n!}\left(\frac{e^{\theta_{n}}\lambda}{n}\right)^{q-n}$ is summable, which means that $\mathbb{E}\left[e^{\theta_{n}(q-n)}\right]$ exists. We have:
$\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right]$ $=\sum_{q<n}\nu_{n}(0)\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}+\sum_{q\geq n}\nu_{n}(0)\frac{\lambda^{n}}{n!}\left(\frac{e^{\theta_{n}}\lambda}{n}\right)^{q-n}$
$=\sum_{q<n}\nu_{n}(0)\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}+\nu_{n}(0)\frac{\lambda^{n}}{n!}\frac{1}{1-\frac{e^{\theta_{n}}\lambda}{n}}$
$=\sum_{q<n}\nu_{n}(0)\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}+\nu_{n}(0)\frac{\lambda^{n}}{n!}\frac{1}{1-\left(\frac{\lambda}{n}\right)^{\frac{\delta}{\delta+1}}}$
$=\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}+\frac{\lambda^{n}}{n!}\frac{1}{1-\left(\frac{\lambda}{n}\right)^{\frac{\delta}{\delta+1}}}}{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}+\frac{\lambda^{n}}{n!}n^{\alpha}}$
$=\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}+\frac{\lambda^{n}}{n!}n^{\alpha}\frac{1}{n^{\alpha}\left(1-\left(\frac{\lambda}{n}\right)^{\frac{\delta}{\delta+1}}\right)}}{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}+\frac{\lambda^{n}}{n!}n^{\alpha}}.$
From here, note that we have the bound $n^{\alpha}\left(1-\left(\frac{\lambda}{n}\right)^{\frac{\delta}{\delta+1}}\right)\geq\frac{\delta+1}{\delta}$. For more details, the reader can refer to Lemma 3 in Appendix D. Apply this bound and from (145), we have
(150) $\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon(q-n)}{1+\delta}}\right]$ $\leq\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}+\frac{\delta+1}{\delta}\frac{\lambda^{n}}{n!}n^{\alpha}}{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}+\frac{\lambda^{n}}{n!}n^{\alpha}}\leq 1+\frac{1}{\delta}.$
The last inequality follows from the fact that if $\frac{a}{b}\leq\frac{c}{d}$ then $\frac{a+c}{b+d}\leq\frac{c}{d}$ for $a,b,c,d>0$.
Now, we will study the asymptotic behavior of the MGF $\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon(q-n)}{1+\delta}}\right]$. Taking limit on both sides and also apply Lemma 3, we have
(151) $\lim_{n\rightarrow\infty}\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right]$ $=\lim_{n\rightarrow\infty}\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}+\frac{\delta+1}{\delta}\frac{\lambda^{n}}{n!}n^{\alpha}}{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}+\frac{\lambda^{n}}{n!}n^{\alpha}}.$
Here, we have to consider the different cases when $\alpha\in(0,1/2)$ and $\alpha\in(1/2,+\infty)$ since there is phase transition at the $\alpha=1/2$ regime.
Case 1 ($\alpha\in(0,1/2)$): In this case, we want to prove that
$\lim_{n\rightarrow\infty}\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right]$ $=0.$
Choosing $m=\left\lfloor n-\frac{n^{1-\alpha}}{2}\right\rfloor$, observe that for $q\geq m$, we have $\frac{\lambda^{q}}{q!}\leq\frac{\lambda^{m}}{m!}\left(\frac{\lambda}{m}\right)^{q-m}$ which implies
(152) $\sum_{m<q<n}\frac{\lambda^{q}}{q!}$ $\leq\frac{\lambda^{m}}{m!}\sum_{m<q<n}\left(\frac{\lambda}{m}\right)^{q-m}<\frac{\lambda^{m}}{m!}\frac{1}{1-\frac{\lambda}{m}}=\frac{\lambda^{m}}{m!}\frac{m}{m-\lambda}\leq\frac{\lambda^{m}}{m!}\frac{n-\frac{n^{1-\alpha}}{2}}{\frac{n^{1-\alpha}}{2}}<\frac{\lambda^{m}}{m!}2n^{\alpha}.$
And so, we can bound the MGF as follows:
$\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right]$ $\leq\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}+\frac{\delta+1}{\delta}\frac{\lambda^{n}}{n!}n^{\alpha}}{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}+\frac{\lambda^{n}}{n!}n^{\alpha}}$
$=\frac{\sum_{0\leq q\leq m}\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}+\sum_{m<q<n}\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}+\frac{\delta+1}{\delta}\frac{\lambda^{n}}{n!}n^{\alpha}}{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}+\frac{\lambda^{n}}{n!}n^{\alpha}}$
[…]
$<e^{-\theta_{n}\frac{n^{1-\alpha}}{2}}+\frac{\frac{\lambda^{m}}{m!}2n^{\alpha}}{\frac{\lambda^{m}}{m!}\frac{n^{1-\alpha}}{4}}+\frac{\frac{\delta+1}{\delta}\frac{\lambda^{n}}{n!}n^{\alpha}}{\frac{\lambda^{n}}{n!}\frac{n^{1-\alpha}}{2}}\text{ from $\theta_{n}>0$ and }\frac{\lambda^{q}}{q!}\geq\frac{\lambda^{n}}{n!}\,\forall\lambda<q\leq n$
$=e^{-\theta_{n}\frac{n^{1-\alpha}}{2}}+8n^{2\alpha-1}+\frac{\delta+1}{\delta}2n^{2\alpha-1}$
Now, we need to bound $\theta_{n}$. For $\theta_{n}=\frac{1}{1+\delta}\log\left(\frac{n}{\lambda}\right)$, we have $\theta_{n}\leq\frac{1}{1+\delta}\left(\frac{n}{\lambda}-1\right)<\frac{1}{1+\delta}\frac{n^{1-\alpha}}{\frac{7n}{8}}=\frac{8}{7(1+\delta)}n^{-\alpha}$. For $\alpha\in\left(0,\frac{1}{2}\right)$, this gives
$\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right]$ $<e^{-\frac{4n^{1-2\alpha}}{7(1+\delta)}}+8n^{2\alpha-1}+\frac{\delta+1}{\delta}2n^{2\alpha-1}.$
Taking limits on both sides, we have
$\lim_{n\rightarrow\infty}\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right]$ $<\lim_{n\rightarrow\infty}\underbrace{e^{-\frac{4n^{1-2\alpha}}{7(1+\delta)}}}_{\rightarrow 0}+\underbrace{8n^{2\alpha-1}}_{\rightarrow 0}+\frac{\delta+1}{\delta}\underbrace{2n^{2\alpha-1}}_{\rightarrow 0}=0.$
By the Sandwich Theorem and since $\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right]\geq 0$, we have $\lim_{n\rightarrow\infty}\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right]=0$.
Case 2 ($\alpha\in(1/2,\infty)$): In this case, we want to prove that
$\lim_{n\rightarrow\infty}\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right]$ $=\frac{\delta+1}{\delta}.$
Observe that for $\alpha\in\left(\frac{1}{2},\infty\right)$, we have
$\lim_{n\rightarrow\infty}\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}}{\frac{\lambda^{n}}{n!}n^{\alpha}}=0.$
Note that we have
(153) $\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}}{\frac{\lambda^{n}}{n!}n^{\alpha}}\leq\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}}{\frac{\lambda^{n}}{n!}n^{\alpha}}\overset{*}{\leq}\frac{\sum_{q=0}^{\infty}\frac{\lambda^{q}}{q!}}{\frac{\lambda^{n}}{n!}n^{\alpha}}\overset{**}{=}n!e^{\lambda}\lambda^{-n}n^{-\alpha}\leq 2\sqrt{2\pi}e^{-n^{1-\alpha}}\left(\frac{\lambda}{n}\right)^{-n}n^{\frac{1}{2}-\alpha}\leq 4\sqrt{2\pi}n^{\frac{1}{2}-\alpha},$
where ()(*)( ∗ ) and ()(**)( ∗ ∗ ) follows from
(154) $\sum_{q=0}^{n-1}\frac{e^{-\lambda}\lambda^{q}}{q!}\leq\sum_{q=0}^{\infty}\frac{e^{-\lambda}\lambda^{q}}{q!}=1.$
In addition, the second last inequality of (153) holds by Stirling’s formula and the last inequality holds as $\left(\frac{\lambda}{n}\right)^{-n}=\left(1-n^{-\alpha}\right)^{-n}\leq 2e^{n^{1-\alpha}}$ for $n\geq n_{0}$. Now, by taking the limit as $n\rightarrow\infty$ and noting that $\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}}{\frac{\lambda^{n}}{n!}n^{\alpha}}\geq 0$, we get
(155) $\lim_{n\rightarrow\infty}\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}}{\frac{\lambda^{n}}{n!}n^{\alpha}}=\lim_{n\rightarrow\infty}\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}}{\frac{\lambda^{n}}{n!}n^{\alpha}}=0.$
Furthermore, observe that $\theta_{n}>0$ and $e^{\theta_{n}(q-n)}\leq 1\,\forall 0\leq q\leq n-1$, we have
(156) $0\displaystyle 0\leq\lim_{n\rightarrow\infty}\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}}{\frac{\lambda^{n}}{n!}n^{\alpha}}\leq\lim_{n\rightarrow\infty}\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}}{\frac{\lambda^{n}}{n!}n^{\alpha}}=0\Rightarrow\lim_{n\rightarrow\infty}\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}}{\frac{\lambda^{n}}{n!}n^{\alpha}}=$ 0 .
From here, we can compute the limit as follows
$\lim_{n\rightarrow\infty}\mathbb{E}_{\nu_{n}}\left[e^{\theta_{n}(q-n)}\right]$ $=\lim_{n\rightarrow\infty}\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}+\frac{\delta+1}{\delta}\frac{\lambda^{n}}{n!}n^{\alpha}}{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}+\frac{\lambda^{n}}{n!}n^{\alpha}}$
$=\frac{\lim_{n\rightarrow\infty}\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}e^{\theta_{n}(q-n)}}{\frac{\lambda^{n}}{n!}n^{\alpha}}+\frac{\delta+1}{\delta}}{\lim_{n\rightarrow\infty}\frac{\sum_{q=0}^{n-1}\frac{\lambda^{q}}{q!}}{\frac{\lambda^{n}}{n!}n^{\alpha}}+1}$
[…]
$=\frac{\delta+1}{\delta}.$
Hence proved. ∎
For the Light Traffic regime (when $\lambda_{n}/n\rightarrow 0$), we have a separate lemma to bound the MGF in this regime. Unlike in Lemma 3, we do not need a rescaling term in front of $q-n$ in this regime.
Lemma 0.
Let $\{\lambda_{n}\}$ be a sequence of positive numbers such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=0$. Let $n\geq\max\{n_{0},20\}$ where $n_{0}$ is an integer such that $\frac{\lambda_{n}}{n}\leq\frac{1}{2e}\,\forall n\geq n_{0}$ and $\nu_{n}$ be the stationary distribution of the $M/M/n$ system with arrival rate $\lambda_{n}$ and unit service rate. We have
(157) $\mathbb{E}_{\nu_{n}}\left[e^{q-n}\right]\leq e^{-\frac{n^{1-\alpha_{n}}}{2}}+4n^{\alpha_{n}-1}+\frac{8n^{2\alpha_{n}-1}}{e}\,\forall n\geq n_{0}.$
Consequently, we have $\lim_{n\rightarrow\infty}\mathbb{E}_{\nu_{n}}\left[e^{q-n}\right]=0$.
Proof.
Since $\lambda_{n}\in(0,n)$, there exists a unique $\alpha_{n}\in\mathbb{R}^{+}$ such that $\lambda_{n}=n-n^{1-\alpha_{n}}$. Furthermore, as $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=0$, we have $\lim_{n\rightarrow\infty}n^{\alpha_{n}}=1$ and $\lim_{n\rightarrow\infty}\alpha_{n}=0$. We bound the MGF $\mathbb{E}_{\nu_{n}}\left[e^{q-n}\right]$ by following the same routine as in the $\alpha\in(0,1/2)$ regime in Lemma 3 as follows. Choosing $m=\left\lfloor n-\frac{n^{1-\alpha_{n}}}{2}\right\rfloor$, observe that for $q>m$, we have
$\frac{\lambda_{n}^{q}}{q!}\leq\frac{\lambda_{n}^{m}}{m!}\left(\frac{\lambda_{n}}{m}\right)^{q-m}\Rightarrow\sum_{m<q<n}\frac{\lambda_{n}^{q}}{q!}\leq\frac{\lambda_{n}^{m}}{m!}\sum_{m<q<n}\left(\frac{\lambda_{n}}{m}\right)^{q-m}$
From the well-known inequality $\sum_{i=1}^{k}x^{i}<\frac{1}{1-x}\forall x<1$ and the fact that $\frac{\lambda_{n}}{m}<1$, we have
(158) $\sum_{m<q<n}\frac{\lambda_{n}^{q}}{q!}$ $<\frac{\lambda_{n}^{m}}{m!}\frac{1}{1-\frac{\lambda_{n}}{m}}=\frac{\lambda_{n}^{m}}{m!}\frac{m}{m-\lambda_{n}}\leq\frac{\lambda_{n}^{m}}{m!}\frac{n-\frac{n^{1-\alpha_{n}}}{2}}{\frac{n^{1-\alpha_{n}}}{2}-1}<\frac{\lambda_{n}^{m}}{m!}2n^{\alpha_{n}}.$
Where the last inequality is equivalent to
$n-\frac{n^{1-\alpha_{n}}}{2}$ $<2n^{\alpha_{n}}\left(\frac{n^{1-\alpha_{n}}}{2}-1\right)=n-2n^{\alpha_{n}}$
$\Leftrightarrow 8$ $<n^{1-2\alpha_{n}}$
$\Leftrightarrow\frac{8}{\left(1-\frac{1}{2e}\right)^{2}}$ $<n$
which is true for all $n\geq 20$ and $\frac{\lambda_{n}}{n}\leq\frac{1}{2e}$. Moreover, we can rewrite the MGF as
$\mathbb{E}_{\nu_{n}}\left[e^{q-n}\right]$ $=\frac{\sum_{q=0}^{n-1}\frac{\lambda_{n}^{q}}{q!}e^{q-n}+\frac{\lambda_{n}^{n}}{n!}\sum_{q=n}^{\infty}\left(\frac{\lambda_{n}}{n}\right)^{q-n}e^{q-n}}{\sum_{q=0}^{n-1}\frac{\lambda_{n}^{q}}{q!}+\frac{\lambda_{n}^{n}}{n!}\frac{1}{1-\frac{\lambda_{n}}{n}}}$
$=\frac{\sum_{q=0}^{n-1}\frac{\lambda_{n}^{q}}{q!}e^{q-n}+\frac{\lambda_{n}^{n}}{n!}\frac{1}{1-\frac{e\lambda_{n}}{n}}}{\sum_{q=0}^{n-1}\frac{\lambda_{n}^{q}}{q!}+\frac{\lambda_{n}^{n}}{n!}\frac{1}{1-\frac{\lambda_{n}}{n}}}\text{ since $\sum_{i=0}^{\infty}x^{i}=\frac{1}{1-x}$ for $|x|<1$ and $\frac{\lambda_{n}}{n}\leq\frac{1}{2e}$}$
$=\frac{\sum_{0\leq q\leq m}\frac{\lambda_{n}^{q}}{q!}e^{q-n}+\sum_{m<q<n}\frac{\lambda_{n}^{q}}{q!}e^{q-n}+\frac{\lambda_{n}^{n}}{n!}\frac{1}{1-\frac{e\lambda_{n}}{n}}}{\sum_{q=0}^{n-1}\frac{\lambda_{n}^{q}}{q!}+\frac{\lambda_{n}^{n}}{n!}\frac{1}{1-\frac{\lambda_{n}}{n}}}$
And so, we can bound the MGF as follows:
[…]
$<e^{-\frac{n^{1-\alpha_{n}}}{2}}+\frac{\frac{\lambda_{n}^{m}}{m!}2n^{\alpha_{n}}}{\frac{\lambda_{n}^{m}}{m!}\frac{en^{1-\alpha_{n}}}{4}}+\frac{\frac{\lambda_{n}^{n}}{n!}\frac{1}{1-\frac{e\lambda_{n}}{n}}}{\sum_{q=0}^{n-1}\frac{\lambda_{n}^{q}}{q!}+\frac{\lambda_{n}^{n}}{n!}\frac{1}{1-\frac{\lambda_{n}}{n}}}\text{ since }\frac{\lambda_{n}^{q}}{q!}>\frac{\lambda_{n}^{m}}{m!}\forall\lambda_{n}\leq q\leq m$
$<e^{-\frac{n^{1-\alpha_{n}}}{2}}+\frac{8n^{2\alpha_{n}-1}}{e}+\frac{\frac{\lambda_{n}^{n}}{n!}\frac{1}{1-\frac{1}{2}}}{\sum_{q=0}^{n-1}\frac{\lambda_{n}^{q}}{q!}+\frac{\lambda_{n}^{n}}{n!}\frac{1}{1-\frac{\lambda_{n}}{n}}}\text{ since }\frac{\lambda_{n}}{n}\leq\frac{1}{2e}\,\forall n\geq n_{0}$
$\leq e^{-\frac{n^{1-\alpha_{n}}}{2}}+\frac{8n^{2\alpha_{n}-1}}{e}+\frac{\frac{2\lambda_{n}^{n}}{n!}}{\frac{\lambda_{n}^{n}}{n!}\frac{n^{1-\alpha_{n}}}{2}}\text{ since }\frac{\lambda_{n}^{q}}{q!}>\frac{\lambda_{n}^{n}}{n!}\forall\lambda_{n}\leq q\leq n\text{ and }n-\lambda_{n}>\frac{n^{1-\alpha_{n}}}{2}$
$=e^{-\frac{n^{1-\alpha_{n}}}{2}}+4n^{\alpha_{n}-1}+\frac{8n^{2\alpha_{n}-1}}{e}.$
By the Sandwich Theorem, we have $\lim_{n\rightarrow\infty}\mathbb{E}_{\nu_{n}}\left[e^{q-n}\right]=0$. Hence proved. ∎
From the MGF bounds in Lemma 3, we can now formally prove Corollary 5 as follows.
Corollary 0 (Restatement of Corollary 5).
Let $\pi_{n,t},\nu_{n}$ be the queue length distribution at time $t$ and the stationary distribution of the $M/M/n$ system with arrival rate $\lambda_{n}=n-n^{1-\alpha_{n}}$ and unit service rate respectively, and let $\varepsilon=1-\frac{\lambda_{n}}{n}$ and $q$ be the random variable denoting the queue length. For $n\geq n_{0},\delta\in(0,+\infty)$ and $n_{0}$ is a constant dependent on the regime, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]\leq\left(1+\chi(\pi_{n,t},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}$
In particular, for $\alpha_{n}=\alpha\in[1,\infty)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}.$
For $\alpha_{n}=\alpha\in(1/2,1)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}$
where $C_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}C_{n}=1$.
For $\alpha_{n}=\alpha\in(0,1/2)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-D_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{e^{-\frac{4n^{1-2\alpha}}{7(1+\delta)}}+\left(10+\frac{2}{\delta}\right)n^{2\alpha-1}}\times e^{-\frac{x}{2(1+\delta)}}$
where $D_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}D_{n}=1/25$.
And finally, if $\lambda_{n}$ is a sequence of arrival rates such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=0$ then we have
$\mathbb{P}_{\pi_{n,t}}\left[q-n>x\right]$ $\leq\left(1+e^{-L_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{e^{-\frac{n^{1-\alpha_{n}}}{2}}+4n^{\alpha_{n}-1}+\frac{8n^{2\alpha_{n}-1}}{e}}\times e^{-\frac{x}{2}}$
where $L_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}L_{n}=1$.
Proof.
First, let’s consider the regime $\alpha_{n}=\alpha\in(0,\infty)$. Let $n_{0}=\max\left\{65,2^{\frac{1}{\alpha}}\right\}$. From Lemma 3 and Corollary 2, we have
$\left|\mathbb{E}_{\pi_{n,t}}\left[e^{\frac{\varepsilon x}{2(1+\delta)}}\right]-\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon x}{2(1+\delta)}}\right]\right|\leq\chi\left(\pi_{n,t},\nu_{n}\right)\sqrt{\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon x}{(1+\delta)}}\right]}.$
Rearrange the inequalities and apply Cauchy-Schwarz inequality, we have
(159) $\mathbb{E}_{\pi_{n,t}}\left[e^{\frac{\varepsilon(q-n)}{2(1+\delta)}}\right]\leq\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon(q-n)}{2(1+\delta)}}\right]+\chi\left(\pi_{n,t},\nu_{n}\right)\sqrt{\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon(q-n)}{(1+\delta)}}\right]}\leq\left(1+\chi\left(\pi_{n,t},\nu_{n}\right)\right)\sqrt{\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon(q-n)}{(1+\delta)}}\right]}$
Now, we shall obtain the tail bound for $q$. Note that by Markov’s inequality and Equation (159), we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon(q-n)>x\right]\leq\mathbb{E}_{\pi_{n,t}}\left[e^{\frac{\varepsilon(q-n)}{2(1+\delta)}}\right]e^{-\frac{x}{2(1+\delta)}}\leq\left(1+\chi\left(\pi_{n,t},\nu_{n}\right)\right)\sqrt{\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon(q-n)}{(1+\delta)}}\right]}e^{-\frac{x}{2(1+\delta)}}$
And so, by applying the mixing results in Theorem 3 and the MGF bounds in Lemma 3, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}\text{ when }\alpha\in[1,+\infty),$
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}\text{ when }\alpha\in(1/2,1),$
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-D_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{e^{-\frac{4n^{1-2\alpha}}{7(1+\delta)}}+\left(10+\frac{2}{\delta}\right)n^{2\alpha-1}}\times e^{-\frac{x}{2(1+\delta)}}\text{ when }\alpha\in(0,1/2).$
Here, $C_{n},D_{n}$ are positive parameters in Theorem 3 such that $\lim_{n\rightarrow\infty}C_{n}=1$ and $\lim_{n\rightarrow\infty}D_{n}=1/25$.
Now, consider the Light Traffic regime, we denote $n_{0}$ to be the integer such that $n_{0}\geq 20$ and $\frac{\lambda_{n}}{n}\leq\frac{1}{2e}\,\forall n\geq n_{0}$. From the Markov’s inequality and Lemma 4, we have
$\mathbb{P}_{\pi_{n,t}}\left[q>n+x\right]$ $\leq\mathbb{E}_{\pi_{n,t}}\left[e^{\frac{q-n}{2}}\right]e^{-\frac{x}{2}}$
$\leq\left(1+\chi\left(\pi_{n,t},\nu_{n}\right)\right)\sqrt{\mathbb{E}_{\nu_{n}}\left[e^{q-n}\right]}e^{-\frac{x}{2}}$
$\leq\left(1+e^{-L_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{e^{-\frac{n^{1-\alpha_{n}}}{2}}+9n^{2\alpha_{n}-1}}\times e^{-\frac{x}{2}}$
and we are done.
C.3. Proof of Corollary 6
In this subsection, we provide a finite-time tail bound for the number of idle servers, which is denoted as $r_{n}=\left[n-q\right]^{+}$.
Corollary 0 (Restatement of Corollary 6).
Let $\pi_{n,t},\nu_{n}$ be the queue length distribution at time $t$ and the stationary distribution of the $M/M/n$ system with arrival rate $\lambda_{n}=n-n^{1-\alpha_{n}}$ and unit service rate respectively. Denote $r_{n}$ be a random variable denoting the number of idle servers, that is $r_{n}=\left[n-q\right]^{+}$ where $q$ is the queue length random variable. For $\alpha_{n}=\alpha\in(1/2,\infty)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[r_{n}>0\right]\leq 4e\pi n^{\frac{1}{2}-\alpha}+2\sqrt{e\pi}n^{\frac{1}{4}-\frac{\alpha}{2}}e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})$
where $\lim_{n\rightarrow\infty}C_{n}=1$ for $\alpha\in(1/2,1)$ and $C_{n}=1$ for $\alpha\in[1,\infty)$.
For $\alpha_{n}=\alpha\in(0,1/2)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[r_{n}>0\right]\geq 1-\kappa n^{\alpha-\frac{1}{2}}e^{-n^{\frac{1}{2}-\alpha}}-e^{-D_{n}t}\chi\left(\pi_{n,0},\nu_{n}\right)$
where $\lim_{n\rightarrow\infty}D_{n}=1/25$.
Proof.
Similarly to the finite-time mean queue length proof, we use also multiply both the numerator and denominator with a function to obtain the desired quantity. To do this, we use an indicator function to capture the quantity $\mathbb{P}_{\pi_{n,t}}\left[r_{n}>0\right]$. Let $1_{0\leq x\leq n-1}$ be the indicator function when the queue length $x$ is less than $n$, we have $\mathbb{E}_{\pi_{n,t}}\left[1_{0\leq x\leq n-1}\right]=\sum_{x=0}^{n-1}\pi_{n,t}(x)=\mathbb{P}_{\pi_{n,t}}[r_{n}>0]$ and $\mathbb{E}_{\nu_{n}}\left[1_{0\leq x\leq n-1}\right]=\sum_{x=0}^{n-1}\nu_{n}(x)=\mathbb{P}_{\nu_{n}}[r_{n}>0]$. We have:
$\chi^{2}\left(\pi_{n,t},\nu_{n}\right)=\sum_{x=0}^{\infty}\frac{\left(\pi_{n,t}(x)-\nu_{n}(x)\right)^{2}}{\nu_{n}(x)}$ $\geq\sum_{x=0}^{n-1}\frac{\left(\pi_{n,t}(x)-\nu_{n}(x)\right)^{2}}{\nu_{n}(x)}$
$\overset{(a)}{\geq}\frac{\left(\sum_{x=0}^{n-1}|\pi_{n,t}(x)-\nu_{n}(x)|\right)^{2}}{\sum_{x=0}^{n-1}\nu_{n}(x)}$
$\overset{(b)}{\geq}\frac{\left(\sum_{x=0}^{n-1}\pi_{n,t}(x)-\sum_{x=0}^{n-1}\nu_{n}(x)\right)^{2}}{\sum_{x=0}^{n-1}\nu_{n}(x)}$
$=\frac{\left(\mathbb{P}_{\pi_{n,t}}\left[r_{n}>0\right]-\mathbb{P}_{\nu_{n}}\left[r_{n}>0\right]\right)^{2}}{\mathbb{P}_{\nu_{n}}\left[r_{n}>0\right]}$
where (a) is the Cauchy-Schwarz inequality and (b) follows from the inequality $|x|+|y|\geq|x+y|\,\forall x,y\in\mathbb{R}$. This implies the bound
(160) $\left|\mathbb{P}_{\pi_{n,t}}\left[r_{n}>0\right]-\mathbb{P}_{\nu_{n}}\left[r_{n}>0\right]\right|$ $\leq\chi\left(\pi_{n,t},\nu_{n}\right)\sqrt{\mathbb{P}_{\nu_{n}}\left[r_{n}>0\right]}$
which is equivalent to
(161) $\mathbb{P}_{\nu_{n}}\left[r_{n}>0\right]-\chi\left(\pi_{n,t},\nu_{n}\right)\sqrt{\mathbb{P}_{\nu_{n}}\left[r_{n}>0\right]}\leq\mathbb{P}_{\pi_{n,t}}\left[r_{n}>0\right]$ $\leq\mathbb{P}_{\nu_{n}}\left[r_{n}>0\right]+\chi\left(\pi_{n,t},\nu_{n}\right)\sqrt{\mathbb{P}_{\nu_{n}}\left[r_{n}>0\right]}.$
Now, we will analyze $\mathbb{P}_{\pi_{n,t}}\left[r_{n}>0\right]$ depending on the regime. When we have $\alpha>1/2$, from Theorem 3 and Theorem 8 in (Jhunjhunwala et al., 2023), we have
$\mathbb{P}_{\pi_{n,t}}\left[r_{n}>0\right]$ $\leq\mathbb{P}_{\nu_{n}}\left[r_{n}>0\right]+\sqrt{\mathbb{P}_{\nu_{n}}\left[r_{n}>0\right]}\chi\left(\pi_{n,t},\nu_{n}\right)$
$\leq 4e\pi n^{\frac{1}{2}-\alpha}+\sqrt{4e\pi n^{\frac{1}{2}-\alpha}}\chi(\pi_{n,0},\nu_{n})e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}$
$=4e\pi n^{\frac{1}{2}-\alpha}+2\sqrt{e\pi}n^{\frac{1}{4}-\frac{\alpha}{2}}e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n}).$
On the other hand, when we have $\alpha\in(0,1/2)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[r_{n}>0\right]$ $\geq\mathbb{P}_{\nu_{n}}\left[r_{n}>0\right]-\sqrt{\mathbb{P}_{\nu_{n}}\left[r_{n}>0\right]}\chi\left(\pi_{n,t},\nu_{n}\right)$
$\geq\mathbb{P}\left[r_{n}>0\right]-\chi\left(\pi_{n,0},\nu_{n}\right)e^{-D_{n}t}\sqrt{\mathbb{P}_{\nu_{n}}\left[r_{n}>0\right]}$
$\geq 1-\kappa n^{\alpha-\frac{1}{2}}e^{-n^{\frac{1}{2}-\alpha}}-e^{-D_{n}t}\chi\left(\pi_{n,0},\nu_{n}\right).$
The second inequality follows from Theorem 3 and the last inequality follows from $\mathbb{P}\left[r_{n}>0\right]\leq 1$ and Theorem 8 in (Jhunjhunwala et al., 2023). ∎
Appendix D Miscellaneous results
In this Section, we provide the detail proof of some miscellaneous results that we use to prove our main results.
Lemma 0.
Given $x\in\mathbb{R}$ such that $|x|\leq 1$, we have
(162) $1+x\leq e^{x}\leq 1+x+x^{2}e^{x}.$
Proof.
The LHS is a well known inequality. For the RHS, consider $f(x)=e^{x}-(1+x+x^{2}e^{x})$, we want to show that $f(x)\leq 0\,\forall|x|\leq 1$. Factorize $f$, we have
$f(x)=e^{x}-(1+x+x^{2}e^{x})=(1+x)\left[e^{x}(1-x)-1\right].$
Since $|x|\leq 1$, we have $1+x\geq 0$ so it is sufficient to show that $g(x)=e^{x}(1-x)-1\leq 0$. Note that $g^{\prime}(x)=-e^{x}x$ and $g^{\prime\prime}(x)=-e^{x}(x+1)$, we have $g^{\prime}(0)=0$ and $g^{\prime\prime}(x)\leq 0\,\forall x\geq-1$. This means that $g$ attains maximum at $x=0$, which implies $g(x)\leq g(0)=0$. Hence proved. ∎
Lemma 0.
Given $x\in\mathbb{R}$ such that $0\leq x\leq\frac{1}{2}$, we have
(163) $e^{x}\leq 1+x+x^{2}.$
Moreover, for $0\leq x\leq\frac{1}{2}$, we also have
(164) $\frac{x}{2}$ $\leq\log\left(x+1\right).$
For $0\leq x\leq 1$, we have
(165) $x-\frac{x^{2}}{2}\leq\log\left(x+1\right)$
Proof.
Recall that for $20\leq x\leq 1/$ 2, we have
$e^{x}$ $=\sum_{k=0}^{\infty}\frac{x^{k}}{k!}=1+x+x^{2}-\frac{x^{2}}{2}+\sum_{k=3}^{\infty}\frac{x^{k}}{k!}$
$\leq 1+x+x^{2}-\frac{x^{2}}{2}+\frac{x^{3}}{6}\left(\sum_{k=0}^{\infty}x^{k}\right)$
$=1+x+x^{2}-\frac{x^{2}}{2}+\frac{x^{3}}{6(1-x)}$
$\leq 1+x+x^{2}-\frac{x^{2}}{2}+\frac{x^{3}}{3}\leq 1+x+x^{2}.$
Furthermore, note that $x\geq 0$, and so
$e^{x}$ $\leq 1+x+x^{2}\leq(x+1)^{2}$
$\Leftrightarrow x$ $\leq 2\log(x+1)$
$\Leftrightarrow\frac{x}{2}$ $\leq\log(x+1).$
On the other hand, let $f(x)=\log(x+1)-\left(x-\frac{x^{2}}{2}\right)$, we have $f^{\prime}(x)=\frac{1}{x+1}-1+x\geq 0\,\forall x\in[0,1]$. Thus, we have $f(x)\geq f(0)=0$, which implies
$x-\frac{x^{2}}{2}\leq\log(x+1)\,\forall x\in[0,1].$
Hence proved. ∎
Lemma 0.
Let $\delta$ be a given positive real number, $n\geq 2$ be a positive integer and denote $\lambda=n-n^{1-\alpha}$ where $\alpha\in(0,+\infty)$. We have
(166) $\frac{1}{n^{\alpha}\left(1-\left(\frac{\lambda}{n}\right)^{\frac{\delta}{\delta+1}}\right)}\leq 1+\frac{1}{\delta}.$
Moreover, we also have
(167) $\lim_{n\rightarrow\infty}\frac{1}{n^{\alpha}\left(1-\left(\frac{\lambda}{n}\right)^{\frac{\delta}{\delta+1}}\right)}=1+\frac{1}{\delta}.$
Proof.
Observe that
$\frac{1}{n^{\alpha}\left(1-\left(\frac{\lambda}{n}\right)^{\frac{\delta}{\delta+1}}\right)}$ $=\frac{1}{n^{\alpha}\left(1-\left(1-\frac{n-\lambda}{n}\right)^{\frac{\delta}{\delta+1}}\right)}$
$=\frac{1}{n^{\alpha}\left(1-\left(1-n^{-\alpha}\right)^{\frac{\delta}{\delta+1}}\right)}$
$\leq\frac{n^{-\alpha}}{1-e^{-\frac{\delta}{\delta+1}n^{-\alpha}}}$
$\leq 1+\frac{1}{\delta}$
where the two inequalities are obtained by applying the LHS of Lemma 1, and thus the inequality is established.
For the limit, we have
$\lim_{n\rightarrow\infty}\frac{1}{n^{\alpha}\left(1-\left(\frac{\lambda}{n}\right)^{\frac{\delta}{\delta+1}}\right)}$ $=\lim_{n\rightarrow\infty}\frac{1}{n^{\alpha}\left(1-\left(1-\frac{n-\lambda}{n}\right)^{\frac{\delta}{\delta+1}}\right)}$
$=\lim_{n\rightarrow\infty}\frac{1}{n^{\alpha}\left(1-\left(1-n^{-\alpha}\right)^{\frac{\delta}{\delta+1}}\right)}$
=$\overset{(a)}{=}\lim_{n\rightarrow\infty}\frac{1}{n^{\alpha}\left(1-e^{-\frac{\delta}{\delta+1}n^{-\alpha}}\right)}$
=$\overset{(b)}{=}\frac{\delta+1}{\delta}\lim_{n\rightarrow\infty}\frac{\frac{\delta}{\delta+1}n^{-\alpha}}{\left(1-e^{-\frac{\delta}{\delta+1}n^{-\alpha}}\right)}$
$=\frac{\delta+1}{\delta}$
where $(a)$ follows from $\lim_{x\rightarrow\infty}\left(1-\frac{1}{x}\right)^{x}=\frac{1}{e}$ and $(b)$ follows from the L’Hopital rule. ∎