### v4

# Finite-Time behaviors of M/M/$n$ Queue: Mixing Time, Mean Queue Length and Tail Bounds

Abstract
Service systems like data centers and ride-hailing are popularly modeled as queueing systems in the literature. Such systems are primarily studied in the steady state due to their analytical tractability. However, almost all applications in real life do not operate in a steady state, so there is a clear discrepancy in translating theoretical queueing results to practical applications. To this end, we provide a finite-time convergence for Erlang-C systems (also known as $M/M/n$ queues), providing a stepping stone towards understanding the transient behavior of more general queueing systems. We obtain a bound on the Chi-square distance between the finite time queue length distribution and the stationary distribution for a finite number of servers. We then use these bounds to study the behavior in the many-server heavy-traffic asymptotic regimes. The Erlang-C model exhibits a phase transition at the so-called Halfin-Whitt regime. We show that our mixing rate matches the limiting behavior in the Super-Halfin-Whitt regime, and matches up to a constant factor in the Sub-Halfin-Whitt regime.
To prove such a result, we employ the Lyapunov-Poincaré approach, where we first carefully design a Lyapunov function to obtain a negative drift outside a finite set. Within the finite set, we develop different strategies depending on the properties of the finite set to get a handle on the mixing behavior via a local Poincaré inequality. A key aspect of our methodological contribution is in obtaining tight guarantees in these two regions, which when combined give us tight mixing time bounds. We believe that this approach is of independent interest for studying mixing in reversible countable-state Markov chains more generally.
1 Introduction
Queueing theory has proved to be a successful tool for studying several service systems like analyzing data centers [46, 48, 47], wireless networks [58, 59, 60], and ride-hailing and EV systems [81, 83, 82]. A popular approach is to study such systems in the steady-state owing to the tractability of analysis. However, these systems rarely operate in a steady state and so, the theoretical analysis does not directly translate into real-life performance guarantees. Furthermore, a typical goal in the performance analysis of service systems is to understand queue length/waiting time behavior. However, it is not possible to obtain the exact queue length distribution except in special cases such as the $M/M/1$ queue [67, 39, 54]. So typically, one uses the steady-state behavior as a proxy for the finite-time behavior. However, this approximation is useful only when the system mixes quickly, i.e. approaches the steady-state behavior quickly. These discrepancies motivate us to analyze the rate of convergence to the steady state. Such a result allows one to better understand the applicability of the steady-state results. As the transient analysis of queueing systems is known to be challenging, we restrict our attention to one of the simplest systems, the Erlang-C model (also known as the $M/M/n$ queue).
Moreover, sometimes it is not even possible to obtain the closed form steady-state distribution of the system in many queuing systems, and therefore one resorts to asymptotic analysis such as heavy traffic analysis [33]. Recent works have characterized the rate of convergence in these regimes, thus obtaining a handle on the pre-asymptotic queue length behavior [35, 9]. Since these approximation errors are larger when one is further away from the asymptotic regime, and therefore one has a better handle on the stationary distribution when one is closer to the asymptotic regime.
However, prior works suggest that the queuing systems mix slower as one gets closer to the asymptotic regimes, for example, in the Heavy Traffic regime [61, 67, 29]. And so, we have a dilemma where the system has a larger steady-state approximation error if it is further away from the asymptotic regime, but mixes slower when it is closer to the asymptotic regime. And so, it is important to characterize the goodness of the steady-state analysis and asymptotic analysis to the finite-time queue length behavior. Therefore, the goal of this paper is to develop a theoretical framework to analyze the mixing time of reversible Markov chains, which entails Erlang-C systems. Having such a framework would allow us to obtain queue length bounds in finite-time and pre-asymptotic regimes (i.e. with a finite number of servers) for the $M/M/n$ queue and finite-time analysis for other various applications such as learning policies for queues [79] or Markov chain perturbations [52, 69].
1.1 Contributions
In this work, we consider the Erlang-C system with $n$ servers whose service time is exponentially distributed with mean $1$ and the arrival rate $\lambda_{n}=n-n^{1-\alpha}$ for some many-server heavy traffic parameter $\alpha>0$. It is known that the Erlang-C system exhibits a phase transition at $\alpha=1/2$, and the regimes $\alpha\in(0,1/2)$ and $\alpha\in(1/2,\infty)$ are called the Sub-Halfin-Whitt and the Super-Halfin-Whitt regime respectively, each with a distinct behavior. Therefore, our results and our analysis are different for each regime. That being said, our contributions can be summarized as follows.
Tight mixing of $M/M/n$: We obtain tight convergence to stationarity results at time $t$ in terms of Chi-square distance for finite time and a finite number of servers. In particular, let $\pi_{n,t}$ and $\nu_{n}$ be the queue length distribution of the $M/M/n$ system at time $t$ and at the steady-state (i.e. $t\rightarrow\infty$) respectively, we obtain the convergence results in the form of
$\chi\left(\pi_{n,t},\nu_{n}\right)\leq e^{-\square t}\chi\left(\pi_{n,0},\nu_{n}\right)$
where the $\square$ term is called the mixing rate of the system and $\chi$ is the square root of the Chi-square distance, which would be defined in Subsection 2.2. Our results are true for all time $t$ and for all $n$ (number of servers) and such finite-time results can be obtained from any initial queue length distribution. Furthermore, our results are tight in the sense that as $n$ goes to infinity, our mixing rate matches the mixing rate of $M/M/1$ and $M/M/\infty$, depending on the regime. In particular, for $\alpha\in(1/2,\infty)$, the system behavior resembles an $M/M/1$ system with service rate $n$. On the other hand, for $\alpha\in(0,1/2)$, the system will have an idle server with a high probability, and this resembles the behavior of an $M/M/\infty$ system.
We describe our finite-time convergence results for each regime in detail as follows.
  • Super-Halfin-Whitt regime ($\alpha\in(1/2,\infty)$): In this regime, we show the convergence rate to stationarity to be $e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}$ where $C_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}C_{n}=1$ for $\alpha\in(1/2,1)$ and $C_{n}=1$ for $\alpha\in[1,\infty)$. Our results are generalizations of the finite-time behavior of $M/M/1$ since they match the limiting behavior of $M/M/n$ in this regime, which is expected to resemble the behavior of $M/M/1$ with a service rate $n$ (whose mixing rate is known to be $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ [54, 73, 67, 29]).
  • Sub-Halfin-Whitt regime ($\alpha\in(0,1/2)$): We show that the mixing rate of $M/M/n$ queue in the Sub-Halfin-Whitt regime is $D_{n}$, where $D_{n}$ is a positive parameter satisfying $\lim_{n\rightarrow\infty}D_{n}=1/25$, which matches the asymptotic behavior up to a constant factor. Moreover, if we have $\lambda_{n}\in\mathbb{Z}^{+}$ then we show that one can match the asymptotic behavior at the limit, i.e. the mixing rate approaches $1$ as in the $M/M/\infty$ system [12]. We conjecture that the condition $\lambda_{n}\in\mathbb{Z}$ is not required, and getting a mixing rate approaching $1$ without this condition will be a future work.
  • Mean Field regime $\left(\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c\in[0,1)\right)$: Finally, in this regime, we show the mixing rate of the system approaches 1c1-c as $n$ goes to infinity. When $c=0$, we have the mixing rate approaches $1$, which matches the mixing rate of $M/M/\infty$ at the asymptotic. Notably, this result does not rely on the arrival rate being an integer to match the mixing rate of $M/M/\infty$ at the asymptotic.
These results are formally stated in Theorem 1. To the best of our knowledge, our work is the first to establish Chi-square convergence results for queueing systems in contrast to most previous works that used TV distance and Wasserstein distance which would not be sufficient to establish finite-time statistics. On the other hand, the obtained Chi-square distance will allow us to obtain finite-time behavior characterizations of $M/M/n$ systems (i.e. the Erlang-C system), such as:
  • Mean queue length: we obtain an upper bound of the distance between the finite-time mean queue length and the steady-state mean queue length. The rate of which the finite-time mean queue length converges to the steady-state mean queue length is dictated by the mixing rate of the system. We formally state this result in Corollary 4.
  • Finite-time concentration bound of the number of customers: we obtain a finite-time concentration bound for the queue length. Our tail bound results match the steady-state and asymptotic behavior in [35, 9]. The queue length concentration bound results are formally stated in Corollary 5.
  • Probability of having an idle server: We obtain a bound on the probability of having an idle server at time $t$ (i.e. the queue length is less than $n$). Moreover, we also recover that as $n$ goes to infinity, this probability goes to 0 in the Super-Halfin-Whitt regime and goes to a value close to $1$ in the Sub-Halfin-Whitt regime. The full result is stated in Corollary 6.
In summary, our finite-time behavior characterizations match the steady-state and asymptotic results previously established in [35, 9] and provide a holistic insight for Erlang-C systems.
1.2 Our approach: The Lyapunov-Poincaré method
In contrast to finite Markov chains, continuous state space Markov chains and Markov diffusions where there is a vast body of mixing literature for these settings [40, 53, 26, 65, 64, 3], there are only a few toolkits to analyze countable state space Markov chains due to a lack chain rule and many methods in the finite state space are no longer applicable due to the infiniteness of the state space. To this end, we aim to develop toolkits to analyze such systems using the Lyapunov drift, which exists when the system is positive recurrent.
Previously, it is well-known that one can obtain exponential ergodicity for a stochastic system if the system admits some Lyapunov drift outside of some finite set [51, 23, 68, 44, 5, 22, 50, 71, 2]. While one can easily get a handle on the mixing outside of the said set using the negative drift, it is much harder to handle the behavior inside the finite set. The special case when the finite set is a singleton (which is the case in an $M/M/1$ queue) is well understood [44]. When the finite set is not a singleton, one typically uses a ”stitching theorem” to combine knowledge of the drift outside the finite set and local mixing behavior inside the finite set to obtain a mixing time bound for the entire system. Prior work has explored the use of minorization [23, 68, 5, 71], contraction [26, 31] within the finite set or obtain the local Poincaré constant by using the radius of the finite set [21, 66, 20], and then a corresponding stitching theorem to obtain mixing rates. However, these approaches either do not lead to a tight mixing rate inside the finite set (especially for high-dimensional settings [63]) or are hard to obtain in practice. To this end, we propose the Lyapunov-Poincaré method, where we combine the Lyapunov drift outside of a finite set with a local Poincaré inequality inside the finite set to establish mixing results. Our work extends the previously established Lyapunov drift framework [51, 23, 68, 44, 5, 22, 50, 66, 71, 2] to allow a more fine-grained mixing analysis, which eventually enables us to obtain tight mixing bounds. In particular, we develop new stitching theorems so that depending on the properties of the finite set, we can better combine the drift properties outside the finite set with the local mixing behavior inside the finite set (which is characterized in terms of local Poincaré inequalities) so as to obtain more fine-grained mixing analysis.
Handling of the finite set: Our work takes novel approaches to obtain tight mixing bounds inside the finite set. To obtain tight mixing bounds inside the finite set, we propose two novel finite set techniques: 1) the canonical path method and 2) the truncation method. These two methods correspond to two different regimes, which allow us to obtain mixing time bounds that match the asymptotic behaviors. These methods can be summarized as follows:
  • Local canonical path method: The canonical path method, previously developed in [34], is a powerful technique used to analyze the mixing time of finite discrete Markov chains by bounding how ”congested” the state space is with respect to moving from one state to another. In particular, the canonical path method estimates the mixing time of a finite discrete Markov chain with the congestion ratio, which scales with the stationary distribution and the sum of the path length of paths passing through a particular edge. As such, the canonical path method can be viewed as a conductance-typed approach to obtain mixing. Notably, the canonical path method obtains a tight mixing bound when the Markov chain is a symmetric random walk on a path graph [6, 40], which is also the behavior of the Erlang-C system in the finite set when an appropriate Lyapunov function is chosen. By applying the local canonical path method, we obtain a tight weighted local Poincare inequality of the Erlang-C system inside the finite set, which allows us to use our stitching theorems to obtain the mixing rate bound for the entire system. We formally state our local canonical path lemma in Lemma 6.
  • Truncation: On the other hand, the aforementioned local canonical path method is only viable when the distribution inside the finite set is roughly uniform. For other finite sets whose distributions are not necessarily uniform, we propose the truncation method to obtain local mixing results inside such sets. In essence, if we can find another system with known Poincaré constant, which behaves identically inside the finite set, then we can use it to get a handle on the local Poincaré constant of the original system. With this method, we obtain tight mixing time bounds up to a constant factor for $M/M/n$ queues in the Mean Field regime, i.e. when $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c\in[0,1)$. For $c=0$ (the so-called Light Traffic regime), we have the mixing rate approaches to $1$, which matches the limiting behavior of the system.
In summary, our results and approach are summarized in Table 1 here.
Regime $M/M/\infty$ Mean Field Sub-HW Halfin-Whitt Super-HW $M/M/1$
$\alpha\approx 0$ $\alpha\in(0,1/2)$ $\alpha=1/2$ $\alpha\in(1/2,1)$ $\alpha\in[1,\infty]$
$\lambda_{n}/n\rightarrow c$
V(q) Linear Constant + Exp Linear + Exp Two-sided Exp Exp
Finite set Singleton Truncation method Canonical path method Singleton
Mixing rate 1 $L_{n}$ $D_{n}$ $\Omega(1)$ $C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$
$e^{-\square t}$ $L_{n}\rightarrow 1-c$ $D_{n}\rightarrow\frac{1}{25}$ $C_{n}\rightarrow 1$
Table 1: Overview of mixing time analysis using the Lyapunov-Poincaré method for $M/M/n$ for $\lambda_{n}=n-n^{1-\alpha},\mu=1$ where $\alpha$ is a parameter and $V(q)$ is the Lyapunov function. The number of waiting jobs is denoted as $\left[q-n\right]^{+}=\max\{q-n,0\}$, and the asymptotic behavior is the queue length distribution behavior as $n\rightarrow\infty$. When $\alpha\approx\infty$, this is called the Classical Heavy Traffic regime. For the Mean Field regime, we write $\alpha\approx 0$ as an intuitive explanation, whereas the actual condition is $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c\in[0,1)$.
1.3 Related works
Unlike many mixing time works which focus on the discrete-time finite state space Markov chains [34, 42, 6, 40, 85], analyzing the mixing behavior of queuing systems is a non-trivial endeavor as it involves handling an infinite, countable state space. Indeed, while there have been prior works on the mixing time of $M/M/n$ queues [67, 29, 86, 80] and other queuing systems [41, 49, 30, 7, 8, 43, 70], each of them can either only obtain tight finite-time bounds that match the asymptotic behavior only for a specific regime or setting, or can only show exponential ergodicity without an explicit rate [30, 7, 8], geometric convergence to a ball [43] or subgeometric convergence [70]. In particular, we shall provide a recap of mixing results for each of these regimes and settings.
$M/M/1$ queue: The paper [54] is the first to characterize the transient behavior of an $M/M/1$ queue. Following this work, [73] uses the spectral representation of birth-and-death processes to obtain the rate of convergence. On the other hand, [67] uses a coupling argument to establish the distance to stationarity for the $M/M/1$ queue where the author upper bounds the distance to stationarity with the probability that the two processes will coalesce within time $T$, which can be further upper bounded by the probability of hitting 0 within time $T$. Unfortunately, both of these techniques are difficult to generalize to the $M/M/n$ queue given its more complicated dynamics.
$M/M/\infty$ queue: The explicit characterization of the transient distribution of the queue length is known for the $M/M/\infty$ queue. Such a result is obtained by considering a two-dimensional Poisson point process representing the arrival time and the service time. A different approach is to bound the Entropy of the system [12]. In this work, the author uses a diffusion approximation of the $M/M/\infty$ queuing system and uses the binomial-Poisson entropic inequalities to establish the rate of convergence.
Now, we will look at the mixing results for specific $M/M/n$ regimes and queueing settings.
$M/M/n$ queue: In [86], the author uses the log-norm to characterize the convergence rate of the birth-and-death process, which applies to the $M/M/n$ setting. This method obtains the tight mixing rate of $(\sqrt{n\mu}-\sqrt{\lambda})^{2}$ for $\alpha\geq 1$ and a sub-optimal mixing rate of $\mu-\frac{\lambda}{n-1}\ll(\sqrt{n\mu}-\sqrt{\lambda})^{2}$ for $\alpha\in(1/2,1)$. The log-norm approach is also applied to analyze queueing models with queue-length-dependent admission control [55], however, with a sub-optimal mixing rate [72]. This suggests that while simple, the log-norm approach is not strong enough to obtain sharp convergence rates for a wide variety of settings.
In the Halfin-Whitt regime, [29] uses the spectral representation of birth-and-death processes, previously established by [36], to obtain the rate of convergence of the $M/M/n$ queue in the Halfin-Whitt regime, i.e. $\alpha=1/2$. Most notably, the authors were able to characterize the phase transition of the mixing rate. In particular, by letting the arrival rate $\lambda_{n}=n-B\sqrt{n}$ and $\mu=1$, the mixing rate transitions from $(\sqrt{n\mu}-\sqrt{\lambda})^{2}$ to $\approx 1$ as $B$ increases from $0\rightarrow\infty$, establishing a phase transition from $M/M/1$-like behavior to $M/M/\infty$-like behavior even in the Halfin-Whitt regime. In their work, [29] establishes a bound on $|P_{ij}(t)-\pi_{j}(\infty)|$ for $|i-j|=O(\sqrt{n})$ and for a sufficiently large $n$, which only provides guarantee for sufficiently close states, but do not state finite-time moment bounds and tail bounds. These limitations motivate us to consider finite time bounds in the Chi-square metric in our work.
However, these prior results using the spectral representation can be used to obtain finite-time Chi-square bounds. While no prior work has explicitly stated such results, these results can be inferred by putting together several works in the literature. We obtain this and state a formal result in Appendix 12. We also present a discussion contrasting this work with our approach, as well as an in-depth comparison of our work with previous works on the $M/M/n$ systems. In particular, we would like to highlight that while our approach is applicable for any reversible Markov chain, the orthogonal polynomial approach is limited to birth-death chains.
Jackson networks: For Jackson networks, [41] shows exponential ergodicity, but their bound using Cheeger’s inequality is not tight in terms of the traffic parameter, and the dependence of the mixing time on the spectral gap of the communication graph is not discussed. The spectral gap of the open Jackson network is established in [49] using the Markov chain decomposition method.
Load-balancing systems: In addition to $M/M/n$ and Jackson networks, mixing results have been previously established for some load-balancing algorithms. The paper [70] obtained sub-geometric convergence for the spatial load-balancing system and [45] obtained $O(1/t)$ convergence rate for the join-the-shortest-queue system with $2$ servers. On the other hand, [43, 10] obtained exponential convergence to a ball in the sub-Halfin-Whitt regime for the join-the-shortest-queue and power-of-$d$ policies.
Birth and death processes: Besides queuing systems, birth and death processes (which entail $M/M/n$ queues) are common countable state space Markov chains. There are several works analyzing the spectral gap of birth and death processes [36, 73, 74, 72, 77, 75]. Additionally, it is known that these systems admit a system of orthogonal polynomials with respect to some spectral measure [19], which can be used to study exponential ergodicity [36, 29], $\ell_{2}$ convergence [14], or hitting time properties [76]. For non-homogeneous birth-death processes, [87] provides an analysis on the convergence of these processes and applies it to finite Markov chains. However, most of these works exploit the specific structure of birth and death processes (such as the spectral representation) which severely limits its applicability. In contrast, we aim to develop a more general framework to analyze the mixing behavior of reversible Markov chains.
1.3.1 Lyapunov drift method
It is well-known that an irreducible Markov chain is positive recurrent if and only if it has a negative drift outside of a bounded set [28], which is famously known as the Foster-Lyapunov Theorem. Since then, several works have used the drift assumption to establish the convergence rate of the Markov chain. A long line of literature [22, 50, 51, 23, 68, 44, 5, 71, 31, 26] establish geometric ergodicity for general Markov chains whenever we have the Lyapunov drift condition outside of some set and some control inside of that set (i.e. the set is a singleton or the set admits a minorization or a contraction condition). Recently, [1] proposes a drift and hitting time approach instead of using the minorization assumption, and in a similar spirit, [11] uses the exponential integrability of the hitting time to obtain Poincaré inequalities. However, these approaches do not allow us to obtain tight convergence rates for our case.
The Lyapunov drift method is known to be very effective in establishing steady-state queueing results [27, 33, 35] and beyond that, convergence and steady-state results in Stochastic Approximation and Optimization [16, 17, 15, 56, 57]. As Lyapunov drift has been well-studied in the queuing literature, we naturally want to leverage this drift to obtain mixing results. Previously, [30] obtained exponential convergence results for Markovian queues but the focus is not on obtaining tight mixing rates. For JSQ, [7, 8] uses the Lyapunov drift mixing framework by [23]. All of these works, however, only simply show exponential ergodicity and do not get a characterization on the mixing rate. This suggests that to get a tight mixing rate bound, a more fine-grained application of the Lyapunov drift method is required.
In addition to the negative drift approach and the convergence in $\ell_{2}$ distance or TV distance, some other works have opted for convergence in Wasserstein space instead [65, 64, 25, 24] since many systems such as the constant step size SGD may not converge in TV distance but does converge in Wasserstein distance [65]. In this metric, [65] proposes the notion of contractive drift rather than the Foster-Lyapunov drift to obtain the finite-time convergence of various systems such as the single server queue and the constant step size SGD. Notably, this notion allows us to bypass the bounded set problem, where the drift and minorization approach is known to have limitations [63]. On the other hand, the approach of using a contraction condition to handle the bounded set (also known as drift and contraction) often leads to a tradeoff in terms of the drift rate and the bounded set size [65]. Instead of using this approach, our work builds on the Lyapunov drift framework to obtain Chi-square convergence, and we argue that our work is another approach to handling the bounded set issue that is typically faced when using drift arguments to obtain mixing results.
We summarize our results and techniques in Table 1. Our work provides mixing time analysis in the Mean Field regime (where the arrival rate $\lambda_{n}$ is chosen such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c$), the Sub-Halfin-Whitt regime (i.e. $\alpha\in(0,1/2)$) and the Super-Halfin-Whitt regime (i.e. $\alpha\in(1/2,\infty)$). Additionally, note that our mixing results also match the asymptotic and the steady-state behavior of the system, that is the system behaves like $M/M/1$ for $\alpha\in(1/2,\infty)$ and $M/M/\infty$ for $\alpha\in(0,1/2)$.
1.4 Notations
To help the readers better understand our work, we define some notations used in our work as follows. For a countable set $\mathcal{S}$, denote the space of functions $f:\mathcal{S}\rightarrow\mathbb{R}$ that are square-integrable by $\ell_{2,\pi}$. Now, define the inner product
$\langle f,g\rangle_{\pi}:=\sum_{x\in\mathcal{S}}\pi(x)f(x)g(x)$
for $f,g\in\ell_{2,\pi}$. From the definition of the inner product, we have $\|f\|^{2}_{2,\pi}:=\langle f,f\rangle_{\pi}$. The distribution $\pi$ is said to be reversible for the operator $\mathcal{L}:\ell_{2,\pi}\rightarrow\ell_{2,\pi}$ if $\mathcal{L}$ is self-adjoint on $\ell_{2,\pi}$, i.e.,
$\langle f,\mathcal{L}g\rangle_{\pi}=\langle\mathcal{L}f,g\rangle_{\pi},\quad\forall f,g\in\ell_{2,\pi}.$
In addition, for a well-behaved function $f$, we denote $\pi(f)=\mathbb{E}_{\pi}(f),\operatorname{Var}_{\pi}(f)=\mathbb{E}_{\pi}(f^{2})-\mathbb{E}_{\pi}(f)^{2}$, $\mathbf{1}$ as the vector of $1$s. For a set $K$, we denote $\pi(K)=\int_{K}d\pi$ and $1_{K}$ as the indicator function for the set $K$ where $1_{K}(x)=1$ if $x\in K$ and $1_{K}(x)=0$ otherwise. All logarithms in this paper are natural logarithms. Finally, we denote $\left[x\right]^{+}=\max\{0,x\}$ and we interpret $\frac{0}{0}$ as 0.
2 Model and Main Results
2.1 Model
Consider the $M/M/n$ queueing system, also known as the Erlang-C system, described as follows. Customers arrive at a central queue following a Poisson process with mean $\lambda$. There are $n$ homogeneous servers that serve the waiting customers in a first-come-first-serve (FCFS) manner. The service time of each customer is independently and exponentially distributed with mean $1/\mu$. Without the loss of generality, we assume $\mu=1$ unless specified otherwise. The number of customers $\{q(t):t\geq 0\}$ in the queue evolves as a continuous time Markov chain (CTMC) with state space $\mathbb{Z}_{\geq 0}$. This CTMC is a birth-and-death process with the birth rate $\lambda$ and the death rate $\min\{n,q\}$ in the state $q(t)=q$. We denote $\mathcal{L}$ as the infinitesimal generator of the CTMC and $P_{t}$ as the corresponding Markov semigroup where $P_{t}=e^{t\mathcal{L}}$. For a given $n$, we denote the stationary distribution of $M/M/n$ queue by $\nu_{n}$ and we omit the subscript of $n$, whenever it is clear from the context.
In addition, we define $M/M/\infty$ as the birth-and-death process with the birth rate $\lambda$ and the death rate $q$ in the state $q(t)=q$. This system behaves as if there are an infinite number of servers and whenever a customer arrives, it is immediately allocated an idle server to get served.
2.2 Chi-square distance
To measure the distance to stationarity, we use a notion of distance called the Chi-square distance, which is formally defined as follows:
Definition 1.
Given two distributions $P,Q$ whose supports are the subset of the countable set $\mathcal{S}$ and $P$ is absolutely continuous w.r.t $Q$ , the Chi-square distance of $P$ with respect to $Q$ is given by
$\chi^{2}\left(P,Q\right)=\sum_{x\in\mathcal{S}}\frac{(P(x)-Q(x))^{2}}{Q(x)}.$ (1)
As $\mathbb{E}_{Q}\left[\frac{P}{Q}\right]=1$, the Chi-squared distance $\chi^{2}(P,Q)$ can also be rewritten as
$\chi^{2}(P,Q)=\sum_{x\in\mathcal{S}}\frac{(P(x)-Q(x))^{2}}{Q(x)}=\sum_{x\in\mathcal{S}}Q(x)\left(\frac{P(x)^{2}}{Q(x)^{2}}-1\right)=\operatorname{Var}_{Q}\left(\frac{P}{Q}\right).$ (2)
The reason that we are interested in the Chi-square distance is that transient mean queue length and tail bounds can be derived from the Chi-square distance to stationarity. Indeed, we have the following Theorem on the variational form of the Chi-square distance.
Proposition 1.
Given two distributions $P,Q$ with countable supports and let $\mathcal{G}_{Q}$ be a collection of functions $g:\mathbb{R}\rightarrow\mathbb{R}$ such that $\mathbb{E}_{Q}\left[g(X)^{2}\right]<\infty\}$, we have
$\chi^{2}(P,Q)=\sup_{g\in\mathcal{G}_{Q}}\frac{\left(\mathbb{E}_{P}[g(X)]-\mathbb{E}_{Q}[g(X)]\right)^{2}}{\operatorname{Var}_{Q}[g(X)]}$ (3)
The proof of Proposition 1 is well-known and is presented in Appendix 8.3. Additionally, Proposition 7.15 in [62] states that the Chi-square distance can be used to give an upper bound on the TV distance as well. From Proposition 1, we can substitute the function $g$ of our interests to obtain specific bounds. In particular, we obtain the following bound on the distance of the moments of the distributions by substituting $g(x)=x^{k}$.
Corollary 1.
Given two distributions $P,Q$ with countable supports. In addition, let $k\in\mathbb{Z}^{+}$ and assume that $\mathbb{E}_{Q}[X^{2k}]<\infty$, then
$\left|\mathbb{E}_{P}\left[X^{k}\right]-\mathbb{E}_{Q}\left[X^{k}\right]\right|\leq\chi(P,Q)\sqrt{\mathbb{E}_{Q}\left[X^{2k}\right]-\mathbb{E}_{Q}\left[X^{k}\right]^{2}}.$ (4)
We can substitute $g(X)=e^{\theta X}$ for an appropriately chosen $\theta\in\mathbb{R}$ to establish a bound on the difference of the Moment Generating Functions (MGF) as stated in the following corollary.
Corollary 2.
Given two distributions $P,Q$ with countable supports and $\theta\in\mathbb{R}^{+}$ such that $\mathbb{E}_{Q}[e^{2\theta X}]<\infty$, then
$\left|\mathbb{E}_{P}\left[e^{\theta X}\right]-\mathbb{E}_{Q}\left[e^{\theta X}\right]\right|\leq\chi\left(P,Q\right)\sqrt{\mathbb{E}_{Q}\left[e^{2\theta X}\right]-\mathbb{E}_{Q}\left[e^{\theta X}\right]^{2}}.$ (5)
The MGF bounds are particularly useful for obtaining exponential tail bounds. With the proper scaling, many classical queueing systems such as $M/M/n$ and JSQ possess an MGF [33, 35], and thus one can derive exponential tail bounds for such systems.
These results will be the bread and butter to obtain the finite-time statistics of our queueing systems of interest. Indeed, let $\pi_{n,t}$ be the queue length distribution at time $t$ and $\nu_{n}$ be the stationary distribution of the $M/M/n$ system, if we can obtain a bound on $\chi(\pi_{n,t},\nu_{n})$ then we can apply these results to obtain finite-time statistics such as mean queue length bounds and tail bounds. To this end, we develop machinery to establish the distance to stationarity bounds for $\pi_{n,t}$, in turn, resulting in finite-time statistics.
2.3 Main Results
Consider an $M/M/n$ queue with arrival rate $\lambda_{n}=n-n^{1-\alpha}$ for some $\alpha\in\mathbb{R}^{+}$ and service rate $\mu=1$ (unless stated otherwise). Such a parametrization of $\lambda_{n}$ is popularly known as the many-server-heavy-traffic regime and queueing systems are generally analyzed for $t,n\rightarrow\infty$. On the contrary, here we focus on a finite $t\geq 0$ and a finite $n\geq 1$. We obtain a bound on $\chi(\pi_{n,t},\nu_{n})$ of the form $e^{-\square t}$, thus, establishing the convergence of the finite-time queue length distribution $\pi_{n,t}$ to the stationary distribution $\nu_{n}$ as $t\rightarrow\infty$. Moreover, we provide a tight bound on the mixing rate, denoted by $\square$. This result is established in the following theorem.
Theorem 1.
Let $\pi_{n,t}$ be the queue length distribution at time $t\geq 0$ of the continuous-time $M/M/n$ system defined above and let the stationary distribution be $\nu_{n}$.
For $\alpha\geq 1,n\geq 2$, we have that:
$\chi(\pi_{n,t},\nu_{n})\leq e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0.$ (6)
For $\alpha\in(1/2,1),n\geq 2$, we have that:
$\chi(\pi_{n,t},\nu_{n})\leq e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0,$ (7)
for some $C_{n}>0$ such that $\lim_{n\rightarrow\infty}C_{n}=1$.
For $\alpha\in(0,1/2),n\geq 2$, we have that:
$\chi(\pi_{n,t},\nu_{n})\leq e^{-D_{n}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0,$ (8)
for some $D_{n}>0$ such that $\lim_{n\rightarrow\infty}D_{n}=1/25$.
Finally, for $\alpha=1/2$ and $n\geq 110$, we have
$\chi(\pi_{n,t},\nu_{n})\leq e^{-H_{n}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0$ (9)
for some $H_{n}>\frac{1}{10861}$ and $\lim_{n\rightarrow\infty}H_{n}=\frac{1}{1781}$.
The proof of this theorem and the exact expressions of $C_{n},D_{n},H_{n}$ are described in Appendix 9, and its proof sketch is summarized in Section 4. Note that we assume $n\geq 2$ so that $\lambda_{n}>0$ (for the convergence rate of $M/M/1$, please refer to Proposition 2), and assume $n\geq 110$ for $\alpha=1/2$ only to obtain an explicit mixing rate for this regime and to get reasonable constants in the proof. Obtaining an explicit bound for all $n\geq 1$ is possible at the cost of worse universal constants in our mixing rate bounds. Now, denote $\tau_{mix}(\varepsilon)=\min_{t}\chi(\pi_{n,t},\nu_{n})\leq\varepsilon$, we obtain the following corollary on the mixing time of the system from the bound on the Chi-square distance.
Corollary 3.
(Mixing time bounds for $M/M/n$ systems) Let $\pi_{n,t}$ be the queue length distribution at time $t$ of the continuous-time $M/M/n$ system defined above and let the stationary distribution be $\nu_{n}$. For $\alpha\geq 1$, we have that:
$\tau_{mix}(\varepsilon)\leq 4n^{2\alpha-1}\log\left(\frac{\chi(\pi_{n,0},\nu_{n})}{\varepsilon}\right).$ (10)
For $\alpha\in(1/2,1)$, we have that:
$\tau_{mix}(\varepsilon)\leq 4C_{n}^{-1}n^{2\alpha-1}\log\left(\frac{\chi(\pi_{n,0},\nu_{n})}{\varepsilon}\right)$ (11)
for some $C_{n}>0$ such that $\lim_{n\rightarrow\infty}C_{n}=1$.
For $\alpha\in(0,1/2)$, we have that:
$\tau_{mix}(\varepsilon)\leq D_{n}^{-1}\log\left(\frac{\chi(\pi_{n,0},\nu_{n})}{\varepsilon}\right)$ (12)
for some $D_{n}>0$ such that $\lim_{n\rightarrow\infty}D_{n}=1/25$.
Finally, for $\alpha=1/2$ and $n\geq 110$, we have
$\tau_{mix}(\varepsilon)\leq H_{n}^{-1}\log\left(\frac{\chi(\pi_{n,0},\nu_{n})}{\varepsilon}\right)$ (13)
for some $H_{n}>\frac{1}{10861}$ and $\lim_{n\rightarrow\infty}H_{n}=\frac{1}{1781}$.
Corollary 3 tells us that the $\varepsilon$-mixing time $\tau_{mix}(\varepsilon)$ has logarithmic dependence on the initial distance and $\varepsilon^{-1}$. Additionally, when the traffic is light (i.e. $\alpha\in(0,1/2)$), it takes $O(1)$ time to mix as $D_{n},H_{n}=\Omega(1)$. However, when the system has heavy traffic (i.e. $\alpha>1/2$), $n^{2\alpha-1}$ would grow to infinity as $n$ grows, and so, it would take a long time to mix in this regime. We defer the proof of Corollary 3 to Appendix 10. Now, it is well-known that there is a phase transition at $\alpha=1/2$ [32], and the behavior in the regimes $\alpha\in(0,1/2)$ and $\alpha\in(1/2,\infty)$ is similar to $M/M/\infty$ and $M/M/1$ respectively. Our Theorem 1 captures such a phase transition, as evident by the mixing-time bounds for the $M/M/1$ and $M/M/\infty$ queues stated below:
Proposition 2.
Let $\pi_{t,1}$ be the queue length distribution at time $t\in\mathbb{R}^{+}$ of the continuous-time $M/M/1$ system with the arrival rate $\lambda$ and service rate $\mu$ such that $0<\lambda<\mu$ and let the stationary distribution be $\nu_{1}$, we have:
$\chi(\pi_{t,1},\nu_{1})\leq e^{-(\sqrt{\mu}-\sqrt{\lambda})^{2}t}\chi(\pi_{0,1},\nu_{1})\,\forall t\geq 0.$
Proposition 3.
Let $\pi_{t,\infty}$ be the queue length distribution at time $t$ of the continuous-time $M/M/\infty$ system with the arrival rate $\lambda$ and the service rate $\mu$ and let the stationary distribution be $\nu_{\infty}$, we have if $\lambda/\mu\in\mathbb{Z}^{+}$ then:
$\chi(\pi_{t,\infty},\nu_{\infty})\leq e^{-\mu t}\chi(\pi_{0,\infty},\nu_{\infty})\,\forall t\geq 0.$
While the above results for $M/M/1$ and $M/M/\infty$ are previously established in [12, 54, 67], it is worth noting that our Lyapunov-Poincaré methodology provides a unified approach to obtain mixing results for $M/M/1$, $M/M/n$, and $M/M/\infty$ queues. The proof of Proposition 2, presented in Appendix 9.1.4, is instructive to understand the case of $\alpha\in(1/2,\infty]$ of Theorem 1. Similarly, the proof of Proposition 3, while it is only applicable for $\lambda\in\mathbb{Z}^{+}$, in Appendix 9.5 is helpful to understand the case of $\alpha\in(0,1/2)$. Observe that the mixing rate of the $M/M/n$ queue for $\alpha\in(1/2,\infty]$ and the $M/M/1$ queue is $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}\approx n^{1-2\alpha}\rightarrow 0$ as $n\rightarrow\infty$. On the other hand, the mixing rate of the $M/M/n$ queue for $\alpha\in(0,1/2)$ and the $M/M/\infty$ queue is $\Theta(1)$ as $n\rightarrow\infty$. Thus, our theorem characterizes the phase transition of the mixing rate from 0 to $\Theta(1)$ at $\alpha=1/2$. Finally, consider the case of $\alpha=1/2$, similar to the case of $\alpha\in(0,1/2)$, our mixing rate is $\Theta(1)$ but is off by a constant. For this case, a precise characterization of the mixing rate in the limit as $t\rightarrow\infty$, and $n\rightarrow\infty$ was provided in [29], and obtaining a finite time behavior for a finite-sized system that exactly matches these limiting results is an interesting future direction.
While we establish a $\Theta(1)$ bound on the mixing rate for $\alpha\in(0,1/2)$, it is off by a constant compared to the mixing rate of $M/M/\infty$ queue, indicating that our mixing rate characterization is loose in this regime. Nonetheless, we believe our characterization of the mixing rate is state-of-the-art. For example, the result of [86] implies a mixing rate of $1-\frac{\lambda_{n}}{n-1}=O\left(n^{-\alpha}\right)\rightarrow 0$ as $n\rightarrow\infty$. Moreover, by imposing additional technical assumptions, we are able to improve our results to obtain tight mixing rates in this regime. To this end, we present two results assuming that either $\lambda_{n}$ is very small (light traffic, i.e. $\lambda_{n}/n\rightarrow 0$) or $\lambda_{n}\in\mathbb{Z}^{+}$, where the latter is an artifact of our methodology as discussed in Section 3.
Proposition 4.
Let $\pi_{n,t},\nu_{n}$ be the queue length distribution at time $t$ and the steady state distribution of the continuous-time $M/M/n$ system with unit service rate respectively and let $\lambda_{n}$ be a sequence of arrival rate such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c$ where $c\in[0,1)$. We have
$\chi(\pi_{n,t},\nu_{n})\leq e^{-L_{n}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0$
for some positive parameter $L_{n}$ such that $\lim_{n\rightarrow\infty}L_{n}=1-c$.
Since $\lambda_{n}\in(0,n)$, there exists a unique positive number $\alpha_{n}$ such that $\lambda_{n}=n-n^{1-\alpha_{n}}$, and so the assumption $\lim_{n\rightarrow\infty}\lambda_{n}/n=c$ means that $\lim_{n\rightarrow\infty}\alpha_{n}\log n=-\log(1-c)$. When $c=0$, we have the mixing rate of the system matches the limiting behavior at the asymptotic (that is, mixing rate approaches $1$ as $n\rightarrow\infty$). On the other hand, for $c\in(0,1)$, Proposition 4 shows that the system admits a mixing rate that is bounded below by a constant that is independent of $n$. Moreover, when $c<\frac{24}{25}$, we get a better constant than the $\frac{1}{25}$ constant in the $\alpha\in(0,1/2)$ regime. One can find the proof of Proposition 4 in Appendix 9.4 and the discussion on this result in comparison with previous works in Appendix 12. However, since we know that the phase transition happens at the Halfin-Whitt regime, this suggests that we could do better by obtaining an analysis that matches the limiting behavior for all $\alpha\in(0,1/2)$, possibly with some additional conditions. And so, we consider the case when the arrival rate is an integer, i.e. there is a sequence $\{\lambda_{k}\}_{k\geq\mathbb{Z}^{+}}$ such that $\lambda_{k}\in\mathbb{Z}\,\forall k\in\mathbb{Z}^{+}$. In this case, we can show that as $\lambda_{k}\rightarrow\infty$, we have the mixing rate approaches $1$ as follows.
Proposition 5.
Let $\pi_{n,t},\nu_{n}$ be the queue length distribution at time $t$ and the steady state distribution of the continuous-time $M/M/n$ system with unit service rate respectively and let $\{\lambda_{n}\}$ be a sequence of integer arrival rates ($\lambda_{n}\in\mathbb{Z}$) such that $\lambda_{n}\geq 0,n-\lambda_{n}\geq 1,\frac{\log\left(n-\lambda_{n}\right)}{\log n}\in(1/2,1)$ and
$\lim_{n\rightarrow\infty}\frac{\log\left(n-\lambda_{n}\right)}{\log n}=1-\alpha$ (14)
where $\alpha<1/2$. Then, we have
$\chi(\pi_{n,t},\nu_{n})\leq e^{-\overline{D}_{n}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0$ (15)
such that $\overline{D}_{n}>0$ and $\lim_{n\rightarrow\infty}\overline{D}_{n}=1$.
The proof of Proposition 5 is presented in Appendix 9.3.3. With additional technical assumptions on $\lambda$, our mixing rate now matches with that of the $M/M/\infty$ queue asymptotically for $\alpha\in(0,1/2)$. Now, when both of these conditions are not satisfied, we can still show mixing but the mixing rate is only tight up to a constant, as shown in Theorem 1. Nevertheless, we believe that these two conditions are rather artificial and we conjecture that these conditions can be lifted to get a mixing rate approaching $1$, matching the mixing rate of the $M/M/\infty$ at the asymptotics.
2.3.1 Finite-time statistics
From the established mixing bounds in Theorem 1, we obtain finite-time statistics like the mean and tail. We start with the mean queue length below.
Corollary 4.
Let $\pi_{n,t}$ be the queue length distribution at time $t$ of the continuous-time $M/M/n$ system with the arrival rate $\lambda_{n}=n-n^{1-\alpha}$ and a service rate $1$ whose stationary distribution be $\nu_{n}$. For $\alpha\geq 1$, we have that
$\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\sqrt{2}\left(n+n^{\alpha}\right)\chi(\pi_{n,0},\nu_{n}).$ (16)
For $\alpha\in(1/2,1)$, we have
$\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\sqrt{2}\left(n+n^{\alpha}\right)\chi(\pi_{n,0},\nu_{n})$ (17)
for some $C_{n}>0$ such that $\lim_{n\rightarrow\infty}C_{n}=1$.
For $\alpha\in(0,1/2)$, we have
$\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-D_{n}t}\sqrt{2}\left(n+n^{\alpha}\right)\chi(\pi_{n,0},\nu_{n})$ (18)
for some $D_{n}>0$ such that $\lim_{n\rightarrow\infty}D_{n}=1/25$.
And finally, if $\lambda_{n}$ is a sequence of arrival rates such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c$ for $c\in[0,1)$, then we have
$\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-L_{n}t}n\chi(\pi_{n,0},\nu_{n})$ (19)
for some $L_{n}>0$ such that $\lim_{n\rightarrow\infty}L_{n}=1-c$.
One can prove Corollary 4 by combining the Theorem 1 and Corollary 1 for $k=1$. We defer the details of the proof to Appendix 11.1. Similar to the above corollary, we can obtain bounds on any moment $k\in\mathbb{Z}^{+}$ by Corollary 1 combined with a $2k$-moment bound on the steady-state distribution $\nu_{n}$. Moreover, in addition to the mean queue length bound, we can establish some tail-bound results to gain some understanding on the concetration of the finite-time behavior of the queue length distribution as follows.
Corollary 5.
Let $\pi_{n,t},\nu_{n}$ be the queue length distribution at time $t$ and the stationary distribution of the $M/M/n$ system with arrival rate $\lambda_{n}=n-n^{1-\alpha_{n}}$ and unit service rate respectively, and let $\varepsilon=1-\frac{\lambda_{n}}{n}$ and $q$ be the random variable denoting the queue length. For $n\geq n_{0},\delta\in(0,+\infty)$ and $n_{0}$ is a constant dependent on the regime, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon_{n}\left(q-n\right)>x\right]\leq\left(1+\chi(\pi_{n,t},\nu_{n})\right)\sqrt{ex}e^{-\frac{x}{2}}$
In particular, for $\alpha_{n}=\alpha\in[1,\infty)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon_{n}\left(q-n\right)>x\right]$ $\leq\left(1+e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{ex}e^{-\frac{x}{2}}.$
For $\alpha_{n}=\alpha\in(1/2,1)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon_{n}\left(q-n\right)>x\right]$ $\leq\left(1+e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{ex}e^{-\frac{x}{2}}$
where $C_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}C_{n}=1$.
For $\alpha_{n}=\alpha\in(0,1/2)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-D_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{0.5e^{-\frac{2n^{1-2\alpha}}{7}}+6n^{2\alpha-1}}\times\sqrt{ex}e^{-\frac{x}{2}}$
where $D_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}D_{n}=1/25$.
Finally, if $\lambda_{n}$ is a sequence of arrival rates such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c$ for $c\in[0,1)$ then we have
$\mathbb{P}_{\pi_{n,t}}\left[q-n>x\right]$ $\leq\left(1+e^{-L_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{e^{-\frac{(1-c)n^{1-\alpha_{n}}}{2}}+2n^{\alpha_{n}-1}+\frac{8n^{2\alpha_{n}-1}}{e(1-c)}}\times e^{-\frac{x}{2}}$
where $\lim_{n\rightarrow\infty}\alpha_{n}=0$ and $L_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}L_{n}=1-c$.
The proof of Corollary 5 is deferred to Appendix 11.2. Note that if $q(t)\geq n$, then q(t)nq(t)-n is the number of waiting customers at time $t$. The results in Corollary 5 tell us what is the tail distribution of the number of waiting customers at time $t$. In the Sub-Halfin-Whitt regime, we have the pre-exponent term approaches 0 as $n\rightarrow\infty$, which suggests a faster decay rate than exponential. Moreover, we can also estimate the probability of having an idle server at time $t$ as follows.
Corollary 6.
Let $\pi_{n,t},\nu_{n}$ be the queue length distribution at time $t$ and the stationary distribution of the $M/M/n$ system with arrival rate $\lambda_{n}=n-n^{1-\alpha_{n}}$ and unit service rate respectively. Denote $r_{n}$ be a random variable denoting the number of idle servers, that is $r_{n}=\left[n-q\right]^{+}$ where $q$ is the queue length random variable. For $\alpha_{n}=\alpha\in(1/2,\infty)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[r_{n}>0\right]\leq 4e\pi n^{\frac{1}{2}-\alpha}+2\sqrt{e\pi}n^{\frac{1}{4}-\frac{\alpha}{2}}e^{-C_{n}(\sqrt{n}-\sqrt{\lambda})^{2}t}\chi(\pi_{n,0},\nu_{n})$
where $\lim_{n\rightarrow\infty}C_{n}=1$ for $\alpha\in(1/2,1)$ and $C_{n}=1$ for $\alpha\in[1,\infty)$.
For $\alpha_{n}=\alpha\in(0,1/2)$, we have
$\mathbb{P}_{\pi_{n,t}}\left[r_{n}>0\right]\geq 1-\kappa n^{\alpha-\frac{1}{2}}e^{-n^{\frac{1}{2}-\alpha}}-e^{-D_{n}t}\chi\left(\pi_{n,0},\nu_{n}\right)$
where $\lim_{n\rightarrow\infty}D_{n}=1/25$.
The proof of Corollary 6 can be found in Appendix 11.3. These results are the finite-time version of the steady-state results established in [35] and thus can be seen as the generalization of such results.
3 The Lyapunov-Poincaré method
To establish the mixing results stated in Subsection 2.3, we dedicate the following Section to introduce the so-called Lyapunov-Poincaré machinery. All the definitions and results in this section are valid for any CTMC with a countable state space $\mathcal{S}$, generator $\mathcal{L}$, and Markov semigroup $P_{t}$.
Assumption 1.
The given CTMC is irreducible, aperiodic, positive recurrent, and reversible.
Assumption 1 implies that there exists a unique stationary distribution $\nu$ such that $\nu(x)\mathcal{L}(x,y)=\nu(y)\mathcal{L}(y,x)$ for all $x,y\in\mathcal{S}$.
Definition 2.
A CTMC admits a Poincaré constant $C_{P}$ if for all test function $f\in\ell_{2,\nu}$, we have
$\operatorname{Var}_{\nu}(f)\leq C_{P}\mathcal{E}(f,f).$ (20)
Here, the Dirichlet form $\mathcal{E}(f,f)$ is defined as:
$\mathcal{E}(f,f)=\langle f,-\mathcal{L}f\rangle_{\nu}=-\sum_{x,y\in\mathcal{S}}\nu(x)\mathcal{L}(x,y)f(x)f(y).$ (21)
It is known that Poincaré inequalities immediately imply exponential ergodicity of the CTMC [3]. We formally state it in the following proposition for completeness.
Proposition 6.
Under Assumption 1, if the system admits the Poincaré inequality (20) with constant $C_{P}$, then
$\chi^{2}\left(\pi_{t},\nu\right)\leq e^{-\frac{2t}{C_{P}}}\chi^{2}(\pi_{0},\nu)$ (22)
where $\pi_{t}$ is distribution at time $t$ and $\nu$ is the stationary distribution.
The proof of Proposition 6 is well-known and can be found in [3] and in Appendix 8.1. Proposition 6 shows that Poincaré inequality with constant $C_{P}$ implies exponential convergence in the chi-square distance with mixing rate $1/C_{P}$. Thus, for the rest of the section, our focus is on establishing the Poincaré constant.
Building on the recent developments on Markov chain mixing [71, 2], we now present the Lyapunov-Poincaré methodology below, which has two main ingredients. The first one is a Foster-Lyapunov like assumption:
Assumption 2.
There exists a Lyapunov function $V:\mathcal{S}\rightarrow[1,\infty)$ along with $b:\mathcal{S}\rightarrow[0,\infty)$ and $\gamma>0$ such that
$\mathcal{L}V(q)\leq-\gamma V(q)+b(q)\,\forall q\in\mathcal{S}.$ (23)
For the special case of $b=B1_{K}$ for some finite $K\subseteq S$ and $B>0$, the above is same as the classical Foster-Lyapunov condition. More generally, let $K=\{q:b(q)>0\}$, then in words, the above assumption ensures a negative drift proportional to $V(q)$ outside $K$. On the other hand, inside $K$, our drift might be positive but it is controlled by the $b(q)$ for $q\in K$. If $K$ is a singleton then we have a slightly modified assumption as follows.
Assumption 3.
There exists a single point $x^{*}\in\mathcal{S}$, a Lyapunov function $V:\mathcal{S}\rightarrow[0,\infty)$ satisfying $V(q)>0\,\forall q\neq x^{*}$ and constants $\gamma,B>0$ such that
$\mathcal{L}V(q)\leq-\gamma V(q)+B1_{\{x^{*}\}}\,\forall q\in\mathcal{S}.$ (24)
Unlike Assumption 2, Assumption 3 allows us to have $V(x^{*})<1$. In many settings to be discussed below, this modification is necessary to obtain the tight mixing rate in the case we have a negative drift outside of a singleton. From here, we have the following proposition.
Proposition 7.
Assume that Assumptions 1 and 3 holds where $\{b(q)>0\}\subseteq\mathcal{S}$ is a singleton, then the CTMC admits the Poincaré constant $C_{P}=\frac{1}{\gamma}$.
While a similar convergence result for TV distance was previously established [44], this result allows us to establish Chi-square convergence, and we present the proof in Appendix 8.2.3. Combining the above result with Proposition 6, one can obtain the Chi-square convergence as in (22). More generally, the finite set $K$ may not be a singleton for most of the queueing systems (including the $M/M/n$ queue), and thus, the Foster-Lyapunov assumption alone fails to provide a Poincaré inequality since we do not know how the system would mix inside $K$. So we need to establish a local mixing result inside $K$, which can be done using a local Poincaré inequality. Such a result is formalized in the following assumption.
Assumption 4.
(Weighted Poincaré) Given the stationary distribution $\nu$, a function $b:\mathcal{S}\rightarrow[0,\infty)$ such that $\sum_{q\in\mathcal{S}}b(q)\nu(q)$ is finite and let $\tau(q)=\frac{b(q)\nu(q)}{\sum_{q\in\mathcal{S}}b(q)\nu(q)}\forall q\in\mathcal{S}$ be a $(b,\nu)$-weighted distribution. Denote $K=\{q\in\mathcal{S}:b(q)>0\}$, a CTMC is said to admit a $b$-weighted Poincaré inequality if there exists a non-negative constant $C_{b}$ such that
$\operatorname{Var}_{\tau}(f)=\left\lVert f-\mathbb{E}_{\tau}[f]\right\rVert_{2,\tau}^{2}\leq\frac{C_{b}}{\nu(K)}\langle f,-\mathcal{L}f\rangle_{\nu}\,\forall f\in\ell_{2,\nu}.$ (25)
Now, we combine Assumption 4 along with the drift assumption (Assumption 2) to obtain the Poincaré constant for the entire system, which is stated in the following Stitching Theorem.
Theorem 2.
(Stitching Theorem) Denote $K=\{q\in\mathcal{S}:b(q)>0\}$ and $\nu_{K}(q)=\frac{\nu(q)}{\nu(K)}\,\forall q\in K$. Under Assumptions 1, 2 and 4, the following inequality holds
$\operatorname{Var}_{\nu}(f)\leq\frac{1+\left(\sum_{q\in\mathcal{S}}b(q)\nu_{K}(q)\right)C_{b}}{\gamma}\langle f,-\mathcal{L}f\rangle_{\nu}\,\forall f\in\ell_{2,\nu},$ (26)
where $C_{b}$ is the weighted local Poincaré constant in Assumption 4, $\nu$ is the stationary distribution of the CTMC and $b:\mathcal{S}\rightarrow[0,\infty)$ is the positive drift term in Assumption 2.
The Stitching Theorem (Theorem 2) is applicable for irreducible, aperiodic, positive recurrent and reversible CTMCs. Our Stitching Theorem is a generalization of Theorem 1 in [71] for countable state space CTMCs and this generalization is crucial in getting the tight mixing rate in the Super-Halfin-Whitt regime all the way to $\alpha>1/2$ instead of $\alpha>2/3$. We defer this discussion to (32) and the proof of Theorem 2 to Appendix 8.2.1. For the special case of $b=B1_{K}$ for some finite $K\subseteq S$ and $B>0$, Assumption 4 gives the local Poincaré inequality
$\operatorname{Var}_{\nu_{K}}(f)\leq\frac{C_{L}}{\nu(K)}\langle f,-\mathcal{L}f\rangle_{\nu},$ (27)
where $\nu_{K}(x)=\frac{\nu(x)}{\sum_{x\in K}\nu(x)}\forall x\in K$. And so, this allows us to recover the result of [71] as stated below.
Corollary 7.
Under Assumptions 1 and 2, assume that $b=B1_{K}$ where $B>0$ and $K\subseteq S$ is some finite set. Also, assume that the system satisfies Assumption 4 with $b$ and constant $C_{L}>0$. Then, the CTMC admits the Poincaré constant $\frac{1+BC_{L}}{\gamma}$.
A version of Corollary 7 is known in [71, 68] for reversible Markov chains and [21, 66] for Markov processes. The readers can find the proof of Corollary 7 in Appendix 8.2.2, which is by applying the result of Theorem 2 for $b=B1_{K}$. The constant $C_{L}$ is called the local Poincaré constant w.r.t. the finite set $K$ of the CTMC. Intuitively, the CTMC mixes fast outside the finite set $K$ due to the negative drift, in addition to the rapid mixing within $K$ ensured by the “local” Poincaré inequality. In many applications, establishing a tight characterization of the local mixing behavior is the key to obtaining the tight mixing rate, and previously, the local Poincaré constant is predominantly obtained using the minorization condition [68, 23, 71]. However, it is known that the approach using minorization condition often yields sub-optimal bounds, and the minorization condition itself is usually hard to obtain as well [63, 1]. Moreover, whenever the bounded set is not a singleton, there will be a ”stitching” error when combining the mixing behavior inside the finite set with the mixing behavior outside the finite set together. Therefore, our Stitching Theorem (Theorem 2) is a necessary generalization to achieve a tight characterization of the local mixing behavior inside the finite set, whereas a naive application of Corollary 7 would not yield such a result. For a more detailed explanation, we refer the readers to the discussion around Equation (32).
Comparing Corollary 7 with Proposition 7, one can see that when we try to ”stitch” negative drift outside of the finite set $K$ with the local mixing condition behavior inside the finite set, we incur a ”stitching error” term $BC_{L}$ which hurts our Poincaré constant. Ideally, to get a tight mixing bound, we would like to make this stitching error term to be as small as possible or vanishes to 0 as $n$ goes to infinity. And so, it is crucial to choose the right Lyapunov function so that we get that right negative drift rate $\gamma$ while minimizing $B$, as well as developing methods to get tight local mixing constant $C_{L}$. We shall elaborate more on these steps in the following Section 4.
4 Proof sketch of mixing results
Recall that the $M/M/n$ queue experiences a phase transition at the Halfin-Whitt regime, i.e. $\alpha=\frac{1}{2}$ [32]. And so, to prove our main results, we will take different approaches and construct different Lyapunov functions for each of these regimes. The intuitions behind our approaches and the reasons why these regimes require two different approaches will become apparent as we go through the subsections below.
Figure 1: Visualization of chosen Lyapunov functions for the Super-Halfin-Whitt regime, Sub-Halfin-Whitt regime and the Mean Field regime. Figure 1(a) corresponds to the piecewise exponential Lyapunov function in the Super-Halfin-Whitt regime. Figure 1(b) corresponds to the linear and exponential Lyapunov function in the Sub-Halfin-Whitt regime. Figure 1(c) corresponds to the constant and exponential Lyapunov function in the Mean-Field regime.
4.1 Regime 1: Super-Halfin-Whitt Regime
Let $K=\{q\in S:b(q)>0\}$, recall that the Lyapunov-Poincaré machinery requires the following:
  • Step 1: Obtain the correct Lyapunov drift outside of some finite set $K$, which ensures rapid mixing outside $K$.
  • Step 2: Obtain local mixing guarantees inside the finite set $K$ by exploiting the properties of $K$ and $b$.
In Step 1, we wish to choose an appropriate Lyapunov function to correctly capture the drift. To gain some intuition, let’s start with the $M/M/1$ queue. The $M/M/1$ queue has the negative drift $\lambda-\mu=-\varepsilon$ everywhere except when the queue is empty. And thus, to obtain exponential convergence, we choose the Lyapunov function $V(q)=e^{\theta q}$. Intuitively, the fluid analog of the $M/M/1$ queue would then satisfy $\dot{V}\approx-\varepsilon\theta V$ providing the negative drift as required. More precisely, we obtain
$\mathcal{L}V\leq-\left(\sqrt{n}-\sqrt{\lambda}\right)^{2}V+b1_{\{0\}}$ (28)
where $b$ is some universal constant. The rate of $\left(\sqrt{n}-\sqrt{\lambda}\right)^{2}$ in the above equation matches previously established mixing results [73, 67] for the $M/M/1$ queue suggesting that $V$ is indeed the correct Lyapunov function. As seen in (28), we have a negative drift everywhere except the singleton $\{0\}$, so we simply apply Proposition 7 to obtain the Poincaré constant, which gives us the desired mixing rate.
For the $M/M/n$ queue, the dynamic is more complicated since there is another upward drift to $\lambda$ when $q<\lambda$ in addition to the downward drift when $q>\lambda$. So, we appropriately adjust our Lyapunov function and define it as follows:
$V(q)=e^{\theta\left[q-(n-1)\right]^{+}+\theta\left[(n-1)-q\right]^{+}}$ (29)
with appropriately chosen $\theta\geq 0$ to accurately capture the negative drift above $\lambda$ and the positive drift below $\lambda$. While $\lambda$ is the center of the drift, we consider the distance of the queue length to n1n-1 in the exponent because of algebraic convenience in the proof. From this choice of Lyapunov function, we obtain the following drift lemma for the Super-Halfin-Whitt regime.
Lemma 1.
Let $V$ be defined as in (29) and $\mathcal{L}$ be the generator of the $M/M/n$ queue in the super Halfin-Whitt regime (i.e. $\alpha>1/2$) with arrival rate $\lambda_{n}=n-n^{1-\alpha}$. Then, there exists $\gamma_{n}\geq\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}$ and a function $b:\mathbb{Z}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ such that:
$\mathcal{L}V(q)\leq-\gamma_{n}V(q)+b(q)1_{K}\,\forall q\in\mathbb{Z}_{\geq 0}$ (30)
where $K=\{[\lfloor 2\lambda_{n}\rfloor-n]^{+},...,n-1\}$. Furthermore, we have $b(q)\leq L\,\forall q\in K$ for some constant $L>0$ and $2\leq n\leq 7$ and for $n>7$, we have
$b(q)=\begin{cases}\frac{n^{1-2\alpha}}{1-n^{-\alpha}}e^{\frac{n^{1-2\alpha}}{1-n^{-\alpha}}}\,\,&\forall q\in K-\{n-1\}\\ n^{1-\alpha}\bigg{(}1+\frac{1}{4(n^{\alpha}-1)}\bigg{)}&\text{if }q=n-1\\ 0\,\,&\forall q\not\in K.\end{cases}$ (31)
The above lemma completes Step 1 of our methodology by establishing a negative drift with rate $\gamma_{n}$ outside the finite set $K$, the detail proof of this Lemma is deferred to Subsection 5.1. Unlike [71], the function $b$ in our drift condition (30) is not a constant function, which allows a more refined analysis and precipitates the use of our Stitching Theorem 2. To understand why the chosen Lyapunov function works for the $M/M/n$ queue, consider the following ODE that is a fluid analog of $M/M/n$ queue: $\dot{q}=\lambda_{n}-\min\{q,n\}$. When $q\geq\lambda_{n}+n^{1-\alpha}$ and $\theta_{1}\approx n^{-\alpha}$, we have $\dot{q}=\lambda_{n}-n=-n^{1-\alpha}$ which gives $\dot{V}\leq-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V=-\Theta(n^{1-2\alpha})V$, which is the desired negative drift. Similarly, when $q\leq\lambda_{n}-n^{1-\alpha}$, we also have $\dot{V}\leq-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V$. This allows us to establish a negative drift with rate $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ outside of a set $K$ with cardinality at most $\lceil 2n^{1-\alpha}-1\rceil$.
Observe that for $\alpha\geq 1$ or $n=1$, $K$ is a singleton as $2n^{1-\alpha}-1\leq 1$. In such a case, we simply apply Proposition 7 to obtain the desired mixing rate $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ without incurring any stitching error. For $\alpha\in(1/2,1)$, however, $K$ is no longer a singleton. And so, we come to Step 2, where we wish to show that the system mixes quickly in the finite set $K$ by showing that the system admits a local Poincaré inequality inside the finite set. To this end, we apply the local canonical path method (described in Lemma 6) to establish the local Poincaré constant $C_{L}=\Theta(n^{1-2\alpha})$ in Lemma 2, which means that $C_{L}$ goes to zero as $n$ goes to infinity and allows us to show rapid mixing in $K$.
However, note that from (31), if we choose $B=\max_{q\in\mathbb{Z}_{\geq 0}}b(q)$ and then naively apply Theorem 1 in [71] (also Corollary 7 in our work), we get $B=b(n-1)=O(n^{1-\alpha})$. Combining with the fact that $C_{L}=\Theta(n^{1-2\alpha})$, we would get the Poincaré constant
$C_{P}=\frac{1+O\left(n^{2-3\alpha}\right)}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}$ (32)
which matches the limiting behavior as $n\rightarrow\infty$ only when $\alpha\in(2/3,\infty)$ rather than for the entire Super-Halfin-Whitt regime $\alpha\in(1/2,\infty)$. Thus, naively applying previously proposed ”stitching theorems” would not yield the desired bound. To address this gap, we need a more refined local mixing analysis, by treating the $b$ term carefully as follows.
Lemma 2.
(Weighted Poincaré for $M/M/n$ in the Super-Halfin-Whitt regime) Under Assumptions 1, 2, let $b$ be given by (31), and $K=\{\lfloor 2\lambda_{n}\rfloor-n,...,n-1\}$. In addition, denote $\nu_{n}$ as the stationary distribution, we have Assumption 4 holds with constant $C_{b}=\Delta_{1}n^{2-4\alpha}+\Delta_{2}n^{3-6\alpha}$ holds for $\alpha\in[1/2,1)$ and for all test function $f\in\ell_{2,\nu_{n}}$ and for some positive constants $\Delta_{1},\Delta_{2}$.
Since $\alpha>1/2$, we have $C_{b}\rightarrow 0$ as $n\rightarrow\infty$, which allows our finite-time mixing bound to match the limiting behavior at the asymptotic. The proof and the precise characterization of the constant $C_{b}$ is reserved in Appendix 5.2.3, which is done by a clever application of the local canonical path method. In the final step, we combine the results in Lemma 1 and Lemma 2 and apply Theorem 2 to obtain the results in Equation (6) and Equation (7) in Theorem 1 and Proposition 2 as follows.
Finally, we bound the $\sum_{q\in K}b(q)\nu_{K}(q)$ term in Theorem 2, which is no more than a constant.
Claim 1.
Let $K=\{\lceil 2\lambda_{n}\rceil-n,...,n-1\}$, we have $\sum_{q\in K}\tau(q)\leq g_{3}(n)$. Here, we have $g_{3}(n)\leq 3.48\,\forall n\geq n_{0}$ and $\lim_{n\rightarrow\infty}g_{3}(n)=1/2$ for $\alpha>1/2$ and $\lim_{n\rightarrow\infty}g_{3}(n)=e/2+1$ for $\alpha=1/2$.
With Lemma 1, Lemma 2 and Claim 1 (whose proof is in Appendix 9.1.2), we obtain the mixing rate of $M/M/n$ in the Super-Halfin-Whitt regime and $M/M/1$ as follows.
Proof of Theorem 1 for $\alpha\in(1/2,1)$ and Proposition 2:.
We will consider two separate cases: the case where $\alpha\geq 1$ or $n=1$ and the case where $\alpha\in(1/2,1)$.
Case 1 ($\alpha\geq 1$ or $n=1$): From Proposition 7, Lemma 1 and observe that in this case, we have $K=\{n-1\}$, we have that the system admits the Poincaré constant $C_{P}=\frac{1}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}$. And so, from Proposition 6, we have
$\chi(\pi_{n,t},\nu_{n})\leq e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n}).$ (33)
Case 2 ($\alpha\in(1/2,1)$): From Lemma 1, Lemma 2, Claim 1 and Theorem 2, we obtain the Poincaré constant to be:
$C_{P}$ $=\frac{1+\left(\sum_{q\in K}b(q)\nu_{K}(q)\right)C_{b}}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}\leq\frac{1+3.48\times\left(384n^{2-4\alpha}+395.93n^{3-6\alpha}\right)}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}$
Let $C_{n}=\frac{1}{1+3.48\times\left(384n^{2-4\alpha}+395.93n^{3-6\alpha}\right)}$, we have that $\lim_{n\rightarrow\infty}C_{n}=1$ since $\alpha>1/2$. From Proposition 6, we have
$\chi(\pi_{n,t},\nu_{n})\leq e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})$ (34)
where $\lim_{n\rightarrow\infty}C_{n}=1$. ∎
Remark: Previously, several works have attempted different methods to establish the local Poincaré inequality. Previous works [71, 68] rely on the minorization assumption to establish some form of local mixing inside $K$, i.e. a local Poincaré inequality as in (27). However, such an approach does not directly work in our setting since the assumption requires every state in the state space to be within reach of any other state, which clearly does not hold for a birth and death process. Such an assumption can be satisfied if one looks at an $m$-step transition matrix for a large enough $m>0$. However, such an approach usually results in a poor minorization constant. In addition, a few other works have obtained local Poincaré inequalities using the diameter of the finite set, such as [66, 20], but generally, the local Poincaré constant bounds in these works are loose. Thus, as our goal is to obtain a tight characterization of the mixing rate, we take a different approach and propose a local version of the canonical path method that exploits the birth and death structure of the transition matrix. In fact, the non-local version has been used to bound the mixing time of random walks on different graph topologies [6].
In summary, the mixing rate $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ in the regime $\alpha\geq 1$ is tight and matches the result in [86]. In addition, the mixing rate for the regime $\alpha\in(1/2,1)$ is asymptotically tight, i.e., it matches with $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ for large $n$. For $\alpha=1/2$, the Poincaré constant inside the finite set $K$ does not vanish. We show that it is bounded above by a constant implying that the mixing rate is bounded below by a universal constant.
4.2 Regime 2: Sub-Halfin-Whitt Regime
As we saw before in Proposition 3, the $M/M/n$ queue with $\alpha\in(0,1/2)$ behaves similarly to the $M/M/\infty$ queue and we use this connection to motivate our Lyapunov function. Thus,, we first discuss the proof strategy for the $M/M/\infty$ queue, followed by that of the $M/M/n$ queue in the Sub-Halfin-Whitt regime.
4.2.1 $M/M/\infty$ system
For an $M/M/\infty$ system with arrival rate $\lambda$ and service rate $\mu$, it is known that the system is a discrete space diffusion analogue of the Ornstein–Uhlenbeck (OU) process [12]. For the latter process, it is also known that it admits the Gaussian Poincaré inequality with constant $1$, which implies unit mixing rate for the OU process and also suggests that the mixing rate for $M/M/\infty$ is $\mu$. To construct an appropriate Lyapunov function, one can look at the eigenfunction that corresponds to the correct mixing rate, which is also the smallest positive real number such that the spectral measure has a jump. For $M/M/\infty$, this eigenfunction is the first non-trivial Charlier polynomial $x-\frac{\lambda}{\mu}$ [19]. Hence, with an appropriate centering, we choose $V(q)=\left|q-\frac{\lambda}{\mu}\right|$ as the Lyapunov function so that it captures the correct eigenvalue equation $\mathcal{L}V(q)=-\mu V(q)$ for all $q$ outside of some finite set $K$ and satisfies the non-negativity condition for Lyapunov functions in Assumption 2.
4.2.2 $M/M/n$ system in the Sub-Halfin-Whitt regime
Observe that for an $M/M/n$ queue, the system behaves like an $M/M/\infty$ queue when $q<n$. Similarly, for $q>n$, the system behaves like an $M/M/1$ queue. And so, we stitch an exponential function for $q>n$ along with $|q-\lambda_{n}|$ for $q<n$ to obtain the following drift lemma.
Lemma 3.
Let $\{\lambda_{n}\}_{n\in\mathbb{Z}^{+}}$ be the sequence of arrival rates of the corresponding $M/M/n$ system such that $\lambda_{n}\in(0,n)\forall n\in\mathbb{Z}^{+}$ and let $\alpha_{n}=1-\frac{\log(n-\lambda_{n})}{\log n}$. Set $V(q)=\zeta e^{\theta(q-\lambda_{n})}\forall q>n$. If $\lambda_{n}\in\mathbb{Z}_{\geq 0}$, set $V(q)=|q-\lambda_{n}|\forall q\leq n$. Otherwise, set
$V(q)=\begin{cases}|q-\lambda_{n}|\,&\forall q\leq n,q\not\in\{\lfloor\lambda_{n}\rfloor,\lceil\lambda_{n}\rceil\}\\ |\lfloor\lambda_{n}\rfloor-\lambda_{n}-1|&\textit{if }q=\lfloor\lambda_{n}\rfloor\\ |\lceil\lambda_{n}\rceil-\lambda_{n}+1|&\textit{if }q=\lceil\lambda_{n}\rceil\end{cases}$ (35)
where $\zeta$ is chosen such that $|n-\lambda_{n}|=\zeta e^{\theta(n-\lambda_{n})}$ and $\theta>0$ is a parameter to be chosen. Then, for $\alpha_{n}\in(0,1/2)\,\forall n\in\mathbb{Z}^{+}$ and $\lim_{n\rightarrow\infty}\alpha_{n}=\alpha\in(0,1/2)$ for some constant $\alpha$, we have there exists $\gamma_{n},b$ with $\gamma_{n}\rightarrow 1$, such that
$\mathcal{L}V\leq-\gamma_{n}V+b1_{K}$ (36)
where $b\leq 2\lambda_{n}$ and $K=\{\lambda_{n}\}$ if $\lambda_{n}\in\mathbb{Z}_{\geq 0}$ and $b\leq\lceil\lambda_{n}\rceil+2$ and $K=\{\lfloor\lambda_{n}\rfloor-1,\lfloor\lambda_{n}\rfloor,\lceil\lambda_{n}\rceil,\lceil\lambda_{n}\rceil+1\}$ otherwise.
The proof of this Lemma is deferred to Appendix 9.3.1. Using the above lemma, the finite set is a singleton when $\lambda_{n}\in\mathbb{Z}_{\geq 0}$, so we directly use Proposition 7 to infer the Poincaré constant, which in turn gives us the mixing rate. Observe that as $\gamma_{n}\rightarrow 1$ in the above lemma, we asymptotically match the mixing rate of the $M/M/\infty$ queue. However, when $\lambda_{n}\notin\mathbb{Z}_{\geq 0}$, the finite set ceases to be a singleton. Nonetheless, $|K|=4$ is independent of $n$, and so we obtain a $\Omega(1/n)$ local Poincaré constant using the canonical path method since there are $\Theta(n)$ transitions per time unit. Although this is not enough to overcome the stitching error to get a mixing rate that approaches $1$, combining this local Poincaré constant with our drift lemma above using the Stitching Theorem (Theorem 2) is sufficient to obtain an $\Omega(1)$ mixing rate. We defer the full proof to Appendix 9.3.
4.3 Regime 3: Mean Field regime
In this subsection, we outline the proof of Proposition 4, which establishes a tight mixing bound for $\lambda_{n}\notin\mathbb{Z}_{\geq 0}$, but with an additional assumption that $\lambda_{n}\ll n$. As the result in the previous subsection did not yield a tight mixing rate for $\lambda_{n}\notin\mathbb{Z}_{\geq 0}$, we propose the following alternative Lyapunov function: $V(q)=e^{\theta\left[q-n\right]^{+}}$ with $\theta\approx\varepsilon$. This choice does not align with the intuition to engineer $V(q)\approx|q-\lambda_{n}|$ for $q<n$ to mimic the $M/M/\infty$ behavior. Indeed, analyzing the drift of $V(q)=e^{\theta\left[q-n\right]^{+}}$ results in a finite set $K=\{0,1,\ldots,n\}$, which is much larger than what we previously had. However, $V(q)$ allows us to exploit the strong drift for $q>n$. Indeed, we establish the drift inequality (23) with $\gamma=\Theta(n^{1-2\alpha})\rightarrow\infty$ as $n\rightarrow\infty$. We then exploit the connection to the $M/M/\infty$ queue to obtain a local Poincaré inequality corresponding to the set $K=\{0,1,\ldots,n\}$ using the following lemma:
Lemma 4.
Let $K=\{m,...,M\}$ be a connected subset of the state space $\mathcal{S}\subset\mathbb{Z}$ of the birth and death process, $\nu$ be the stationary distribution of the CTMC and $\nu_{K}(x)=\nu(x)/\sum_{x\in K}\nu(x)\,\forall x\in\mathcal{S}$. Furthermore, let $\mathcal{L}$ be the generator of the birth and death process and assume that $\operatorname{Var}_{\nu}(f)\leq C_{P}\langle f,-\mathcal{L}f\rangle\nu\,\forall f\in\ell_{2,\nu}$, we have
$\operatorname{Var}_{\nu_{K}}(f)\leq C_{P}\sum_{x:x,x+1\in K}Q_{K}(x,x+1)\left(f(x)-f(x+1)\right)^{2}\,\forall f\in\ell_{2,\nu_{K}}$ (37)
where $Q_{K}(x,y)=\nu_{K}(x)\mathcal{L}(x,y)\,\forall x,y\in\mathcal{S}$.
In words, the above lemma allows us to obtain local mixing results if we know the mixing behavior of another system that subsumes the dynamic inside the finite set. The $M/M/\infty$ queue indeed subsumes the dynamics inside the finite set $K\in\{0,1,\ldots,n\}$. The above lemma then allows us to establish the local Poincaré constant of $\approx 1$. We call this approach the “truncation argument” to obtain a local Poincaré inequality. We then apply Theorem 2 and use the fact $\lambda_{n}\ll n$ to conclude that the Poincaré constant for the $M/M/n$ queue is $\approx 1$ which completes the proof. The proof details of Proposition 4 and Lemma 4 are deferred to Appendix 9.4 and 9.4.2 respectively.
5 Proof details of key technical lemmas
To help readers better understand how our Lyapunov-Poincaré machinery work, we shall provide a detailed proof for the Super-Halfin-Whitt with its key technical lemmas as follows. Recall in our proof sketch in Subsection 4.1 that we need to establish Lyapunov drift lemma (namely Lemma 1) and obtain the local mixing bound (Lemma 2), which consists of the following main ingredients:
  1. 1.
    Show that the stationary distribution in $K$ is roughly uniform.
  2. 2.
    Prove the local canonical path method.
  3. 3.
    Apply the canonical path method to obtain the weighted local mixing bound.
Similarly for the Sub-Halfin-Whitt regime, one can apply a similar recipe to obtain the mixing bound for the system in that regime with an appropriate Lyapunov function.
5.1 Proof of the Lyapunov drift (Lemma 1)
Proof.
Recall that $V(q)=e^{\theta[q-(n-1)]^{+}+\theta[(n-1)-q]^{+}}$, where we fix $\theta=\log\sqrt{\frac{n}{\lambda_{n}}}$. Now, when $q\geq n$, we have $V(q)=e^{\theta(q-(n-1))}$. We then have
$\mathcal{L}V(q)$ $=\left(\lambda_{n}e^{\theta}+\frac{n}{e^{\theta}}-(\lambda_{n}+n)\right)V(q)=\left(e^{\theta}-1\right)\left(\lambda_{n}-\frac{n}{e^{\theta}}\right)V(q)=-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V(q).$ (38)
The choice of $\theta$ is evident in the third equality as it minimizes $\left(e^{\theta}-1\right)\left(\lambda_{n}-\frac{n}{e^{\theta}}\right)$. For $q<n$, we have $V(q)=e^{\theta(n-1-q)}$ and now consider two cases: $n=1\leq n\leq 7$, and $n\geq 8$.
Case 1 $(1\leq n\leq 7)$: In this case, we have $K=\{0,1,...,n-1\}$. From (38), we have $\mathcal{L}V\leq-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V\,\,\forall x\not\in K$. If $x\in K$ then $q\leq n-1$ and since $n\leq 7$ we have that there is only a finite number of pairs $(n,q)$. Thus, we have
$\mathcal{L}V(q)\leq-\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}V(q)+L1_{K}$
for some constant $L$, where it suffices to consider the maximum value of $\mathcal{L}V(q)$ taken over all choice of $q\in K$ and $1\leq n\leq 7$. Hence, the case $1\leq n\leq 7$ is done.
Case 2 $(n>7)$: We have $K=\{\lfloor 2\lambda_{n}\rfloor-n,\ldots,n\}$. This gives $V(q)=e^{\theta(n-1-q)}$ for $q\leq n-1$. So, for q<n1q<n-1, we have
$\mathcal{L}V(q)$ $=\left(e^{\theta}-1\right)\left(q-\frac{\lambda_{n}}{e^{\theta}}\right)V(q)=\left(e^{\theta}-1\right)\bigg{(}q-(2\lambda_{n}-n)+\underbrace{2\lambda_{n}-n-\frac{\lambda_{n}}{e^{\theta}}}_{\leq\lambda_{n}-\sqrt{\lambda_{n}n}}\bigg{)}V(q)$
$\leq-\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}V(q)+\left(e^{\theta}-1\right)\left(q-(2\lambda_{n}-n)\right)V(q)$
For $2\lambda_{n}-n\leq q\leq n-2$ and $\alpha>1/2$, we have $V(q)=e^{\theta(n-1-q)}\leq e^{2\theta n^{1-\alpha}}$. Furthermore, we have
$\theta\leq e^{\theta}-1=\sqrt{\frac{n}{\lambda_{n}}}-1=\frac{\frac{n-\lambda_{n}}{\lambda_{n}}}{\sqrt{\frac{n}{\lambda_{n}}}+1}\leq\frac{n^{1-\alpha}}{\sqrt{n\lambda_{n}}+\lambda_{n}}=\frac{n^{-\alpha}}{\sqrt{1-n^{-\alpha}}+1-n^{-\alpha}}\leq\frac{1}{2(n^{\alpha}-1)}.$ (39)
Thus, we have $V(q)\leq e^{\frac{n^{1-\alpha}}{n^{\alpha}-1}}$ from (39). And so for $q\in K-\{n-1\}$
$\underbrace{\left(e^{\theta}-1\right)}_{O(n^{-\alpha})}\underbrace{\left(q-(2\lambda_{n}-n)\right)}_{O(n^{1-\alpha})}\underbrace{V(q)}_{O(1)}\leq\frac{1}{2(n^{\alpha}-1)}\times 2n^{1-\alpha}\times e^{\frac{n^{1-\alpha}}{n^{\alpha}-1}}\leq\frac{n^{1-2\alpha}}{1-n^{-\alpha}}e^{\frac{n^{1-2\alpha}}{1-n^{-\alpha}}}=b(q).$
And so, we have
$\mathcal{L}V(q)\leq-\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}V(q)+\frac{n^{1-2\alpha}}{1-n^{-\alpha}}e^{\frac{n^{1-2\alpha}}{1-n^{-\alpha}}}.$ (40)
For q=n1q=n-1, we have
$\mathcal{L}V(n-1)$ $=(\lambda_{n}+n-1)\left(e^{\theta}-1\right)\leq 2\lambda_{n}\times\frac{1}{2(n^{\alpha}-1)}=n^{1-\alpha}$
$=-\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}\underbrace{V(n-1)}_{=1}+\bigg{(}n^{1-\alpha}+\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}\bigg{)}$
$\leq-\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}V(n-1)+n^{1-\alpha}\bigg{(}1+\frac{1}{4(n^{\alpha}-1)}\bigg{)}$
Hence we have for $q\leq n$:
$\mathcal{L}V(q)\leq-\gamma_{n}V(q)+b(q)1_{K}.$
with $\gamma_{n}=\left(\sqrt{n}-\sqrt{\lambda_{n}}\right)^{2}$ and $b$ as defined in (31), which completes the proof. ∎
5.2 Proof of Lemma 2
In this subsection, we will establish the $b$-weighted Poincaré results where $b$ is the positive drift term in Lemma 1.
5.2.1 The stationary distribution inside $K$ is roughly uniform
To get a control on the finite set behavior, we first need to study the steady state distribution inside the finite set, that is to show that $\nu_{K}$ is a roughly uniform distribution whose support is the set $K$. This lemma lays the groundwork for us to apply the local canonical path method.
Lemma 5.
Let $\lambda_{n}=n-n^{1-\alpha},K=\{\lfloor 2\lambda_{n}\rfloor-n,...,n-1\}$ and $\nu_{K}(x)=\frac{\nu(x)}{\sum_{x\in K}\nu(x)}\forall x\in K$ and $\nu_{K}(x)=0$ otherwise. For $\alpha\in(1/2,1)$ and $n\geq 110$, we have $L_{K}(n)\times n^{\alpha-1}\leq\nu_{K}(x)\leq U_{K}(n)\times n^{\alpha-1}$ where $\lim_{n\rightarrow\infty}L_{K}(n)=\lim_{n\rightarrow\infty}U_{K}(n)=1/2$.‘
The main idea of the proof is to apply Stirling’s approximation and appropriate bounds. We defer the full statement and the full proof of the lemma to Appendix 9.1.1.
5.2.2 Local canonical path lemma
To this end, we first need to develop a toolkit to analyze the local mixing property of the birth-and-death process in a bounded state space.
Lemma 6.
(Local canonical path method) For any birth and death process with the generator $\mathcal{L}$, a subset $K$ of the state space $\mathbb{Z}_{\geq 0}$ and the stationary distribution $\nu$, we have the following inequality holds:
$\operatorname{Var}_{\nu_{K}}(f)\leq C_{L}\sum_{k:k,k+1\in K}Q_{K}(k,k+1)(f(k+1)-f(k))^{2}$ (41)
where the constant $C_{L}$ is defined as:
$C_{L}=\max_{k:k,k+1\in K}\frac{\sum_{x,y\in K:x\leq k\leq y-1}|x-y|\nu_{K}(x)\nu_{K}(y)}{Q_{K}(k,k+1)}\,\forall f\in\ell_{2,\nu_{K}}$ (42)
where $\nu_{K}(x)=\frac{\nu(x)}{\sum_{x\in K}\nu(x)}\,\forall x\in K$ and $\nu_{K}(x)=0$ otherwise and $Q_{K}(u,v)=\nu_{K}(u)\mathcal{L}(u,v)=\nu_{K}(v)\mathcal{L}(v,u)\forall u,v\in\mathbb{Z}_{\geq 0}$.
Here, $C_{L}$ can also be understood as the congestion ratio, which is lower bounded by the inverse mixing rate [40]. When $\nu_{K}$ is roughly uniform and the transitions are roughly symmetric, the constant $C_{L}$ is proportion to the sum of the path length of the path passes through an edge $e=(k,k+1)$ that maximizes the constant. As the name suggests, we prove this lemma by analyzing the path structure as follows.
Proof.
Proof of Lemma 6: We emulate the proof for the canonical path method as in [6] to handle the continuous-time Markov chain. Denote $\gamma_{x,y}$ as one of the paths connecting $x,y$, note that $\mathcal{L}(x,y)>0$ if and only if |xy|=1|x-y|=1, thus all path in $K$ is a sequence of adjacent states. For any test function $f\in\ell_{2,\nu_{K}}$ and let $Q_{K}(x,y)=\nu_{K}(x)\mathcal{L}(x,y)\forall x\neq y$, we have:
$\operatorname{Var}_{\nu_{K}}(f)$ $=\sum_{x<y\in K}\nu_{K}(x)\nu_{K}(y)\left(f(x)-f(y)\right)^{2}$
$\overset{(a)}{\leq}\sum_{x<y\in K}|x-y|\nu_{K}(x)\nu_{K}(y)\left(\sum_{x\leq k\leq y-1}(f(k)-f(k+1))^{2}\right)$
$=\sum_{x<y\in K}|x-y|\nu_{K}(x)\nu_{K}(y)\left(\sum_{x\leq k\leq y-1}\frac{(f(k)-f(k+1))^{2}Q_{K}(k,k+1)}{Q_{K}(k,k+1)}\right)$
$\overset{(b)}{\leq}\sum_{k:k,k+1\in K}\left(\frac{\sum_{x,y\in K:x\leq k\leq y-1}|x-y|\nu_{K}(x)\nu_{K}(y)}{Q_{K}(k,k+1)}\right)\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1).$
Here, (a) follows from the Cauchy-Schwarz inequality and (b) follows from the interchange of sums. Hence proved. $\square$
5.2.3 Proof of Lemma 2
Finally, observe that $f\in\ell_{2,\nu_{n}}$ trivially implies $f\in\ell_{2,\nu_{K}}$ and $f\in\ell_{2,\nu_{K}}$ is equivalent to $f\in\ell_{2,\tau}$ for $\tau(q)=b(q)\nu_{K}(q)\,\forall q\in\mathcal{S}$, given that $b$ is sufficiently well-behaved. This essentially tells us that if a test function $f$ is valid for the global distribution then it also works for the weighted distribution and the local distribution.
Proof.
Proof of Lemma 2: Let $\tau$ be a measure such that $\tau(q)=b(q)\nu_{K}(q)\,\forall q\in\mathbb{Z}_{\geq 0}$, where $b$ is defined as in (31). Note that from Lemma 1 and Lemma 5 and the fact that $n\geq n_{0}$, we have:
$\tau(q)$ $\leq\left(1+\frac{1}{4(n^{\alpha}-1)}\right)\frac{e^{\frac{1}{12\lfloor n-2n^{1-\alpha}\rfloor}}\sqrt{1-2n^{-\alpha}-n^{-1}}}{2\left(1-\frac{1}{(n^{\alpha}-2)^{2}}\right)^{n}}\text{ for }q=n-1,$ (43)
$\tau(q)$ $\leq\frac{e^{\frac{n^{1-2\alpha}}{1-n^{1-\alpha}}}}{1-n^{-\alpha}}\frac{e^{\frac{1}{12\lfloor n-2n^{1-\alpha}\rfloor}}\sqrt{1-2n^{-\alpha}-n^{-1}}}{2\left(1-\frac{1}{(n^{\alpha}-2)^{2}}\right)^{n}}n^{-\alpha}\text{ for }q\in K-\{n-1\}.$ (44)
In addition, we also have
$\tau(q)$ $\geq\left(1+\frac{1}{4(n^{\alpha}-1)}\right)\frac{\left(1-\frac{1}{(n^{\alpha}-2)^{2}}\right)^{n}\sqrt{1-2n^{-\alpha}-n^{-1}}}{e^{\frac{1}{12\lfloor n-2n^{1-\alpha}\rfloor}}\left(2+n^{\alpha-1}\right)}\text{ for }q=n-1,$ (45)
$\tau(q)$ $\geq\frac{e^{\frac{n^{1-2\alpha}}{1-n^{-\alpha}}}}{1-n^{-\alpha}}\frac{\left(1-\frac{1}{(n^{\alpha}-2)^{2}}\right)^{n}\sqrt{1-2n^{-\alpha}-n^{-1}}}{e^{\frac{1}{12\lfloor n-2n^{1-\alpha}\rfloor}}\left(2+n^{\alpha-1}\right)}n^{-\alpha}\text{ for }q\in K-\{n-1\}.$ (46)
From Lemma 1 and Lemma 5, we have
$\operatorname{Var}_{\tau}(f)$ $=\frac{1}{2}\sum_{x,y\in K}\tau(x)\tau(y)(f(x)-f(y))^{2}$
$=\underbrace{\sum_{x\in K}\tau(n-1)\tau(x)\left(f(n-1)-f(x)\right)^{2}}_{T_{1}}+\underbrace{\frac{1}{2}\sum_{x,y\in K-\{n-1\}}\tau(x)\tau(y)(f(x)-f(y))^{2}}_{T_{2}}.$
The terms $T_{1},T_{2}$ can be bounded by the following claims.
Claim 2.
Let $g_{1}:\mathbb{N}\rightarrow\mathbb{R}^{+}$ be a function to be defined later, we have $T_{1}\leq g_{1}(n)\times n^{2-4\alpha}\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1)$. Here, we have $g_{1}(n)\leq 384\forall n\geq n_{0}$ and
$\lim_{n\rightarrow\infty}g_{1}(n)$ =2 $=2\text{ for }\alpha>1$/2, (47)
$\lim_{n\rightarrow\infty}g_{1}(n)$ $=2e^{3}\text{ for }\alpha=1/2.$ (48)
Next, using the local canonical path method, we have the following claim:
Claim 3.
Let $g_{2}:\mathbb{N}\rightarrow\mathbb{R}^{+}$ be a function to be defined later, we have $T_{2}\leq g_{2}(n)\times n^{3-6\alpha}\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1)$. Here, we have $g_{2}(n)\leq 395.93\,\forall n\geq n_{0}$ and
$\lim_{n\rightarrow\infty}g_{2}(n)$ $=0\text{ for }\alpha>1/2,$ (49)
$\lim_{n\rightarrow\infty}g_{2}(n)$ $=e^{5}\text{ for }\alpha=1/2.$ (50)
The full proofs of these claims can be found in Appendix 9.1.3. Combining the claims gives us
$\operatorname{Var}_{\tau}(f)=T_{1}+T_{2}$ $\leq\left(g_{1}(n)n^{2-4\alpha}+g_{2}(n)n^{3-6\alpha}\right)\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1)$
$=\frac{g_{1}(n)n^{2-4\alpha}+g_{2}(n)n^{3-6\alpha}}{\nu_{n}(K)}\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}\nu_{n}(k)\mathcal{L}(k,k+1)$
$\leq\frac{g_{1}(n)n^{2-4\alpha}+g_{2}(n)n^{3-6\alpha}}{\nu_{n}(K)}\sum_{k=0}^{\infty}\left(f(k)-f(k+1)\right)^{2}\nu_{n}(k)\mathcal{L}(k,k+1)$
$=\frac{g_{1}(n)n^{2-4\alpha}+g_{2}(n)n^{3-6\alpha}}{\nu_{n}(K)}\langle f,-\mathcal{L}f\rangle_{\nu_{n}}$
holds for all $f\in\ell_{2,\nu_{n}}$. Hence proved. $\square$
6 Conclusion and Future Work
In this paper, we establish bounds on the Chi-squared distance between the finite-time and the steady-state queue length distribution for the $M/M/n$ queue. We demonstrate that the mixing rate is tight (possibly up to a constant) in the many-server-heavy-traffic regimes. We also obtain bounds on the mean, moments, and tail of the queue length for a finite time and finite $n$. To prove these results, we build a Lyapunov-Poincaré framework and also propose two different ways to obtain the local Poincaré inequality which are of independent interest. In particular, our proposed local canonical path method and the truncation method have the potential to provide stronger bounds than the classical drift and minorization method [23, 68].
While we focus on analyzing a birth and death chain in this paper, note that our Lyapunov - Poincaré methodology is flexible enough to be applied to any reversible CTMC. Generalizing our results to other reversible Markov chains would pose the main challenge of constructing a suitable Lyapunov function resulting in suitable negative drift outside a suitable finite set. Nonetheless, the results in Section 3 are versatile, and one can hope to employ this methodology to obtain mixing results beyond birth and death chains. Beyond queueing theory, our finite set approaches can be further applied to other areas such as Learning, Generative AI and Sampling [66, 84, 26, 13].
7 Acknowledgement
This work was partially supported by NSF grants EPCN-2144316 and CMMI-2140534 and the Georgia Tech ARC-ACO Fellowship. The authors thank Professor David Goldberg for insightful discussions and for suggesting relevant references. The authors also kindly thank the anonymous reviewers for their helpful reviews and feedback.